### v2

# Teleportation=Translation: Continuous recovery of black hole information

Abstract
[…]
1 Introduction
The black hole information paradox exposes a fundamental tension between the apparent thermality of Hawking radiation, as obtained in semiclassical gravity, and the unitary evolution required by quantum theory Hawking1976 . Although Hawking’s original calculation indicates that a semiclassical black hole should evaporate through thermal radiation in a manner that seems to violate unitarity, a number of recent developments suggest the possibility of information recovery even within the semiclassical framework, provided the paradox is analyzed with more refined conceptual and technical tools Almheiri2020 ; Raju2022 ; Akil:2025bhip .
As a primary example of such tools, AdS/CFT offers a semiclassical framework that uses replica wormholes and quantum extremal surfaces to successfully derive the Page curve Almheiri2019 ; Almheiri2020Page ; Penington2022 . This result strongly suggests that unitarity is preserved, although it has raised difficult questions about whether gravity theories require ensemble averaging GhoshKrishnan2021 or how the factorization of spacetime works HarlowJafferis2020 ; Marolf2020 . To address the mathematical foundations of these results, quantum information and operator-algebraic analyses have been used to clarify the precise assumptions behind the entropy computations Raju2022 ; Kibe2022 .
While these analyses clarify the entropy computations, calculating the Page curve is distinct from explaining the physical mechanism of retrieval. Focusing on the specific mechanism of retrieval, Hayden and Preskill reformulated the problem as a quantum decoding task. They demonstrated that if a black hole is old and acts as a fast scrambler, any new information thrown into it can be retrieved rapidly, provided the observer has access to the early radiation HaydenPreskill . This recovery process is operationally very similar to quantum teleportation. Building on this operational viewpoint, van den Heijden and Verlinde (vdH-V) reformulated the information problem entirely within the language of operator algebras vdHV . In their framework, the canonical shift $\Gamma$ plays the role of a teleportation operator, leading to the identification Teleportation=Translation.
However, as acknowledged in vdHV , a direct realization of this protocol in the continuum limit faces fundamental obstructions. The local operator algebras in quantum field theory are Type III von Neumann factors, distinguished by the absence of a tracial state Connes80 ; Witten . This algebraic feature reflects the physical reality that vacuum entanglement is effectively unbounded. Consequently, the finite maximally entangled resources prerequisite for standard teleportation are mathematically ill-defined in this setting.
In this work, we construct a smooth unitary interpolation to connect the discrete canonical shift with continuous spacetime evolution. To handle the Type III structure, we employ the Haagerup–Kosaki framework Connes80 ; Kosaki84 and lift the inclusion to a Type II crossed product via Takesaki duality Connes80 ; TakesakiII . Since a trace-preserving conditional expectation is well-defined in this setting, we can construct a strongly continuous unitary path connecting the identity to the canonical shift. We define the continuous generator $\widetilde{G}$ as the unique self-adjoint operator associated with this flow.
Building on this construction, we derive our main result by proving that this generator is physically identified with the modular momentum. Specifically, we establish a rigorous identity asserting that the generator of the teleportation flow equals twice the difference of the modular Hamiltonians. The factor of 2 is dictated by Borchers’ commutation relations Borchers92 , ensuring consistency with the scaling of modular flows. Our proof establishes this equality on a dense core of analytic modular vectors and demonstrates the equality of the closed operators via Nelson’s analytic vector theorem Nelson59 . Furthermore, we quantify the stability of this identification under small deformations using non-commutative $L^{p}$-theory Kosaki84 ; Kato . Physically, this result provides a precise operator-algebraic realization of the Teleportation=Translation proposal, demonstrating that information recovery manifests as a continuous geometric translation. Finally, to facilitate the verification of this operator identity in complex physical models where direct analysis is challenging, we propose a concrete observable test based on correlation functions.
The remainder of this paper is organized as follows. In Sec. 2, we summarize the necessary elements of Tomita–Takesaki theory and introduce the Jones projection together with the canonical shift $\Gamma$. In Sec. 3, we construct the continuous unitary interpolation by lifting the Type III inclusion to the crossed product via the Haagerup–Kosaki framework. Sec. 4 is devoted to the proof of our main result, identifying the generator of this flow with the modular momentum ($\widetilde{G}=2P$). We conclude with a discussion of the results in Sec. 5.
Technical details are deferred to the appendices. Appendix A collects the notation used throughout the paper. Appendix B illustrates the failure of naive interpolation in finite dimensions and provides a concrete continuous unitary path. Finally, Appendix C details the Haagerup-Takesaki construction of faithful conditional expectations in Type III settings via operator-valued weights.
2 The Discrete Protocol: Canonical Shift
2.1 Preliminaries: Tomita–Takesaki Theory
We work on a fixed Gelfand-Naimark-Segal (GNS) Hilbert space $\mathcal{H}$ built from a faithful normal state $\omega$ (with cyclic and separating vector $|\Omega\rangle$) for the von Neumann algebra $\mathcal{M}$ TakesakiI . Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of von Neumann algebras. For an algebra $\mathcal{X}$ (where $\mathcal{X}=\mathcal{M}$ or $\mathcal{N}$), the core objects of Tomita–Takesaki theory TakesakiII are:
  • The Tomita operator $S_{\mathcal{X}}$, defined on the dense domain $\mathcal{X}|\Omega\rangle$ by $S_{\mathcal{X}}x|\Omega\rangle=x^{*}|\Omega\rangle$ for $x\in\mathcal{X}$.
  • Its polar decomposition $S_{\mathcal{X}}=J_{\mathcal{X}}\Delta_{\mathcal{X}}^{1/2}$, where $J_{\mathcal{X}}$ is the anti-unitary modular conjugation and $\Delta_{\mathcal{X}}$ is the positive self-adjoint modular operator.
  • The modular Hamiltonian $K_{\mathcal{X}}$, defined by $\Delta_{\mathcal{X}}=e^{-2\pi K_{\mathcal{X}}}$.
  • The modular automorphism group $\sigma_{t}^{\mathcal{X}}(a)=\Delta_{\mathcal{X}}^{it}a\Delta_{\mathcal{X}}^{-it}$.
These objects satisfy $J_{\mathcal{X}}\mathcal{X}J_{\mathcal{X}}=\mathcal{X}^{\prime}$, where $\mathcal{X}^{\prime}$ is the commutant of $\mathcal{X}$.
2.2 The Canonical Shift Protocol
The discrete vdH-V protocol is built from three components vdHV .
Step 1: Conditional Expectation. Let $E:\mathcal{M}\to\mathcal{N}$ be a normal, faithful conditional expectation that is $\omega$-preserving: $\omega(E(a))=\omega(a)$ for all $a\in\mathcal{M}$. This map effectively discards information in $\mathcal{M}$ that is not in $\mathcal{N}$.
Step 2: Jones Projection. In the GNS representation where $|a\rangle:=a|\Omega\rangle$, $E$ is implemented by the Jones projection $e_{N}$, an operator $e_{N}\in\mathcal{B}(\mathcal{H})$ satisfying
$e_{N}|a\rangle=|E(a)\rangle.$ (1)
$E$ being a $\omega$-preserving conditional expectation ensures that $e_{N}$ is a well-defined self-adjoint projection ($e_{N}^{2}=e_{N}$, $e_{N}^{\dagger}=e_{N}$) and satisfies $e_{N}ae_{N}=E(a)e_{N}$ for all $a\in\mathcal{M}$. The discrete (0 or 1) nature of this projection is the source of the protocol’s discreteness.
Step 3: Canonical Shift. The protocol is enacted by the unitary operator
$U_{\Gamma}:=J_{\mathcal{M}}J_{\mathcal{N}},$ (2)
which induces the automorphism (the shift) $\Gamma(a)=U_{\Gamma}aU_{\Gamma}^{\dagger}$. This map $\Gamma$ acts as a teleportation map, sending the relative commutant $\mathcal{N}^{\prime}\cap\mathcal{M}$ (information in $\mathcal{M}$ hidden from $\mathcal{N}$) to the relative commutant $\mathcal{M}^{\prime}\cap\mathcal{M}_{1}$ (information in the basic extension $\mathcal{M}_{1}:=\langle\mathcal{M},e_{N}\rangle$ hidden from $\mathcal{M}$):
$\Gamma(\mathcal{N}^{\prime}\cap\mathcal{M})=\mathcal{M}^{\prime}\cap\mathcal{M}_{1}.$ (3)
Information is not lost, but relocated from one relative commutant to another. Crucially, while explicitly defined algebraically, this shift operator $U_{\Gamma}$ corresponds to a geometric translation in the emergent spacetime—a connection that will be rigorously formulated in the continuum limit in the subsequent sections.
3 From a Discrete Protocol to a Continuous Path
Our core strategy is to soften the discrete Jones projection $e_{N}$ by constructing a continuous family of maps $E_{s}$ and corresponding operators $e_{N}(s)$ for $s\in[0,1]$, such that $E_{0}=\mathrm{id}$ and $E_{1}=E$. We first briefly discuss the challenge of (dynamical) idempotency in Sec. 3.1. Then, we present the pedagogical completely positive interpolation (Path A). Although this path ultimately fails the idempotency requirement in the infinite-dimensional (Type III) setting, discussing it is crucial for understanding the necessity of the more advanced construction (Sec. 3.2). We then lift the problem to the semifinite crossed-product (Haagerup–Kosaki envelope) and construct the canonical, physical interpolating path via $L^{p}$-space interpolation theory (Path B), satisfying the rigorous boundary conditions required for our proof (Sec. 3.3). Finally, using the Tomita–Takesaki machinery in the semifinite setting, we build the unitary path $\widetilde{U}(s)$ and rigorously establish the generator $\widetilde{G}$ as a tangent vector, proving its self-adjointness via analytic vectors (Sec. 3.4).
3.1 The Physical Requirement: Dynamical Idempotency
To define a continuous unitary path $U(s)=J_{\mathcal{M}}J_{\mathcal{N}(s)}$, we must first construct a continuous path of von Neumann subalgebras $\mathcal{N}(s)$ such that $\mathcal{N}(0)=\mathcal{M}$ and $\mathcal{N}(1)=\mathcal{N}$. Such a path of algebras would, in turn, define a path of conditional expectations $E_{s}:\mathcal{M}\to\mathcal{N}(s)$. The crucial link between the map $E_{s}$ and the subalgebra $\mathcal{N}(s)$ is idempotency. Furthermore, for the parameter $s$ to be interpreted as a physical coarse-graining or information resolution flow, one should expect a condition even stronger than simple “static” idempotency ($E_{s}^{2}=E_{s}$). A true physical flow should satisfy a “dynamical” idempotency or semigroup property, such as $E_{s^{\prime}}\circ E_{s}=E_{s^{\prime}}$ for $s^{\prime}\geq s$. This condition, common in processes like the renormalization group, ensures the path is consistent: coarse-graining to level $s$ and then further to $s^{\prime}$ is identical to coarse-graining directly to $s^{\prime}$. While this dynamical property implies the static one, the static property ($E_{s}^{2}=E_{s}$) remains the minimum mathematical requirement for $E_{s}$ to be a conditional expectation at all, as established by the following theorem (Tomiyama theorem Tomiyama57 ).
Definition 3.1.
Dynamic idempotency for a family of maps $\{E_{s}\}_{s\in I}$ means:
$E_{s}^{2}=E_{s}\qquad\text{(static idempotency)},$
$E_{s^{\prime}}\circ E_{s}=E_{s^{\prime}}\quad\text{for }s^{\prime}\geq s\qquad\text{(semigroup/absorption property)}.$ (4)
While static idempotency is sufficient to define a projection at a single point, the dynamical condition is essential for interpreting the path as a coherent teleportation protocol. Physically, this ensures that the sequence of coarse-graining operations is consistent: projecting to an intermediate scale $s$ and then to a coarser scale $s^{\prime}$ must be identical to projecting directly to $s^{\prime}$. In the context of information recovery, this implies a strict nesting of information. Without this property, the intermediate stages of the protocol would lack a hierarchical structure, meaning that recovered information at step $s$ could be inexplicably lost or scrambled in subsequent steps. Thus, dynamical idempotency is the rigorous condition that prevents information leakage and ensures the monotonicity of the recovery flow from the black hole interior ($\mathcal{N}$) to the exterior ($\mathcal{M}$).
Theorem 3.2.
(Idempotent CP maps are conditional expectations) Let $\mathcal{M}$ be a von Neumann algebra and $\omega$ a faithful normal state on $\mathcal{M}$. Let $E:\mathcal{M}\to\mathcal{M}$ be a normal, unital, completely positive (CP) map satisfying $E^{2}=E$ and $\omega\circ E=\omega$. Then $\mathcal{N}:=E(\mathcal{M})$ is a von Neumann subalgebra of $\mathcal{M}$, and $E$ is the faithful normal conditional expectation from $\mathcal{M}$ onto $\mathcal{N}$.
This theorem sets our target: we must find a continuous path of idempotent CP maps.
3.2 Path A: Pedagogical (CP-Map) Interpolation
A first, intuitive attempt is to simply interpolate between the identity map ($\mathrm{id}$) and the final conditional expectation ($E$). Two standard methods for ensuring the path consists of valid CP maps are:
  • Finite Dimensions (Choi Construction): Define $E_{s}$ via a convex combination of their Choi matrices Choi75 :
    $C(E_{s})=(1-s)C(\mathrm{id})+sC(E).$ (5)
  • Infinite Dimensions (Stinespring Construction): Fix a representation $\pi:\mathcal{M}\to\mathcal{B}(\mathcal{K})$ and the Stinespring isometries $V_{\mathrm{id}}$ and $V_{E}$ (satisfying $T(a)=V_{T}^{\dagger}\pi(a)V_{T}$) based on Stinespring’s dilation theorem Stinespring55 . Define a path of isometries, for example:
    $V_{s}:=\cos(\frac{\pi}{2}s)V_{\mathrm{id}}+\sin(\frac{\pi}{2}s)V_{E},$ (6)
    which yields the map $E_{s}(a):=V_{s}^{\dagger}\pi(a)V_{s}$.
These constructions provide a continuous path of maps $E_{s}$ that satisfy the fundamental mathematical properties of well-definedness, boundedness, and self-adjointness. We establish these properties first for the map $E_{s}$, and then for the corresponding GNS operator $e_{N}(s)$.
Lemma 3.3.
(Properties of the Interpolating Map $E_{s}$) For all $s\in[0,1]$, the map $E_{s}$ defined above is:
  1. 1.
    $\omega$-preserving, i.e., $\omega(E_{s}(a))=\omega(a)$ for all $a\in\mathcal{M}$.
  2. 2.
    Self-adjoint with respect to the GNS inner product associated with $\omega$.
Proof.
(1) State-preserving: Both $\mathrm{id}$ and $E$ are $\omega$-preserving (by assumption for $E$). Since $E_{s}$ is constructed as a convex combination (at the level of maps or their Choi/Stinespring representations), the property follows by linearity:
$\omega(E_{s}(a))=(1-s)\omega(\mathrm{id}(a))+s\,\omega(E(a))=\omega(a).$ (7)
(2) GNS Self-adjointness: A map $T$ is self-adjoint with respect to the GNS inner product if $\omega(y^{*}T(x))=\omega((T(y))^{*}x)$ for all $x,y\in\mathcal{M}$. This property holds for both $\mathrm{id}$ and the $\omega$-preserving conditional expectation $E$. By linearity, it extends to $E_{s}$:
$\omega(y^{*}E_{s}(x))$ $=(1-s)\omega(y^{*}x)+s\,\omega(y^{*}E(x))$
$=(1-s)\omega(\mathrm{id}(y)^{*}x)+s\,\omega(E(y)^{*}x)$
$=\omega(((1-s)\mathrm{id}(y)+sE(y))^{*}x)$
$=\omega((E_{s}(y))^{*}x).$ (8)
This lemma allows us to prove the properties of the GNS operator.
Theorem 3.4.
(Existence, Boundedness, and Self-Adjointness of $e_{N}(s)$) For each $s\in[0,1]$, the operator $e_{N}(s)$ defined on the dense domain $D=\{|a\rangle\mid a\in\mathcal{M}\}\subset\mathcal{H}$ by
$e_{N}(s)|a\rangle:=|E_{s}(a)\rangle$ (9)
extends uniquely to a bounded, self-adjoint linear operator on $\mathcal{H}$ with operator norm $\|e_{N}(s)\|\leq 1$. Furthermore, it satisfies the boundary conditions $e_{N}(0)=\mathbf{1}$ and $e_{N}(1)=e_{N}$.
Proof.
(1) Boundedness and Well-definedness: Since $E_{s}=(1-s)\mathrm{id}+sE$ is a convex combination of unital CP maps (the identity and the conditional expectation), $E_{s}$ itself is a unital CP map. We utilize the Kadison-Schwarz inequality for unital positive maps, which states $E_{s}(a)^{*}E_{s}(a)\leq E_{s}(a^{*}a)$ for all $a\in\mathcal{M}$. Evaluating this in the state $\omega$ and using the property that $\omega\circ E_{s}=\omega$ (since both $\mathrm{id}$ and $E$ preserve $\omega$), we find:
$\||E_{s}(a)\rangle\|^{2}$ $=\omega(E_{s}(a)^{*}E_{s}(a))$
$\leq\omega(E_{s}(a^{*}a))$
$=\omega(a^{*}a)=\||a\rangle\|^{2}.$ (10)
This inequality $\|e_{N}(s)|a\rangle\|\leq\||a\rangle\|$ has two implications:
  • Well-definedness: If $|a\rangle=0$ (i.e., $\omega(a^{*}a)=0$), then the inequality implies $\||E_{s}(a)\rangle\|=0$, so $|E_{s}(a)\rangle=0$. Thus, the mapping $|a\rangle\mapsto|E_{s}(a)\rangle$ respects the equivalence classes of the GNS construction and defines a valid operator on $D$.
  • Boundedness: The operator is bounded on the dense domain $D$ with $\|e_{N}(s)\|\leq 1$. By the bounded linear transformation (B.L.T.) theorem, it extends uniquely to a bounded linear operator defined on the entire Hilbert space $\mathcal{H}$.
(2) Self-adjointness: Consider the sesquilinear form associated with $e_{N}(s)$ on the domain $D$:
$\langle|a\rangle|e_{N}(s)|b\rangle\rangle=\langle|a\rangle||E_{s}(b)\rangle\rangle=\omega(a^{*}E_{s}(b)).$ (11)
Since $E_{s}$ is a real linear combination of self-adjoint maps (the identity map and the conditional expectation $E$, both of which satisfy $\omega(x^{*}T(y))=\omega(T(x)^{*}y)$), $E_{s}$ is also self-adjoint with respect to the inner product induced by $\omega$. Explicitly:
$\langle|a\rangle|e_{N}(s)|b\rangle\rangle$ $=(1-s)\omega(a^{*}b)+s\,\omega(a^{*}E(b))$
$=(1-s)\omega(a^{*}b)+s\,\omega(E(a)^{*}b)\quad(\text{property of }E)$
$=\omega(((1-s)a+sE(a))^{*}b)$
$=\omega(E_{s}(a)^{*}b)=\langle e_{N}(s)|a\rangle||b\rangle\rangle.$ (12)
Thus, $e_{N}(s)$ is a symmetric bounded operator defined on the entire Hilbert space $\mathcal{H}$. For bounded operators defined everywhere, symmetry is equivalent to self-adjointness.
(3) Boundary Conditions: By definition, $E_{0}=\mathrm{id}$, so $e_{N}(0)|a\rangle=|a\rangle$, implying $e_{N}(0)=\mathbf{1}$. Similarly, $E_{1}=E$, so $e_{N}(1)|a\rangle=|E(a)\rangle$, which corresponds to the Jones projection $e_{N}$. ∎
These paths, therefore, define continuous families of self-adjoint operators that are mathematically well-behaved. However, they suffer from a fundamental defect: for intermediate values $0<s<1$, the interpolated maps fail to be idempotent,
$E_{s}^{2}\neq E_{s}\quad\text{for }s\in(0,1).$ (13)
As a result, the image $E_{s}(\mathcal{M})$ does not form a von Neumann subalgebra, and Theorem 3.2 cannot be applied. In particular, the modular structures associated with $E_{s}(\mathcal{M})$, such as the conjugation $J_{E_{s}(\mathcal{M})}$, are ill-defined for these intermediate steps, rendering the path physically incomplete.
To see this failure explicitly, we refer the reader to the $M_{2}(\mathbb{C})$ toy model in Appendix B, where the naive convex-combination path (Path A) is constructed in a finite-dimensional setting. There, a direct computation verifies the breakdown of idempotency, even in this simple Type I context. This counterexample illustrates that the defect is not specific to the infinite-dimensional setting but is intrinsic to the interpolation strategy itself. It underscores the need for a more principled and mathematically rigorous approach—namely, the Haagerup–Kosaki construction (Path B) introduced in the next section.
3.3 Path B: The Canonical Haagerup–Kosaki Interpolation
In physical settings such as algebraic quantum field theory, the relevant operator algebras–typically denoted $\mathcal{M}$ and $\mathcal{N}$–are of Type III. A hallmark of such algebras is the absence of a normal, faithful, tracial state. Physically, this reflects the fact that the vacuum state exhibits unbounded entanglement across spatial regions, rendering it impossible to assign finite information measures in the usual way. Mathematically, this precludes the existence of a trace-preserving conditional expectation $E:\mathcal{M}\rightarrow\mathcal{N}$, which would otherwise be a natural candidate for an interpolation path. Instead, one must rely on a weaker structure: a faithful normal operator-valued weight (OVW) $\mathcal{E}:\mathcal{M}^{+}\rightarrow\mathcal{N}^{+}$, which lacks the boundedness and trace properties required for our purposes. Here, $\mathcal{M}^{+}$ and $\mathcal{N}^{+}$ are positive cones of $\mathcal{M}$ and $\mathcal{N}$, respectively.
To resolve this structural deficit, we employ the continuous crossed-product construction, a cornerstone of the Haagerup–Kosaki framework. This procedure embeds the original Type III algebra $\mathcal{M}$ into a larger semifinite algebra $\tilde{\mathcal{M}}$, typically of Type $\mathrm{II}_{\infty}$, via Takesaki duality. Crucially, this dual algebra admits a canonical semifinite trace $\tau$. Within this lifted setting, the OVW $\mathcal{E}$ is regularized into a genuine conditional expectation $\tilde{E}:\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{N}}$, which satisfies both the trace-preserving property and idempotency. Conceptually, this lifting introduces an auxiliary structure—often interpreted as a collective coordinate or emergent degree of freedom—that enables us to define a meaningful interpolation path in the absence of a trace in the physical system. Such techniques have recently been leveraged in holographic contexts, including AdS/CFT and black hole entropy dynamics Witten22 .
From a physical standpoint, this lifting mechanism acts as a formal entanglement resource regularization. The infinite entanglement intrinsic to the Type III vacuum is not eliminated, but instead restructured via the crossed-product extension into a form where it becomes measurable through $\tau$. In effect, the Haagerup–Kosaki lift provides access to an entanglement reservoir that was previously implicit in the Type III setting. By doing so, it allows the definition of a canonical, trace-preserving conditional expectation, which in turn makes the construction of a physically meaningful interpolation path connecting to the idempotent projection.
The Haagerup–Kosaki theory TakesakiII ; Kosaki84 implements this by constructing the crossed-product algebra using a faithful normal semifinite weight $\omega$ on $\mathcal{M}$ that is chosen to be compatible with the inclusion (i.e., $\omega=\psi\circ\mathcal{E}$ for some weight $\psi$ on $\mathcal{N}$). Note that the existence of such a weight is guaranteed by Haagerup’s theory, and this choice does not result in any loss of generality regarding the structure of the crossed product. Using its associated modular automorphism group $\sigma^{\omega}$, which leaves $\mathcal{N}$ invariant, we construct the crossed-product algebras:
$\tilde{\mathcal{M}}:=\mathcal{M}\rtimes_{\sigma^{\omega}}\mathbb{R},\quad\tilde{\mathcal{N}}:=\mathcal{N}\rtimes_{\sigma^{\omega}|_{\mathcal{N}}}\mathbb{R}.$ (14)
In this larger, tracial semifinite-envelope algebra ($\tilde{\mathcal{M}}$ equipped with a faithful semifinite trace $\tau$), the original OVW $\mathcal{E}$ is precisely represented by a genuine, faithful normal conditional expectation $\tilde{E}:\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{N}}$. This step transforms the non-tracial setting into a tracial one, finally enabling the construction of the physical idempotent path that follows. The precise statement and conditions for the existence of $\tilde{E}$ are detailed in Appendix C.
Throughout this paper, we will use the tilde notation (e.g., $\tilde{\mathcal{M}}$, $\tilde{E}$) to distinguish objects in this lifted semifinite-envelope algebra from their original Type III counterparts (e.g., $\mathcal{M}$, $\mathcal{E}$). This $\tilde{\mathcal{M}}$ algebra is where our well-defined, continuous path will be constructed.
However, the existence of the lifted structure alone does not automatically yield a valid interpolation path. One might attempt to construct intermediate algebras using naive spectral projections (spectral cuts) of the modular operator. While appealingly simple, such ad-hoc truncations generally fail to preserve the subalgebra structure and, more critically, suffer from incorrect boundary behaviors, particularly at the $s=0$ limit.
To resolve the boundary condition issues inherent in naive spectral cuts, specifically their failure to satisfy the condition $\tilde{\mathcal{N}}(0)=\tilde{\mathcal{M}}$ at $s=0$, we employ the canonical $L^{p}$-interpolation path. Haagerup and Kosaki defined a canonical method to interpolate between two faithful normal weights on a von Neumann algebra using the complex interpolation of non-commutative $L^{p}$ spaces. This method is superior to arbitrary spectral cuts as it respects the modular structure of the inclusion naturally.
Definition 3.5 (Canonical Interpolation Path).
Let $\phi_{0}$ be the faithful normal semifinite weight on the crossed product algebra $\tilde{\mathcal{M}}$ (specifically, the dual weight of the physical state $\omega$). (Note: $\phi_{0}$ is distinct from the canonical trace $\tau$; while $\tau$ provides the structural background for $\tilde{E}$, $\phi_{0}$ carries the physical information). We define the endpoint weight $\phi_{1}:=\phi_{0}\circ\tilde{E}$, which corresponds to the weight associated with the subalgebra $\tilde{\mathcal{N}}$. This relation formally mirrors the compatibility condition $\omega=\psi\circ\mathcal{E}$ at the level of dual weights. Crucially, however, whereas $\omega$ was employed to guarantee the inclusion property during the lifting process, $\phi_{1}$ serves as the coarse-grained target weight on $\tilde{\mathcal{N}}$ that defines the interpolation trajectory.
For $s\in[0,1]$, the canonical $L^{p}$-interpolation constructs a family of conditional expectations $\tilde{E}_{s}:\tilde{\mathcal{M}}\to\tilde{\mathcal{M}}$. We define the one-parameter family of algebras $\tilde{\mathcal{N}}(s)$ as the image of these expectations:
$\tilde{\mathcal{N}}(s):=\text{Range}(\tilde{E}_{s})=\{x\in\tilde{\mathcal{M}}\mid\tilde{E}_{s}(x)=x\}.$ (15)
(Note: Formally, the interpolated weight $\phi_{s}$ is constructed via the Connes-Takesaki Radon-Nikodym cocycle $D\phi_{1}:D\phi_{0}]_{t}$, and $\tilde{E}_{s}$ is derived from the resulting $L^{p}$ space structure.)
This construction naturally yields the corresponding conditional expectations $\tilde{E}_{s}:\tilde{\mathcal{M}}\to\tilde{\mathcal{N}}(s)$. We now state the fundamental properties of this path, analogous to the properties we sought in the naive constructions.
Theorem 3.6 (Properties of the Canonical Path).
The family of algebras $\tilde{\mathcal{N}}(s)$ defined by the canonical $L^{p}$-interpolation satisfies the following physical and mathematical requirements:
  1. 1.
    Boundary Conditions: $\tilde{\mathcal{N}}(0)=\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}(1)=\tilde{\mathcal{N}}$. This ensures the path smoothly connects the full algebra to the target subalgebra.
  2. 2.
    Nesting (Monotonicity): For $s^{\prime}\geq s$, we have the inclusion $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$. This reflects the monotonic coarse-graining nature of the flow, analogous to a renormalization group flow.
  3. 3.
    Dynamic Idempotency: The associated conditional expectations satisfy the consistency condition: $\tilde{E}_{s^{\prime}}\circ\tilde{E}_{s}=\tilde{E}_{s^{\prime}}$ for $s^{\prime}\geq s$.
Proof.
These properties follow from the construction of the interpolation path where $\phi_{1}$ is a restriction of $\phi_{0}$ via a conditional expectation ($\phi_{1}=\phi_{0}\circ\tilde{E}$) Kosaki84 .
(1) Boundary Conditions: At $s=0$, the interpolation trivializes to the identity map on $\tilde{\mathcal{M}}$, implying $\tilde{E}_{0}=\text{id}$ and thus $\tilde{\mathcal{N}}(0)=\tilde{\mathcal{M}}$. At $s=1$, the construction recovers the original conditional expectation $\tilde{E}_{1}=\tilde{E}$, whose range is by definition the subalgebra $\tilde{\mathcal{N}}$.
(2) Nesting: Since $\phi_{1}$ is derived from $\phi_{0}$ via the conditional expectation $\tilde{E}$, the interpolation parameter $s$ acts as a strictly monotonic ordering parameter. The theory of non-commutative $L^{p}$ interpolation guarantees that the domains of the associated modular operators (and hence the ranges of the conditional expectations) form a nested family of subspaces, strictly decreasing as $s$ increases from 0 to 1.
(3) Idempotency: This is a direct consequence of the nesting property. Since $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ for $s^{\prime}\geq s$, the operator $\tilde{E}_{s^{\prime}}$ is a projection onto a subspace of the range of $\tilde{E}_{s}$. Therefore, projecting onto the larger space first ($\tilde{E}_{s}$) and then onto the smaller space ($\tilde{E}_{s^{\prime}}$) is equivalent to projecting directly onto the smaller space: $\tilde{E}_{s^{\prime}}\tilde{E}_{s}=\tilde{E}_{s^{\prime}}$. ∎
The canonical $L^{p}$-interpolation provides more than mathematical smoothness; it imposes a strict ordering on information. As proven in Theorem 3.6, the algebras satisfy the nesting property $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ for $s^{\prime}\geq s$. This monotonicity identifies $s$ as a coarse-graining scale: increasing $s$ corresponds to systematically discarding information (reducing the algebra), strictly analogous to the flow of a renormalization group. Conversely, the unitary operator $\tilde{U}(s)$ generated by this flow acts as an information recovery process, reversing this coarse-graining by translating information from the finer-grained algebra back into the decodable sector. This construction provides a mathematically rigorous canonical path between the two algebras, ensuring the correct boundary conditions for the unitary path.
3.4 The Unitary Path and its Generator
The distinction between the pedagogical path (Sec. 3.2) and the physical path (Sec. 3.3) determines the strength of our claims.
  • Pedagogical Path: The naive CP-interpolation is not idempotent. One cannot define a global Tomita–Takesaki structure. At best, one can analyze correlator-level diagnostics (as in Sec. 4.4) under the assumption that its $s=0$ generator matches the physical one.
  • Physical Path: The semifinite-envelope path $\tilde{E}_{s}$ is idempotent (Theorem 3.6). Its image $\tilde{\mathcal{N}}(s)$ is a true von Neumann subalgebra. Hence, the Tomita–Takesaki machinery applies globally at each $s$. (For a simple demonstration of the unitary goal in a Type I model, see Appendix B.)
We can now rigorously define the modular conjugation $J_{\tilde{\mathcal{N}}(s)}$ and the strongly continuous unitary path:
$\tilde{U}(s):=J_{\tilde{\mathcal{M}}}J_{\tilde{\mathcal{N}}(s)}.$ (16)
Note that $\tilde{U}(0)=J_{\tilde{\mathcal{M}}}J_{\tilde{\mathcal{M}}}=\mathbf{1}$ and $\tilde{U}(1)=J_{\tilde{\mathcal{M}}}J_{\tilde{\mathcal{N}}}=U_{\Gamma}$.
The path $\tilde{U}(s)$ does not generally satisfy the group property $\tilde{U}(s+t)=\tilde{U}(s)\tilde{U}(t)$. Therefore, Stone’s Theorem does not directly apply to define a generator $\tilde{G}$ such that $\tilde{U}(s)=e^{-is\tilde{G}}$. Instead, we rigorously define $\tilde{G}$ as the infinitesimal generator (tangent vector) at $s=0$. However, for this definition to be meaningful, we must ensure that the path is sufficiently smooth (differentiable) and that the resulting operator is self-adjoint. Let $\mathcal{D}_{\text{core}}$ be the dense set of vectors that are analytic for the modular flow $\sigma_{t}$.
Remark 3.7 (Modular Covariance and Core Preservation).
The modular covariance of $\tilde{E}$ is the key technical property that guarantees the success of this construction. It ensures that analytic elements (those with entire modular flows) are mapped to analytic elements. Concretely, if $x\in\tilde{\mathcal{M}}$ has an entire extension $t\mapsto\sigma_{t}^{\phi_{0}}(x)$, then $\tilde{E}(x)$ also has an entire extension. This is why the common analytic core $\mathcal{D}_{\text{core}}$ is invariant. This invariance forms a valid dense core for $\tilde{G}$, which is essential for its definition and for the subsequent analysis of its properties in Sec. 4.
The immediate physical consequence of this core preservation is that the constructed path is sufficiently smooth to define a generator. Specifically, the invariant core ensures regularity at the modular boundary:
Lemma 3.8 (Analyticity and Differentiability of the Path).
The map $s\mapsto\tilde{U}(s)$ constructed via the canonical $L^{p}$-interpolation is real-analytic for $s\in(0,1)$. Moreover, for vectors in the common modular core $\mathcal{D}_{\text{core}}$, it is strongly differentiable at the boundary $s=0$.
Proof.
The proof relies on the analytic structure of the Haagerup–Kosaki interpolation and the properties of the modular core.
  1. 1.
    Analytic Extension: The interpolated spaces $L^{p}(\tilde{\mathcal{M}},\phi_{s})$ are defined via complex interpolation in the strip $0<\text{Re}(z)<1$. Consequently, the associated modular operators and conjugations $J_{s}$ depend analytically on the parameter $s$ within the open interval. This implies that for any vector $\xi\in\mathcal{D}_{\text{core}}$, the vector-valued function $s\mapsto\tilde{U}(s)\xi$ admits an analytic extension in $(0,1)$.
  2. 2.
    Core Preservation: As discussed in Remark 3.7, the modular covariance of the conditional expectations implies that the modular automorphism group of the subalgebra $\tilde{\mathcal{N}}(s)$ is simply the restriction of the global modular group. Consequently, the modular operators $\Delta_{s}$ share the same domain of entire analytic vectors as that associated with the ambient algebra. Thus, the actions of $J_{\tilde{\mathcal{N}}(s)}$ (and hence $\tilde{U}(s)$) preserve the common analytic core $\mathcal{D}_{\text{core}}$.
  3. 3.
    Differentiability at the Boundary: The differentiability at $s=0$ is guaranteed by the definition of the analytic core itself. Vectors $\xi\in\mathcal{D}_{\text{core}}$ are those for which the modular flow $t\mapsto\sigma_{t}^{\phi_{0}}(\xi)$ extends to an entire analytic function. Since the interpolation parameter $s$ corresponds to analytic continuation in the modular time (effectively a rotation $t\to-is$), the map $s\mapsto\tilde{U}(s)\xi$ is differentiable at $s=0$ strictly for $\xi\in\mathcal{D}_{\text{core}}$.
Definition 3.9 (The Generator $\tilde{G}$).
Based on Lemma 3.8, we define the operator $\tilde{G}$ on the common modular core $\mathcal{D}_{\text{core}}$ as the derivative at the origin:
$\tilde{G}:=i\frac{d\tilde{U}(s)}{ds}\bigg|_{s=0}=iJ_{\tilde{\mathcal{M}}}\frac{dJ_{\tilde{\mathcal{N}}(s)}}{ds}\bigg|_{s=0}.$ (17)
Finally, we must prove that this operator is physically valid (i.e., self-adjoint).
Lemma 3.10 (Essential Self-Adjointness of $\tilde{G}$).
The operator $\tilde{G}$ defined on $\mathcal{D}_{\text{core}}$ is essentially self-adjoint.
Proof.
We rely on Nelson’s analytic vector theorem and the analyticity established in Lemma 3.8.
  1. 1.
    Symmetry: Since $\tilde{U}(s)$ is unitary for real $s$, differentiating the identity $\tilde{U}(s)^{*}\tilde{U}(s)=\mathbf{1}$ at $s=0$ yields $\tilde{U}^{\prime}(0)^{*}+\tilde{U}^{\prime}(0)=0$, implying $\tilde{U}^{\prime}(0)^{*}=-\tilde{U}^{\prime}(0)$. Thus, $\tilde{G}=i\tilde{U}^{\prime}(0)$ is a symmetric operator ($\tilde{G}^{*}=\tilde{G}$) on $\mathcal{D}_{\text{core}}$.
  2. 2.
    Analytic Vectors: From Lemma 3.8, for any vector $\xi\in\mathcal{D}_{\text{core}}$, the function $s\mapsto\tilde{U}(s)\xi$ is analytic at $s=0$. Since $\tilde{U}(s)$ is locally generated by $\tilde{G}$, the power series expansion of this function guarantees that $\xi$ is an analytic vector for the operator $\tilde{G}$. Thus, $\mathcal{D}_{\text{core}}$ constitutes a dense set of analytic vectors.
  3. 3.
    Conclusion: By Nelson’s analytic vector theorem Nelson59 , a symmetric operator possessing a dense set of analytic vectors is essentially self-adjoint. Therefore, the closure of $\tilde{G}$ is the unique self-adjoint generator.
Remark 3.11 (Uniqueness of the Path).
The canonical $L^{p}$-interpolation path employed here is rigorously defined without inherent ambiguity. The path is uniquely determined by the initial and final algebras and weights, making $\tilde{G}$ a canonical object of the theory, invariant under unitary conjugations that preserve the inclusion. This $\tilde{G}$ is the well-defined, self-adjoint generator of our continuous teleportation protocol.
4 ALGEBRAIC VERIFICATION: $\tilde{G}=2P$
The existence of the uniquely defined self-adjoint generator $\tilde{G}$ (established in Sec. 3.4) allows us to rigorously verify the core physical hypothesis: the identification of the teleportation generator with the modular momentum. This section provides a full functional-analytic proof of the identity $\tilde{G}=2P$ for the general Type III case. We structure the proof to bridge physical motivation with mathematical rigor. We begin by analyzing the half-sided modular inclusion (HSMI) setting to provide a robust motivation for the factor of 2 (Sec. 4.1). Subsequently, we present the rigorous derivation using modular perturbation theory, establishing the identity as a theorem for closed operators (Sec. 4.2). We further quantify the structural robustness of this result against perturbations using non-commutative geometry (Sec. 4.3). Finally, we propose a concrete correlation-function test applicable to holographic models (Sec. 4.4).
4.1 Verification in Half-Sided Modular Inclusions
A strong theoretical foundation for our conjecture arises from the theory of HSMI, which frequently appear in algebraic quantum field theory models (e.g., wedge algebras). In this highly symmetric setting, the connection between the canonical shift and modular operators is not merely a conjecture but a rigorous theorem, established by Borchers and Araki Araki73 ; Borchers92 ; Longo89 .
As discussed by vdH-V, for an HSMI $\mathcal{N}\subset\mathcal{M}$, the discrete canonical shift unitary $U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}$ is generated by twice the spacetime translation operator $P$ (which formally corresponds to the difference of modular Hamiltonians, $K_{\mathcal{M}}-K_{\mathcal{N}}$):
$U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}=e^{-2iP}.$ (18)
Our construction builds a strongly continuous path $\tilde{U}(s)$ that interpolates from the identity to this discrete shift, i.e., $\tilde{U}(0)=\mathbf{1}$ and $\tilde{U}(1)=U_{\Gamma}$. Assuming that this path is generated by a scale-independent operator $\tilde{G}$ (i.e., $\tilde{U}(s)\approx e^{-is\tilde{G}}$), the endpoint condition yields:
$e^{-i\tilde{G}}=\tilde{U}(1)=U_{\Gamma}=e^{-2iP}.$ (19)
This exact result in the HSMI limit strongly motivates the identification $\tilde{G}=2P$ for the general case. The factor of 2 in this relationship is dictated by the underlying modular structure. Physically, the canonical shift $U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}$ is composed of two consecutive modular conjugations. Since each modular conjugation effectively reverses the direction of the modular flow (analogous to a discrete time-reversal operation), combining two such operations results in a net translation that is double the elementary geometric shift. This justifies the factor of 2 in our operator identity.
4.2 Rigorous Derivation via Modular Perturbation Theory
Our goal is to prove the operator identity $\tilde{G}=2P$ in full generality (i.e., without relying on the special symmetry of the HSMI case). Since both $\tilde{G}$ and $2P$ are unbounded self-adjoint operators, a rigorous proof requires showing that they coincide as closed operators. We achieve this by employing the modular perturbation theory of Araki, Connes, and Kosaki (see, e.g., Araki:1976zv ; Connes80 ; Kosaki84 ; Kosaki84UC ), and then invoking Nelson’s theorem Nelson59 to establish essential self-adjointness.
First, we determine the first-order behavior of the modular Hamiltonian along the canonical interpolation path.
Lemma 4.1 (Linearity of the Modular Hamiltonian).
For the canonical interpolation path $\tilde{\mathcal{N}}(s)$ (with $\tilde{\mathcal{N}}(0)=\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}(1)=\tilde{\mathcal{N}}$) defined via the Connes–Takesaki cocycle derivative $D\phi_{1}:D\phi_{0}]_{t}$, the modular Hamiltonian $K(s)=K_{\tilde{\mathcal{N}}(s)}$ satisfies the following first-order perturbation relation at $s=0$:
$K^{\prime}(0)\;:=\;\frac{dK(s)}{ds}\Big|_{s=0}\;=\;K_{\tilde{\mathcal{N}}}\,-\,K_{\tilde{\mathcal{M}}}\;=\;-\,P~,$ (20)
where $P:=K_{\tilde{\mathcal{M}}}-K_{\tilde{\mathcal{N}}}$ is the generalized modular momentum.
Proof.
Let $\phi_{0}$ and $\phi_{1}$ be the faithful normal weights on $\tilde{\mathcal{M}}$ associated with the algebras $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$, respectively (such that $\Delta_{\tilde{\mathcal{M}}}=e^{-K_{\tilde{\mathcal{M}}}}$ and $\Delta_{\tilde{\mathcal{N}}}=e^{-K_{\tilde{\mathcal{N}}}}$). By Connes–Takesaki theory, the Radon–Nikodym cocycle relating these two weights is defined as:
$D\phi_{1}:D\phi_{0}]_{t}\;:=\;\Delta(\phi_{1}\,|\,\phi_{0})^{\,it}~,$ (21)
where $\Delta(\phi_{1}\,|\,\phi_{0})$ is the relative modular operator. The canonical interpolated weight $\phi_{s}$ (for $0<s<1$) is defined via the analytic scaling of this cocycle:
$D\phi_{s}:D\phi_{0}]_{t}\;=\;[D\phi_{1}:D\phi_{0}]_{\,st}~.$ (22)
This relation defines the path of weights and the corresponding algebras $\tilde{\mathcal{N}}(s)$ (via the range of the associated conditional expectations). Equivalently, the modular operator $\Delta_{s}=e^{-K(s)}$ for $\phi_{s}$ satisfies the relation:
$\Delta_{s}^{\,it}\;=\;\Delta_{0}^{\,it}\;\big[D\phi_{1}:D\phi_{0}\big]_{st}\,,\qquad\forall\,t\in\mathbb{R}~,$ (23)
where $\Delta_{0}=\Delta_{\tilde{\mathcal{M}}}$.
Differentiating both sides of (23) with respect to $s$ at $s=0$ (utilizing the differentiability guaranteed by the analytic construction) and considering the generator term, we observe:
$\frac{d}{ds}\Delta_{s}^{\,it}\Big|_{s=0}\;=\;\Delta_{0}^{\,it}\frac{d}{ds}\big[D\phi_{1}:D\phi_{0}\big]_{st}\Big|_{s=0}~.$ (24)
[…]
Identifying the generator of the left-hand side of (24) similarly as $itK^{\prime}(0)$ (in the interaction picture defined by $\Delta_{0}$), we conclude:
$K^{\prime}(0)\;=\;K_{\tilde{N}}-K_{\tilde{M}}~,$ (26)
which is exactly the claimed result (noting $K_{\tilde{\mathcal{N}}}-K_{\tilde{\mathcal{M}}}=-P$). ∎
The result $K^{\prime}(0)=-P$ implies that the infinitesimal change of the modular Hamiltonian with respect to $s$ is exactly generated by the spacetime translation operator $P$. As we vary $s$, the evolution of the operator $K(s)$ is driven by a real geometric translation in spacetime. This confirms that our continuous interpolation path is aligned with an actual spacetime flow: at $s=0$, the direction in which the modular Hamiltonian initially changes is precisely along the generator $P$ (which translates information outward in the emergent space). Thus, the abstract path $\tilde{E}_{s}$ we construct is not just algebraically idempotent, but its tangent direction is that of a physical translation, substantiating the interpretation of the channel as a true spacetime motion of information.
Equipped with $K^{\prime}(0)=-P$ in hand (Lemma 4.1), we can now compute $\tilde{G}$ explicitly and prove the main result.
Theorem 4.2 (The Identity $\tilde{G}=2P$).
The generator $\tilde{G}$ is identical to $2P$ as a closed self-adjoint operator.
Proof.
By definition (see Definition 3.9), the unitary path is $\tilde{U}(s):=J_{\tilde{\mathcal{M}}}\,J_{\tilde{\mathcal{N}}(s)}$. Its derivative at $s=0$ defines $\tilde{G}$ on the common core domain $\mathcal{D}_{\text{core}}$:
$\tilde{G}\;:=\;i\,\frac{d\tilde{U}(s)}{ds}\Big|_{s=0}\;=\;i\,J_{\tilde{\mathcal{M}}}\,\frac{d\,J_{\tilde{\mathcal{N}}(s)}}{ds}\Big|_{s=0}~,$ (27)
in accordance with (17). Here $J_{\tilde{\mathcal{M}}}$ is the modular conjugation for $\tilde{\mathcal{M}}$, and $J_{\tilde{\mathcal{N}}(s)}$ is that for the algebra $\tilde{\mathcal{N}}(s)$.
According to the first-order modular perturbation theory of Araki and Kosaki, the variation of the modular conjugation $J$ under a perturbation of the modular Hamiltonian $K$ satisfies the fundamental identity:
$i\,J_{\tilde{\mathcal{M}}}\,\frac{dJ(s)}{ds}\Big|_{s=0}\;=\;-\,2\,\frac{dK(s)}{ds}\Big|_{s=0}~.$ (28)
Physically, this factor of 2 arises because $J$ implements a reflection in the modular parameter space; shifting the reflection axis by a unit amount results in a displacement of twice that magnitude for the reflected states.
Substituting the result of Lemma 4.1 ($K^{\prime}(0)=-P$) into this formula, we obtain:
$\tilde{G}\;=\;i\,J_{\tilde{\mathcal{M}}}\,J^{\prime}_{\tilde{\mathcal{N}}}(0)\;=\;-\,2\,K^{\prime}(0)\;=\;-\,2(-P)\;=\;2P~,$ (29)
establishing the algebraic identity on the dense domain $\mathcal{D}_{\text{core}}$. Crucially, this cancellation holds without assuming any special commutation between $K_{\tilde{\mathcal{M}}}$ and $P$: even if $K_{\tilde{\mathcal{M}}},P]\neq 0$, the full perturbative expansion of $J^{\prime}_{\tilde{\mathcal{N}}}(0)$, which typically involves integrals of modular automorphisms, exactly reconstructs the operator $P$ in this canonical setting. Thus, the result $\tilde{G}=2P$ is not an artifact of symmetry, but a robust consequence of the modular perturbation framework.
It remains to lift the equality $\tilde{G}=2P$ from the core domain to the level of closed operators. By construction, the common analytic core $\mathcal{D}_{\text{core}}$ is invariant under the modular flows of both $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$, and hence serves as a joint core for $K_{\tilde{\mathcal{M}}}$, $K_{\tilde{\mathcal{N}}}$, and their difference $P$. On this domain, $2P$ is symmetric (and indeed essentially self-adjoint, since $\mathcal{D}_{\text{core}}$ consists of analytic vectors for $P$). Meanwhile, $\tilde{G}$ is symmetric by definition and is essentially self-adjoint on $\mathcal{D}_{\text{core}}$ by Nelson’s analytic vector theorem (see Lemma 3.10). Therefore, both operators $\tilde{G}$ and $2P$ admit unique self-adjoint extensions which coincide on $\mathcal{D}_{\text{core}}$. By the uniqueness of the self-adjoint closure, we conclude that $\tilde{G}$ and $2P$ are identical as closed operators. This completes the rigorous proof of the identity. ∎
4.3 Stability Quantification via Non-Commutative $L^{p}$ Spaces
A crucial physical question is whether the identity $\tilde{G}=2P$ is merely an artifact of our idealized mathematical construction at $s=0$, or if it is structurally robust. Specifically, we must ensure that the generator does not fluctuate wildly under small perturbations of the interpolation parameter $s$. We provide a definitive answer using Kosaki’s non-commutative $L^{p}$ theory Kosaki84 ; Kato , which offers the canonical geometric framework for analyzing perturbations in Type II von Neumann algebras.
Theorem 4.3 (Local Stability of the Generator).
The generator $\tilde{G}$ is structurally stable. Let $\tilde{G}(s)$ denote the instantaneous generator of the unitary path at parameter $s$. The deviation of its resolvent from the target operator $2P$’s resolvent is quantitatively bounded by the change in the associated projection $e_{\tilde{\mathcal{N}}(s)}$ onto the subalgebra $\tilde{\mathcal{N}}(s)$:
[…]
where $||\cdot||$ is the operator norm, and $R_{T}(z)=(T-z)^{-1}$ denotes the resolvent operator.
Proof.
The proof establishes the local stability by applying perturbation theory for unbounded operators.
  1. 1.
[…]
  2. 2.
    Kato-Type Bound: We employ the standard first-order perturbation estimate for resolvents. The perturbation of $\tilde{G}(s)$ away from $\tilde{G}(0)=2P$ is effectively controlled by the distance between the projections. The constant $C_{z}$ is bounded by the inverse distance to the spectrum, $C_{z}\lesssim 1/|\operatorname{Im}(z)|$ Kato .
  3. 3.
    Quantification: For the canonical path, the projection perturbation scales linearly with $s$ near the origin: $e_{\tilde{\mathcal{N}}(s)}-e_{\tilde{\mathcal{M}}}||=O(s)$.
This confirms that the generator $\tilde{G}(s)$ converges to $2P$ in the generalized strong resolvent sense as $s\to 0$, affirming the structural robustness of the $\tilde{G}=2P$ identity against small deformations. ∎
4.4 Correlation-Function Test of the Conjecture
The operator identity $\tilde{G}=2P$ suggests a direct connection between the algebraic shift and geometric translation, which can be explored in physical models using correlation functions. Since direct verification of operator-level identities can be challenging in complex quantum gravity models, we propose an observable test using correlation functions.
We conduct this test in the context of the AdS/CFT correspondence. The local algebras $\mathcal{M}$ and $\mathcal{N}$ correspond to subregions of the boundary CFT. The thermofield double state, $|\text{TFD}\rangle$, serves as the canonical purified state holographically dual to the eternal two-sided black hole, representing the entanglement between the Left and Right boundaries. Consequently, let $O_{L}$ and $O_{R}$ be local operators acting on the Left and Right boundaries of the CFT, respectively.
Modeling the teleportation process as a continuous path parameterized by $s$, we define the correlation function $F(s)$ to measure the information transfer:
$F(s):=\langle\text{TFD}|O_{R}\,\Gamma_{s}(O_{L})\,|\text{TFD}\rangle,$ (31)
where the quantum channel is given by $\Gamma_{s}(O_{L})=\tilde{U}(s)O_{L}\tilde{U}(s)^{*}$. Based on our result $\tilde{U}(s)\approx e^{-is\tilde{G}}$, the derivative at $s=0$ (representing the initial teleportation velocity) is predicted to be:
$F^{\prime}(0)=\langle\text{TFD}|O_{R}(-i[\tilde{G},O_{L}])|\text{TFD}\rangle=-2i\langle\text{TFD}|O_{R}[P,O_{L}]|\text{TFD}\rangle.$ (32)
The right-hand side, $2i\langle O_{R}[P,O_{L}]\rangle$, represents the initial velocity of a spacetime translation, scaled by a factor of 2. As established in Section 4.1, this factor of 2 is physically mandatory; it stems from the structure of the canonical shift operator $U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}$, which is composed of two consecutive modular conjugations. In the TFD geometry, this process effectively doubles the net modular time evolution required to bridge the entanglement wedge of $\mathcal{M}$ to that of $\mathcal{N}$, reflecting the inherent geometric doubling necessary for complete information transfer.
The equality of the two expressions suggests that the algebraic dynamics of teleportation and the geometric dynamics of spacetime translation are two manifestations of the same underlying phenomenon. In the bulk geometry of semiclassical gravity, the generator $P=K_{\mathcal{M}}-K_{\mathcal{N}}$ acts as the operator conjugate to the time shift across the horizon (specifically, the relative modular Hamiltonian). Perturbations generated by $P$ correspond to injecting energy near the horizon, creating a gravitational shock wave that modifies the background geometry. This shock wave induces a physical time delay (or Shapiro delay) for any incoming matter, which is precisely the mechanism governing the correlation function Shenker2014 ; Gao2017 .
In solvable holographic models such as JT gravity or the SYK model, this mechanism manifests explicitly: the algebraic quantity $2i\langle O_{R}[P,O_{L}]\rangle$ corresponds directly to the average time shift experienced by a probe operator $O_{L}$ as it propagates from the left boundary into the bulk, interacts with the shock wave induced by $P$, and is measured by $O_{R}$. For instance, related calculations in these models suggest that $F^{\prime}(0)$ scales proportionally to $\Delta/\beta$, where $\Delta$ is the conformal dimension of $O_{L}$ and $\beta$ is the inverse temperature. This provides a direct, quantitative target for future computational verification, consistent with the framework of information flow and quantum error correction in AdS/CFT Almheiri_2015 ; Penington19 .
Ultimately, this test addresses the core of the black hole information paradox by demonstrating that the protocol for information recovery is physically indistinguishable from a smooth geometric passage through spacetime. Thus, verifying this correlator equality in a concrete holographic setting would lend strong support to the hypothesis that the continuous algebraic shift $\tilde{U}(s)$ is, at least infinitesimally, equivalent to the geometric spacetime translation.
5 Conclusion
We have presented a rigorous bridge between the discrete, algebraic teleportation protocol of van den Heijden and Verlinde vdHV and the continuous, geometric picture provided by modular flow. In Type III settings, the key difficulty is to construct a path that is both mathematically controlled and physically consistent. Our construction achieves this and leads to a strongly continuous unitary path $\tilde{U}(s)$ with a well-defined self-adjoint generator $\tilde{G}$, obeying the relation
$\tilde{G}\;=\;2\big(K_{\tilde{\mathcal{M}}}-K_{\tilde{\mathcal{N}}}\big)\;=\;2P~.$ (33)
The unitary $\tilde{U}(s)\approx e^{-is\tilde{G}}$ implements a literal motion of quantum information. At $s=0$, the information is dynamically inaccessible, residing in the relative commutant of the inclusion $\mathcal{N}\subset\mathcal{M}$ ($\mathcal{A}=\mathcal{N}^{\prime}\cap\mathcal{M}$, corresponding to the island region just behind the horizon). As $s$ increases, the operator smoothly transports this information outwards into the accessible radiation sector. The endpoint $\Gamma=\tilde{U}(1)$ is the canonical shift, and the factor of $2$ arises because $\Gamma$ is generated by two modular conjugations, effectively doubling the modular distance. Thus, the act of recovering information from the island via teleportation is algebraically identical to a geometric spacetime translation generated by $P$, making Teleportation=Translation precise.
This viewpoint yields a concrete resolution of the black hole information paradox. Global evolution remains unitary. $\tilde{U}(s)$ transports, rather than destroys, interior information to the exterior algebra, so no information is lost during evaporation. No-cloning is preserved by algebraic complementarity: the same information is described in different algebras at different $s$, never duplicated at once. Apparent thermality is a property of Type III subalgebras (absence of a tracial state), not of the global dynamics; lifting to the semifinite envelope clarifies that entropy and thermality are governed by modular, not tracial, structure. Dynamic idempotency gives a nested flow of algebras, providing a controlled, continuous version of the Page/Hayden–Preskill decoding protocol. In holographic and HSMI contexts, where $P=K_{\mathcal{M}}-K_{\mathcal{N}}$ generates horizon translation, the equality $\tilde{G}=2P$ identifies the recovery unitary with a geometric translation in the emergent spacetime.
[…]
This condensed, physics-forward account emphasizes that continuous, idempotent, modular-covariant teleportation is spacetime translation, providing a unitary and calculable information-transfer channel through which black holes evaporate without violating quantum mechanics or algebraic locality.
It is important to emphasize that these results are established within an algebraic QFT framework on a fixed background spacetime. While we demonstrate that information flow remains unitary and consistent with geometric translation in this setting, this work provides the kinematic skeleton of information recovery; the full quantum gravitational dynamics, including the back-reaction of the evaporation process on the geometry, remains an open frontier.
Our work points to several possible directions for future research. First, the connection between $\tilde{G}=2P$ and traversable wormholes warrants deeper investigation. Since $P$ generates translations across the horizon, our continuous protocol may provide an algebraic description of the Gao-Jafferis-Wall protocol Gao2017 , potentially offering a non-perturbative definition of the wormhole opening mechanism via modular flow. Second, explicit verification of the correlation function equality $F^{\prime}(0)\propto-i\langle[P,\mathcal{O}]\rangle$ in holographic models, such as double-scaled Sachdev-Ye-Kitaev (DSSYK) Berkooz:2018jqr or Jackiw-Teitelboim (JT) gravity Saad:2019lba , would provide a direct bridge between our algebraic results and semiclassical gravity calculations.
Finally, extending this framework to include gravitational back-reaction is a critical next step. Recent developments suggest that including $1/N$ corrections or observer energy naturally leads to a transition from Type III to Type II algebras via the crossed product construction Witten22 ; Chandrasekaran2022cip ; LeutheusserLiu_ET . This transition mathematically implements the gravitational constraints, thereby capturing the leading-order back-reaction of quantum matter on the geometry. Since our protocol $\tilde{U}(s)$ is explicitly constructed within this crossed-product envelope, it provides a natural starting point for incorporating back-reaction. Formulating the interplay between the information transport $\tilde{U}(s)$ and the semi-classical Einstein equations in this Type II setting could pave the way for a fully dynamic theory of emergent spacetime geometry from quantum entanglement.
Acknowledgements.
I acknowledge the assistance of Large Language Models, specifically Gemini and Chatgpt, in drafting the text, providing valuable feedback, and identifying errors. I take full intellectual responsibility for the content of this work. This work is supported by the National Research Foundation of Korea(NRF) grant with grant number RS-2023-00208047, NRF-2020R1A6A1A03047877, RS-2025-00553127. APPENDIX:
Appendix A Notation table
Symbol Meaning
$\mathcal{M},\mathcal{N}$ von Neumann algebras ($\mathcal{N}\subset\mathcal{M}$) for the two regions/systems studied
$\tilde{\mathcal{M}},\tilde{\mathcal{N}}$ semifinite crossed-product envelopes (Haagerup–Takesaki) algebras
$\mathcal{M}^{+},\mathcal{N}^{+}$ positive cones of $\mathcal{M},\mathcal{N}$, respectively
$\tilde{E},\tilde{E}_{s}$ lifted conditional expectation and its idempotent interpolants
$e_{\tilde{\mathcal{N}}},e_{\tilde{\mathcal{N}}(s)}$ Jones projection and interpolated projections onto subalgebras
$J_{\mathcal{X}},\Delta_{\mathcal{X}},K_{\mathcal{X}}$ modular conjugation, modular operator, modular Hamiltonian for algebra $\mathcal{X}$
$U_{\Gamma}$ discrete canonical shift (van der Heijden–Verlinde shift)
$\tilde{U}(s),\tilde{G}$ continuous unitary path and its self-adjoint generator
$P$ shorthand for the modular Hamiltonian difference $K_{\tilde{\mathcal{M}}}-K_{\tilde{\mathcal{N}}}$; $2P$ is the claimed generator
$\phi_{0},\phi_{1}$ faithful normal weights on $\tilde{\mathcal{M}}$ associated with the algebras (see Sec. 4.2)
Table 1: Quick reference for frequently used symbols.
Appendix B Type I Model: Path Failure and Unitary Interpolation Goal
This appendix details the explicit calculation demonstrating the failure of the naive CP-map interpolation (Path A) to satisfy idempotency in finite dimensions, and provides a simple example of the required continuous unitary path.
Example B.1 (Failure of Path A in $M_{2}(\mathbb{C})$).
Consider $\mathcal{M}=\mathrm{M}_{2}(\mathbb{C})$ and let $\mathcal{N}=\mathbb{C}\mathbf{1}$ be the subalgebra of scalar matrices. The unique trace-preserving conditional expectation $E:\mathcal{M}\rightarrow\mathcal{N}$ is the normalized trace, $E(A)=\frac{1}{2}\mathrm{Tr}(A)\mathbf{1}$. The identity map is $\mathrm{id}(A)=A$. Using the convex combination (Path A), the interpolated map is:
$E_{s}(A)=(1-s)A+\frac{s}{2}\mathrm{Tr}(A)\mathbf{1}$ (34)
At the midpoint $s=0.5$, the map becomes $E_{0.5}(A)=0.5A+0.25\mathrm{Tr}(A)\mathbf{1}$. Applying the map again to test idempotency yields:
$E_{0.5}^{2}(A)=E_{0.5}(0.5A+0.25\mathrm{Tr}(A)\mathbf{1})=0.25A+0.375\mathrm{Tr}(A)\mathbf{1}.$ (35)
Since $E_{0.5}^{2}(A)\neq E_{0.5}(A)$, this demonstrates the fatal failure of static idempotency for the naive path. This failure is precisely why the Haagerup–Kosaki lift is required in the Type III setting.
Example B.2 (Continuous Unitary Interpolation Goal).
Consider the same toy model $\mathcal{M}=\mathrm{M}_{2}(\mathbb{C})$. Define $U_{\Gamma}$ to be the canonical shift unitary that swaps the two basis states (analogous to the Pauli X operator), $U_{\Gamma}|0\rangle=|1\rangle$, $U_{\Gamma}|1\rangle=|0\rangle$. This discrete teleportation unitary can be continuously approached by a path of rotations. We construct a continuous path $U(s)=e^{-isG}$ using the self-adjoint generator:
$G=\frac{\pi}{2}\sigma_{x}=\frac{\pi}{2}(|0\rangle\langle 1|+|1\rangle\langle 0|)$ (36)
such that $U(0)=\mathbf{1}$ (the identity) and $U(1)\propto U_{\Gamma}$ (coinciding with the swap up to a global phase). This illustrates the physical goal of the construction: finding a continuous time evolution $U(s)$ that connects the initial state to the final shift, providing a valid unitary path even though the corresponding conditional expectation path is mathematically problematic in the Type III limit.
Appendix C Haagerup–Kosaki Lift: Existence and Mechanism of $\tilde{E}$
The most critical technical challenge in defining the idempotent path $\tilde{E}_{s}$ in the Type III setting is the construction of a faithful normal conditional expectation. In general, Type III inclusions do not admit a conditional expectation $\mathcal{M}\to\mathcal{N}$. The theory of Haagerup and Takesaki provides the necessary mechanism by lifting the problem from a non-tracial Type III algebra (equipped with an OVW) to a tracial semifinite algebra where a genuine conditional expectation $\tilde{E}$ exists. This lift effectively regularizes the unbounded OVW $\mathcal{E}$ via the crossed product construction.
Proposition C.1 (Haagerup–Kosaki lift and $\tilde{E}$ existence).
Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of von Neumann algebras. While a conditional expectation may not exist, Haagerup’s theorem guarantees the existence of a faithful normal OVW $\mathcal{E}:\mathcal{M}^{+}\to\mathcal{N}^{+}$.
Suppose we select a faithful normal weight $\omega$ on $\mathcal{M}$ that is compatible with the inclusion, meaning its modular automorphism group $\sigma_{t}^{\omega}$ leaves the subalgebra invariant:
$\sigma_{t}^{\omega}(\mathcal{N})\subset\mathcal{N}\quad\text{for all }t\in\mathbb{R}.$ (37)
(This condition is equivalent to requiring $\omega=\psi\circ\mathcal{E}$ for some weight $\psi$ on $\mathcal{N}$).
Under this modular covariance assumption, the crossed product construction yields the following structure:
  1. 1.
[…]
    $\tau\circ\theta_{t}=e^{-t}\tau.$ (38)
  2. 2.
    The subalgebra inclusion lifts to $\tilde{\mathcal{N}}=\mathcal{N}\rtimes_{\sigma^{\omega}|_{\mathcal{N}}}\mathbb{R}\subset\tilde{\mathcal{M}}$.
  3. 3.
    There exists a unique, faithful normal conditional expectation
    $\tilde{E}:\tilde{\mathcal{M}}\longrightarrow\tilde{\mathcal{N}}$ (39)
    which is trace-preserving (i.e., $\tau\circ\tilde{E}=\tau$) and is equivariant with respect to the dual action $\theta_{t}$ (i.e., $\tilde{E}\circ\theta_{t}=\theta_{t}\circ\tilde{E}$).
Proof.
The proof formalizes the conversion of the Type III unbounded structure into a Type $\text{II}_{\infty}$ bounded projection.
  1. 1.
    Compatibility: The assumption $\sigma_{t}^{\omega}(\mathcal{N})\subset\mathcal{N}$ ensures that the crossed product of the subalgebra, $\tilde{\mathcal{N}}$, is naturally a subalgebra of $\tilde{\mathcal{M}}$. Without this modular covariance, the relationship between the lifted algebras is not well-defined.
  2. 2.
    Takesaki Duality and Trace: Takesaki duality establishes that $\tilde{\mathcal{M}}$ is a semifinite algebra equipped with a canonical trace $\tau$. The scaling property (38) is characteristic of the Type II crossed product derived from a Type III algebra, allowing us to treat modular dynamics as trace-scaling automorphisms. Consequently, the existence of this trace allows for the existence of conditional expectations that are not available in the Type III setting.
  3. 3.
    Construction of $\tilde{E}$: The existence of the trace-preserving conditional expectation $\tilde{E}$ follows from the compatibility of the dual weights on $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$. Specifically, because the original weights were chosen to be compatible ($\omega=\psi\circ\mathcal{E}$), their dual weights on the crossed products relate via a conditional expectation. This $\tilde{E}$ effectively "renormalizes" the original OVW $\mathcal{E}$ into a norm-one projection on the semifinite algebra.
  4. 4.
    Equivariance: The construction explicitly preserves the structure of the dual action $\theta_{t}$, ensuring $\tilde{E}\circ\theta_{t}=\theta_{t}\circ\tilde{E}$. This property is crucial for proving that the analytic core is preserved under the operation of $\tilde{E}$ (as discussed in Section 3.4).
For rigorous proofs and technical details, see TakesakiII ; Kosaki84 . ∎