### v3

# Teleportation=Translation: Continuous recovery of black hole information

Abstract
[…]
1 Introduction
The black hole information paradox exposes a fundamental tension between the apparent thermality of Hawking radiation, derived from semiclassical gravity, and the unitary evolution required by quantum theory Hawking1976 . Although Hawking’s original calculation suggests that a semiclassical black hole evaporates via thermal radiation in a manner violating unitarity, recent developments indicate that information recovery is possible even within the semiclassical framework, provided the paradox is analyzed with refined conceptual and technical tools Almheiri2020 ; Raju2022 ; Akil:2025bhip .
In particular, the AdS/CFT correspondence has provided a semiclassical framework utilizing replica wormholes and quantum extremal surfaces to derive the Page curve Almheiri2019 ; Almheiri2020Page ; Penington2022 . While these results strongly support the preservation of unitarity, they raise critical questions regarding the interpretation of the gravitational path integral as an ensemble average SaadShenkerStanford and its compatibility with spacetime factorization HarlowJafferis2020 ; Marolf2020 . To address these foundational issues, quantum information and operator-algebraic approaches have been increasingly employed to clarify the underlying assumptions behind these entropy computations Raju2022 ; Kibe2022 .
However, calculating the entropy curve does not, by itself, explain the dynamical mechanism of information retrieval. To focus on the retrieval process, Hayden and Preskill reformulated the problem as a quantum decoding task. They demonstrated that by modeling the black hole as a fast scrambler, any new information falling in can be retrieved rapidly, provided the observer has access to the early radiation HaydenPreskill . This recovery process is operationally analogous to quantum teleportation. Building on this operational viewpoint, van den Heijden and Verlinde (vdH-V) reformulated the problem entirely within the language of operator algebras vdHV . Mathematical foundations for such teleportation protocols have been developed in Conlon2022 . In their framework, the canonical shift $\Gamma$ plays the role of a teleportation operator, leading to the identification Teleportation=Translation.
Nonetheless, as acknowledged in vdHV , a direct realization of this protocol in the continuum limit faces fundamental obstructions. Local operator algebras in quantum field theory are Type III von Neumann factors, distinguished by the absence of a tracial state Connes80 ; Witten . This algebraic feature reflects the physical reality that vacuum entanglement is effectively unbounded. Consequently, the maximally entangled resources prerequisite for standard teleportation are mathematically ill-defined in this setting.
In this work, we construct a smooth unitary interpolation to connect the discrete canonical shift with continuous spacetime evolution. A central challenge in this construction is ensuring that this interpolation path is canonical rather than arbitrary. To resolve this in the Type III context, we employ the Haagerup–Kosaki framework Connes80 ; Kosaki84 and lift the inclusion to a Type II crossed product via Takesaki duality Connes80 ; TakesakiII . Crucially, the conditional expectation in this setting is uniquely determined by the invariance of the reference weights defining the $L^{p}$-interpolation Kosaki84 ; Takesaki72 . This rigidity guarantees that our unitary path is intrinsic to the modular structure of the system. Furthermore, by exploiting the analytic properties of modular flow, we ensure the path is sufficiently smooth to define a continuous generator $\widetilde{G}$ as the unique self-adjoint operator associated with this flow.
Building on this construction, we derive our main result by proving that this generator bears a direct physical identification with the modular momentum. Specifically, we establish an exact operator identity asserting that the generator of the teleportation flow equals twice the difference of the modular Hamiltonians ($\widetilde{G}=2P$). The factor of 2 is dictated by Borchers’ commutation relations Borchers92 , ensuring consistency with the scaling of modular flows. Our proof establishes this equality on a dense core of analytic modular vectors and demonstrates the equality of the closed operators via Nelson’s analytic vector theorem Nelson59 . Furthermore, we quantify the stability of this identification under small deformations using non-commutative $L^{p}$-theory Kosaki84 ; Kato . On a physical level, this result provides a precise operator-algebraic realization of the Teleportation=Translation proposal, demonstrating that information recovery manifests as a continuous geometric translation. Finally, to facilitate the verification of this operator identity in complex physical models where direct analysis is challenging, we propose a concrete observable test based on correlation functions.
The remainder of this paper is organized as follows. In Sec. 2, we summarize the necessary elements of Tomita–Takesaki theory and introduce the Jones projection together with the canonical shift $\Gamma$. In Sec. 3, we construct the continuous unitary interpolation by lifting the Type III inclusion to the crossed product via the Haagerup–Kosaki framework. Sec. 4 is devoted to the proof of our main result, identifying the generator of this flow with the modular momentum ($\widetilde{G}=2P$). We conclude with a discussion of the results in Sec. 5.
Technical details are deferred to the appendices. Appendix A collects the notation used throughout the paper. Appendix B illustrates the failure of naive interpolation in finite dimensions and provides a concrete continuous unitary path. Finally, Appendix C details the Haagerup-Takesaki construction of faithful conditional expectations in Type III settings via operator-valued weights.
2 The Discrete Protocol: Canonical Shift
2.1 Preliminaries: Tomita–Takesaki Theory
We work on a fixed Gelfand-Naimark-Segal (GNS) Hilbert space $\mathcal{H}$ built from a faithful normal state $\omega$ (with cyclic and separating vector $|\Omega\rangle$) for the von Neumann algebra $\mathcal{M}$ TakesakiI . Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of von Neumann algebras. For an algebra $\mathcal{X}$ (where $\mathcal{X}=\mathcal{M}$ or $\mathcal{N}$), the core objects of Tomita–Takesaki theory TakesakiII are:
  • The Tomita operator $S_{\mathcal{X}}$, defined on the dense domain $\mathcal{X}|\Omega\rangle$ by $S_{\mathcal{X}}x|\Omega\rangle=x^{*}|\Omega\rangle$ for $x\in\mathcal{X}$.
  • Its polar decomposition $S_{\mathcal{X}}=J_{\mathcal{X}}\Delta_{\mathcal{X}}^{1/2}$, where $J_{\mathcal{X}}$ is the anti-unitary modular conjugation and $\Delta_{\mathcal{X}}$ is the positive self-adjoint modular operator.
  • The modular Hamiltonian $K_{\mathcal{X}}$, defined by $\Delta_{\mathcal{X}}=e^{-2\pi K_{\mathcal{X}}}$.
  • The modular automorphism group $\sigma_{t}^{\mathcal{X}}(a)=\Delta_{\mathcal{X}}^{it}a\Delta_{\mathcal{X}}^{-it}$.
These objects satisfy $J_{\mathcal{X}}\mathcal{X}J_{\mathcal{X}}=\mathcal{X}^{\prime}$, where $\mathcal{X}^{\prime}$ is the commutant of $\mathcal{X}$.
2.2 The Canonical Shift Protocol
The discrete vdH-V protocol is built from three components vdHV .
Step 1: Conditional Expectation. Let $E:\mathcal{M}\to\mathcal{N}$ be a normal, faithful conditional expectation that is $\omega$-preserving: $\omega(E(a))=\omega(a)$ for all $a\in\mathcal{M}$. This map effectively discards information in $\mathcal{M}$ that is not in $\mathcal{N}$.
Step 2: Jones Projection. In the GNS representation where $|a\rangle:=a|\Omega\rangle$, $E$ is implemented by the Jones projection $e_{\mathcal{N}}$, an operator $e_{\mathcal{N}}\in\mathcal{B}(\mathcal{H})$ satisfying
$e_{\mathcal{N}}|a\rangle=|E(a)\rangle.$ (1)
$E$ being a $\omega$-preserving conditional expectation ensures that $e_{\mathcal{N}}$ is a well-defined self-adjoint projection ($e_{\mathcal{N}}^{2}=e_{\mathcal{N}}$, $e_{\mathcal{N}}^{\dagger}=e_{\mathcal{N}}$) and satisfies $e_{\mathcal{N}}ae_{\mathcal{N}}=E(a)e_{\mathcal{N}}$ for all $a\in\mathcal{M}$. The discrete (0 or 1) nature of this projection is the source of the protocol’s discreteness.
Step 3: Canonical Shift. The protocol is enacted by the unitary operator
$U_{\Gamma}:=J_{\mathcal{M}}J_{\mathcal{N}},$ (2)
which induces the automorphism (the shift) $\Gamma(a)=U_{\Gamma}aU_{\Gamma}^{\dagger}$. This map $\Gamma$ acts as a teleportation map, sending the relative commutant $\mathcal{N}^{\prime}\cap\mathcal{M}$ (information in $\mathcal{M}$ hidden from $\mathcal{N}$) to the relative commutant $\mathcal{M}^{\prime}\cap\mathcal{M}_{1}$ (information in the basic extension $\mathcal{M}_{1}:=\langle\mathcal{M},e_{\mathcal{N}}\rangle$ hidden from $\mathcal{M}$):
$\Gamma(\mathcal{N}^{\prime}\cap\mathcal{M})=\mathcal{M}^{\prime}\cap\mathcal{M}_{1}.$ (3)
Information is not lost, but relocated from one relative commutant to another. Crucially, while explicitly defined algebraically, this shift operator $U_{\Gamma}$ corresponds to a geometric translation in the emergent spacetime—a connection that will be precisely established in the continuum limit in the subsequent sections.
3 From a Discrete Protocol to a Continuous Path
Our core strategy is to soften the discrete Jones projection $e_{\mathcal{N}}$ by constructing a continuous family of maps $E_{s}$ and corresponding operators $e_{\mathcal{N}}(s)$ for $s\in[0,1]$, such that $E_{0}=\mathrm{id}$ and $E_{1}=E$. We first briefly discuss the challenge of (dynamical) idempotency in Sec. 3.1. Then, we present the pedagogical completely positive interpolation (Path A). Although this path ultimately fails the idempotency requirement in the infinite-dimensional (Type III) setting, discussing it is crucial for understanding the necessity of the more advanced construction (Sec. 3.2). We then lift the problem to the semifinite crossed-product (Haagerup–Kosaki envelope) and construct the canonical, physical interpolating path via $L^{p}$-space interpolation theory (Path B), satisfying the rigorous boundary conditions required for our proof (Sec. 3.3). Finally, using the Tomita–Takesaki machinery in the semifinite setting, we build the unitary path $\widetilde{U}(s)$ and analytically establish the generator $\widetilde{G}$ as a tangent vector, proving its self-adjointness via analytic vectors (Sec. 3.4).
3.1 The Physical Requirement: Dynamical Idempotency
To define a continuous unitary path $U(s)=J_{\mathcal{M}}J_{\mathcal{N}(s)}$, we must first construct a continuous path of von Neumann subalgebras $\mathcal{N}(s)$ such that $\mathcal{N}(0)=\mathcal{M}$ and $\mathcal{N}(1)=\mathcal{N}$. Such a path of algebras would, in turn, define a path of conditional expectations $E_{s}:\mathcal{M}\to\mathcal{N}(s)$. The crucial link between the map $E_{s}$ and the subalgebra $\mathcal{N}(s)$ is idempotency. Furthermore, for the parameter $s$ to be interpreted as a physical coarse-graining or information resolution flow, one should expect a condition even stronger than simple “static” idempotency ($E_{s}^{2}=E_{s}$). A true physical flow should satisfy a “dynamical” idempotency or semigroup property, such as $E_{s^{\prime}}\circ E_{s}=E_{s^{\prime}}$ for $s^{\prime}\geq s$. This condition, common in processes like the renormalization group, ensures the path is consistent: coarse-graining to level $s$ and then further to $s^{\prime}$ is identical to coarse-graining directly to $s^{\prime}$. While this dynamical property implies the static one, the static property ($E_{s}^{2}=E_{s}$) remains the minimum mathematical requirement for $E_{s}$ to be a conditional expectation at all, as established by the following theorem (Tomiyama theorem Tomiyama57 ).
Definition 3.1.
Dynamic idempotency for a family of maps $\{E_{s}\}_{s\in I}$ means:
$E_{s}^{2}=E_{s}\qquad\text{(static idempotency)},$
$E_{s^{\prime}}\circ E_{s}=E_{s^{\prime}}\quad\text{for }s^{\prime}\geq s\qquad\text{(semigroup/absorption property)}.$ (4)
While static idempotency is sufficient to define a projection at a single point, the dynamical condition is essential for interpreting the path as a coherent teleportation protocol. Physically, this ensures that the sequence of coarse-graining operations is consistent: projecting to an intermediate scale $s$ and then to a coarser scale $s^{\prime}$ must be identical to projecting directly to $s^{\prime}$. In the context of information recovery, this implies a strict nesting of information. Without this property, the intermediate stages of the protocol would lack a hierarchical structure, meaning that recovered information at step $s$ could be inexplicably lost or scrambled in subsequent steps. Thus, dynamical idempotency is the rigorous condition that prevents information leakage and ensures the monotonicity of the recovery flow from the black hole interior ($\mathcal{N}$) to the exterior ($\mathcal{M}$).
Theorem 3.2.
(Idempotent CP maps are conditional expectations) Let $\mathcal{M}$ be a von Neumann algebra and $\omega$ a faithful normal state on $\mathcal{M}$. Let $E:\mathcal{M}\to\mathcal{M}$ be a normal, unital, completely positive (CP) map satisfying $E^{2}=E$ and $\omega\circ E=\omega$. Then $\mathcal{N}:=E(\mathcal{M})$ is a von Neumann subalgebra of $\mathcal{M}$, and $E$ is the faithful normal conditional expectation from $\mathcal{M}$ onto $\mathcal{N}$.
This theorem sets our target: we must find a continuous path of idempotent CP maps.
3.2 Path A: Pedagogical (CP-Map) Interpolation
A first, intuitive attempt to connect the identity map ($\mathrm{id}$) to the target conditional expectation ($E$) is to construct a simple linear interpolation. The primary requirement is that every intermediate map must be a valid quantum channel, i.e., a unital CP map.
The most direct construction is a convex combination of the maps themselves:
$E_{s}=(1-s)\mathrm{id}+sE,\quad s\in[0,1].$ (5)
Since both $\mathrm{id}$ and $E$ are unital CP maps, their convex combination $E_{s}$ is automatically a unital CP map for all $s$, regardless of the underlying algebra type.
  • Finite Dimensions: This corresponds to linearly interpolating the Choi matrices Choi75 , preserving positivity throughout the path.
  • Infinite Dimensions: For general von Neumann algebras (including Type III), Stinespring’s dilation theorem Stinespring55 guarantees that any such CP map $E_{s}$ can be physically realized via an isometry $V_{s}$ on a larger Hilbert space (i.e., $E_{s}(x)=V_{s}^{\dagger}\pi(x)V_{s}$).
These constructions provide a continuous path of maps $E_{s}$ that satisfy the fundamental mathematical properties of well-definedness, boundedness, and self-adjointness. We establish these properties first for the map $E_{s}$, and then for the corresponding GNS operator $e_{\mathcal{N}}(s)$.
Lemma 3.3.
(Properties of the Interpolating Map $E_{s}$) For all $s\in[0,1]$, the map $E_{s}$ defined above is:
  1. 1.
    $\omega$-preserving, i.e., $\omega(E_{s}(a))=\omega(a)$ for all $a\in\mathcal{M}$.
  2. 2.
    Self-adjoint with respect to the GNS inner product associated with $\omega$.
Proof.
(1) State-preserving: Both $\mathrm{id}$ and $E$ are $\omega$-preserving (by assumption for $E$). Since $E_{s}$ is constructed as a convex combination (at the level of maps or their Choi/Stinespring representations), the property follows by linearity:
$\omega(E_{s}(a))=(1-s)\omega(\mathrm{id}(a))+s\,\omega(E(a))=\omega(a).$ (6)
(2) GNS Self-adjointness: A map $T$ is self-adjoint with respect to the GNS inner product if $\omega(y^{*}T(x))=\omega((T(y))^{*}x)$ for all $x,y\in\mathcal{M}$. This property holds for both $\mathrm{id}$ and the $\omega$-preserving conditional expectation $E$. By linearity, it extends to $E_{s}$:
$\omega(y^{*}E_{s}(x))$ $=(1-s)\omega(y^{*}x)+s\,\omega(y^{*}E(x))$
$=(1-s)\omega(\mathrm{id}(y)^{*}x)+s\,\omega(E(y)^{*}x)$
$=\omega(((1-s)\mathrm{id}(y)+sE(y))^{*}x)$
$=\omega((E_{s}(y))^{*}x).$ (7)
This lemma allows us to prove the properties of the GNS operator.
Theorem 3.4.
(Existence, Boundedness, and Self-Adjointness of $e_{\mathcal{N}}(s)$) For each $s\in[0,1]$, the operator $e_{\mathcal{N}}(s)$ defined on the dense domain $D=\{|a\rangle\mid a\in\mathcal{M}\}\subset\mathcal{H}$ by
$e_{\mathcal{N}}(s)|a\rangle:=|E_{s}(a)\rangle$ (8)
extends uniquely to a bounded, self-adjoint linear operator on $\mathcal{H}$ with operator norm $\|e_{\mathcal{N}}(s)\|\leq 1$. Furthermore, it satisfies the boundary conditions $e_{\mathcal{N}}(0)=\mathbf{1}$ and $e_{\mathcal{N}}(1)=e_{\mathcal{N}}$.
Proof.
(1) Boundedness and Well-definedness: Since $E_{s}=(1-s)\mathrm{id}+sE$ is a convex combination of unital CP maps (the identity and the conditional expectation), $E_{s}$ itself is a unital CP map. We utilize the Kadison-Schwarz inequality for unital positive maps, which states $E_{s}(a)^{*}E_{s}(a)\leq E_{s}(a^{*}a)$ for all $a\in\mathcal{M}$. Evaluating this in the state $\omega$ and using the property that $\omega\circ E_{s}=\omega$ (since both $\mathrm{id}$ and $E$ preserve $\omega$), we find:
$\||E_{s}(a)\rangle\|^{2}$ $=\omega(E_{s}(a)^{*}E_{s}(a))$
$\leq\omega(E_{s}(a^{*}a))$
$=\omega(a^{*}a)=\||a\rangle\|^{2}.$ (9)
This inequality $\|e_{\mathcal{N}}(s)|a\rangle\|\leq\||a\rangle\|$ has two implications:
  • Well-definedness: If $|a\rangle=0$ (i.e., $\omega(a^{*}a)=0$), then the inequality implies $\||E_{s}(a)\rangle\|=0$, so $|E_{s}(a)\rangle=0$. Thus, the mapping $|a\rangle\mapsto|E_{s}(a)\rangle$ respects the equivalence classes of the GNS construction and defines a valid operator on $D$.
  • Boundedness: The operator is bounded on the dense domain $D$ with $\|e_{\mathcal{N}}(s)\|\leq 1$. By the bounded linear transformation (B.L.T.) theorem, it extends uniquely to a bounded linear operator defined on the entire Hilbert space $\mathcal{H}$.
(2) Self-adjointness: Consider the sesquilinear form associated with $e_{\mathcal{N}}(s)$ on the domain $D$:
$\langle|a\rangle|e_{\mathcal{N}}(s)|b\rangle\rangle=\langle|a\rangle||E_{s}(b)\rangle\rangle=\omega(a^{*}E_{s}(b)).$ (10)
Since $E_{s}$ is a real linear combination of self-adjoint maps (the identity map and the conditional expectation $E$, both of which satisfy $\omega(x^{*}T(y))=\omega(T(x)^{*}y)$), $E_{s}$ is also self-adjoint with respect to the inner product induced by $\omega$. Explicitly:
$\langle|a\rangle|e_{\mathcal{N}}(s)|b\rangle\rangle$ $=(1-s)\omega(a^{*}b)+s\,\omega(a^{*}E(b))$
$=(1-s)\omega(a^{*}b)+s\,\omega(E(a)^{*}b)\quad(\text{property of }E)$
$=\omega(((1-s)a+sE(a))^{*}b)$
$=\omega(E_{s}(a)^{*}b)=\langle e_{\mathcal{N}}(s)|a\rangle||b\rangle\rangle.$ (11)
Thus, $e_{\mathcal{N}}(s)$ is a symmetric bounded operator defined on the entire Hilbert space $\mathcal{H}$. For bounded operators defined everywhere, symmetry is equivalent to self-adjointness.
(3) Boundary Conditions: By definition, $E_{0}=\mathrm{id}$, so $e_{\mathcal{N}}(0)|a\rangle=|a\rangle$, implying $e_{\mathcal{N}}(0)=\mathbf{1}$. Similarly, $E_{1}=E$, so $e_{\mathcal{N}}(1)|a\rangle=|E(a)\rangle$, which corresponds to the Jones projection $e_{\mathcal{N}}$. ∎
These paths, therefore, define continuous families of self-adjoint operators that are mathematically well-behaved. However, they suffer from a fundamental defect: for intermediate values $0<s<1$, the interpolated maps fail to be idempotent,
$E_{s}^{2}\neq E_{s}\quad\text{for }s\in(0,1).$ (12)
As a result, the image $E_{s}(\mathcal{M})$ does not form a von Neumann subalgebra, and Theorem 3.2 cannot be applied. In particular, the modular structures associated with $E_{s}(\mathcal{M})$, such as the conjugation $J_{E_{s}(\mathcal{M})}$, are ill-defined for these intermediate steps, rendering the path physically incomplete.
To see this failure explicitly, we refer the reader to the $M_{2}(\mathbb{C})$ toy model in Appendix B, where the naive convex-combination path (Path A) is constructed in a finite-dimensional setting. There, a direct computation verifies the breakdown of idempotency, even in this simple Type I context. This counterexample illustrates that the defect is not specific to the infinite-dimensional setting but is intrinsic to the interpolation strategy itself. It underscores the need for a more principled and mathematically rigorous approach—namely, the Haagerup–Kosaki construction (Path B) introduced in the next section.
3.3 Path B: The Canonical Haagerup–Kosaki Interpolation
Ideally, to resolve the idempotency failure identified in Path A, one would construct a continuous family of genuine conditional expectations. However, the existence of such trace-preserving expectations is obstructed by the intrinsic nature of the physical algebra. In physical settings such as algebraic quantum field theory, the relevant operator algebras–typically denoted $\mathcal{M}$ and $\mathcal{N}$–are of Type III. A hallmark of such algebras is the absence of a normal, faithful, tracial state. Physically, this reflects the fact that the vacuum state exhibits unbounded entanglement across spatial regions, rendering it impossible to assign finite information measures in the usual way. Mathematically, this precludes the existence of a trace-preserving conditional expectation $E:\mathcal{M}\rightarrow\mathcal{N}$. Instead, one must rely on a faithful normal operator-valued weight (OVW) $\mathcal{E}:\mathcal{M}^{+}\rightarrow\widehat{\mathcal{N}}^{+}$, which lacks the boundedness and trace properties required for our purposes. Here, $\mathcal{M}^{+}$ and $\mathcal{N}^{+}$ are the positive cones of the respective algebras, while $\widehat{\mathcal{N}}^{+}$ denotes the extended positive cone of $\mathcal{N}$, a domain that mathematically accommodates the infinite values (UV divergences) inherent in the Type III setting.
To resolve this structural deficit, we employ the continuous crossed-product construction, a cornerstone of the Haagerup–Kosaki framework. This procedure embeds the original Type III algebra $\mathcal{M}$ into a larger semifinite envelope $\tilde{\mathcal{M}}$ (typically of Type $\mathrm{II}_{\infty}$) via Takesaki duality. Crucially, this dual algebra admits a canonical semifinite trace $\tau$, which provides the necessary tracial background absent in the physical algebra. The connection between the envelope and the base algebra is formally established by the canonical structural map $\underline{\mathcal{E}}:\tilde{\mathcal{M}}^{+}\to\widehat$$\mathcal{M}$$^{+}$. As the fundamental operator-valued weight of the crossed-product, $\underline{\mathcal{E}}$ serves as a structural bridge that relates the canonical trace $\tau$ to the physical weights on $\mathcal{M}$.
Within this lifted setting, the original inclusion $\mathcal{N}\subset\mathcal{M}$ and its associated OVW $\mathcal{E}$ are regularized into a genuine, trace-preserving conditional expectation $\tilde{E}:\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{N}}$. While the structural weight $\underline{\mathcal{E}}$ ensures the mathematical consistency of the lifting itself, the conditional expectation $\tilde{E}$ provides the idempotent projection required to define a valid interpolation path. Following the insight of Kosaki’s theorem, the existence of such a trace-preserving projection allows the static inclusion to be dynamically resolved: it enables the construction of a continuous family of maps connecting the identity on $\tilde{\mathcal{M}}$ to the projection $\tilde{E}$ on $\tilde{\mathcal{N}}$. This mathematical mechanism transforms the algebraic inclusion into a continuous flow, providing the foundation for our information recovery protocol.
From a physical standpoint, this crossed-product construction serves as a formal regularization of entanglement resources. Conceptually, this construction introduces an auxiliary structure—often interpreted as a collective coordinate or an emergent degree of freedom—typically associated with the energy or the clock of an auxiliary observer Witten22 ; Chandrasekaran2022cip . This additional degree of freedom allows the infinite entanglement intrinsic to the Type III vacuum to be measured against the canonical trace $\tau$. In effect, the Haagerup–Kosaki lift accesses an entanglement reservoir implicit in the physical algebra, restructuring it into a form where a meaningful information flow can be defined. By doing so, it bridges the gap between the intractable Type III structure and the physically transparent Type II setting, where information recovery manifests as a smooth geometric process.
The Haagerup–Kosaki theory TakesakiII ; Kosaki84 implements this by constructing the crossed-product algebra using a faithful normal semifinite weight $\omega$ on $\mathcal{M}$ that is chosen to be compatible with the inclusion (i.e., $\omega=\psi\circ\mathcal{E}$ for some weight $\psi$ on $\mathcal{N}$). Note that the existence of such a weight is guaranteed by Haagerup’s theory, and this choice does not result in any loss of generality regarding the structure of the crossed product. Using its associated modular automorphism group $\sigma^{\omega}$, which leaves $\mathcal{N}$ invariant, we construct the crossed-product algebras:
$\tilde{\mathcal{M}}:=\mathcal{M}\rtimes_{\sigma^{\omega}}\mathbb{R},\quad\tilde{\mathcal{N}}:=\mathcal{N}\rtimes_{\sigma^{\omega}|_{\mathcal{N}}}\mathbb{R}.$ (13)
In this larger, tracial semifinite-envelope algebra ($\tilde{\mathcal{M}}$ equipped with a faithful semifinite trace $\tau$), the original OVW $\mathcal{E}$ is realized as a genuine, faithful normal conditional expectation $\tilde{E}:\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{N}}$. This transformation from a non-tracial to a tracial setting facilitates the subsequent construction of the physical idempotent path. The formal statement and the necessary conditions for the existence of $\tilde{E}$ are detailed in Appendix C.
Throughout this paper, we employ the tilde notation (e.g., $\tilde{\mathcal{M}}$, $\tilde{E}$) to distinguish objects in this lifted semifinite-envelope algebra from their original Type III counterparts (e.g., $\mathcal{M}$, $\mathcal{E}$). This algebra $\tilde{\mathcal{M}}$ serves as the workspace where our well-defined, continuous path is constructed.
However, the existence of this lifted structure does not, by itself, guarantee a valid interpolation path. One might naively attempt to construct intermediate algebras via spectral projections (or spectral cuts) of the modular operator. While intuitively appealing, such ad-hoc truncations generally fail to preserve the subalgebra structure and, more critically, exhibit pathological boundary behavior. Specifically, any finite spectral cut excludes the tails of the modular spectrum, leaving the algebra effectively open and unable to recover the identity in the $s\to 0$ limit. Consequently, no matter how the limit is taken, such paths cannot close the gap to the full algebra $\tilde{\mathcal{M}}$, resulting in a fundamental discontinuity at $s=0$.
To resolve these boundary issues—specifically to enforce the condition $\tilde{\mathcal{N}}(0)=\tilde{\mathcal{M}}$—we adopt the canonical $L^{p}$-interpolation path derived from the theory of non-commutative $L^{p}$ spaces Kosaki84 . Unlike arbitrary spectral cuts, this framework provides a rigorous method to interpolate between the algebra $\tilde{\mathcal{M}}$ (identified with $L^{\infty}$) and its predual space of normal functionals (identified with $L^{1}$).
Intuitively, the interpolation parameter $p\in[1,\infty]$ connects these spaces, where the intermediate elements behave like fractional powers of states and operators. In our specific context, we reparametrize this interpolation using $s=1/p\in[0,1]$. Crucially, by interpolating between the reference weight $\tilde{\phi}_{0}$ and its restriction $\tilde{\phi}_{1}=\tilde{\phi}_{0}\circ\tilde{E}$, we model the continuous loss of information. This transition follows the modular flow, ensuring that the non-commutative structure is preserved via analytic continuation, providing the unique rigidity required for our canonical path. Based on this structure, we define the interpolation path as follows.
Definition 3.5 (Canonical Interpolation Path).
Let $\tilde{\phi}_{0}$ be the faithful normal semifinite dual weight on the crossed product algebra $\tilde{\mathcal{M}}$, constructed from the physical weight $\omega$ on $\mathcal{M}$. We designate $\tilde{\phi}_{0}$ as the reference weight. [^1] Note that $\tilde{\phi}_{0}$ is distinct from the canonical trace $\tau$; while $\tau$ provides the structural background for the lifted setting, $\tilde{\phi}_{0}$ represents the physical reference weight.
We define the endpoint weight $\tilde{\phi}_{1}:=\tilde{\phi}_{0}\circ\tilde{E}$, which represents the weight restricted to the target subalgebra $\tilde{\mathcal{N}}$. This relation implies that $\tilde{\phi}_{1}$ is not an independent state, but rather a coarse-grained version of $\tilde{\phi}_{0}$ filtered through $\tilde{E}$. It also formally lifts the compatibility condition $\omega=\psi\circ\mathcal{E}$ to the level of dual weights on the crossed product. Crucially, whereas $\omega$ ensures the inclusion property during the lifting process, $\tilde{\phi}_{1}$ serves as the target weight on $\tilde{\mathcal{N}}$ that governs the interpolation trajectory.
For $s\in[0,1]$, the canonical $L^{p}$-interpolation constructs a one-parameter family of conditional expectations $\tilde{E}_{s}:\tilde{\mathcal{M}}\to\tilde{\mathcal{M}}$. The corresponding algebras $\tilde{\mathcal{N}}(s)$ are defined as the range of these expectations:
$\tilde{\mathcal{N}}(s):=\text{Range}(\tilde{E}_{s})=\{x\in\tilde{\mathcal{M}}\mid\tilde{E}_{s}(x)=x\}.$ (14)
The intermediate expectations $\tilde{E}_{s}$ are uniquely determined by the analytic continuation of the Connes–Takesaki Radon–Nikodym cocycle $D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}$. Specifically, if $h=d\tilde{\phi}_{1}/d\tilde{\phi}_{0}$ denotes the non-commutative Radon–Nikodym derivative, the interpolation is driven by the relation $D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}=h^{it}$. This relation underpins the analyticity of the Haagerup $L^{p}$-interpolation path at the boundaries (particularly at $s=0$) and ensures that the interpolated weights $\tilde{\phi}_{s}$ are uniquely determined by the resulting non-commutative $L^{p}$ space structure.
The construction in Definition 3.5 has a clear physical interpretation in terms of non-commutative geometry. The Radon–Nikodym derivative relation implies that the interpolated weight $\tilde{\phi}_{s}$ satisfies the exact power-law relation regarding its density:
$\frac{d\tilde{\phi}_{s}}{d\tilde{\phi}_{0}}\;=\;h^{s}\;=\;\left(\frac{d\tilde{\phi}_{1}}{d\tilde{\phi}_{0}}\right)^{s}.$ (15)
Heuristically, this signifies that $\tilde{\phi}_{s}$ acts as a non-commutative geometric mean between the reference weight $\tilde{\phi}_{0}$ and the coarse-grained weight $\tilde{\phi}_{1}$. Although formal products of weights are not defined in the operator algebra, we may conceptually visualize this interpolation as:
$\tilde{\phi}_{s}\;\sim\;\tilde{\phi}_{1}^{\,s}\,\tilde{\phi}_{0}^{\,1-s}~.$ (16)
This geometric averaging property ensures that the path $\tilde{\mathcal{N}}(s)$ follows the geodesic of the underlying modular structure, minimizing the information distance between the full algebra and the subalgebra.
This construction naturally yields the corresponding conditional expectations $\tilde{E}_{s}:\tilde{\mathcal{M}}\to\tilde{\mathcal{N}}(s)$. We now state the fundamental properties of this path, analogous to the properties we sought in the naive constructions.
Theorem 3.6 (Properties of the Canonical Path).
The family of algebras $\tilde{\mathcal{N}}(s)$ defined by the canonical $L^{p}$-interpolation satisfies the following physical and mathematical requirements:
  1. 1.
    Boundary Conditions: $\tilde{\mathcal{N}}(0)=\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}(1)=\tilde{\mathcal{N}}$. This ensures the path smoothly connects the full algebra to the target subalgebra.
  2. 2.
    Nesting (Monotonicity): For $s^{\prime}\geq s$, we have the inclusion $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$. This reflects the monotonic coarse-graining nature of the flow, analogous to a renormalization group flow.
  3. 3.
    Dynamic Idempotency: The associated conditional expectations satisfy the consistency condition: $\tilde{E}_{s^{\prime}}\circ\tilde{E}_{s}=\tilde{E}_{s^{\prime}}$ for $s^{\prime}\geq s$.
Proof.
These properties follow directly from the analytic structure of the Haagerup-Kosaki interpolation. Unlike naive spectral truncations, this path relies on complex interpolation between the weights $\tilde{\phi}_{0}$ and $\tilde{\phi}_{1}$, ensuring structural continuity via the modular automorphism group.
(1) Boundary Conditions: The interpolation relies on the analytic continuation of the Radon-Nikodym derivative $h$ defined above. At $s=0$, the interpolation yields the unperturbed reference weight $\tilde{\phi}_{0}$, implying that the associated projection is the identity map ($\tilde{E}_{0}=\text{id}_{\tilde{\mathcal{M}}}$). Conversely, at $s=1$, the construction converges in norm to the restricted weight $\tilde{\phi}_{1}$, thereby recovering the original conditional expectation $\tilde{E}_{1}=\tilde{E}$ and its range $\tilde{\mathcal{N}}$. This smooth limit behavior eliminates the discontinuity at $s=0$ that plagues spectral cut methods.
(2) Nesting: This property arises from the definition $\tilde{\phi}_{1}=\tilde{\phi}_{0}\circ\tilde{E}$. The operator $h$ essentially acts as a density operator representing the projection $\tilde{E}$. The interpolation corresponds to powers $h^{s}$, which serve as a soft filter whose intensity increases with $s$. Since $0\leq h\leq 1$ (in the appropriate operator sense), increasing $s$ strictly contracts the support of the associated expectations. Thus, for $s^{\prime}\geq s$, the range of $\tilde{E}_{s^{\prime}}$ is contained within the range of $\tilde{E}_{s}$.
(3) Idempotency: This follows directly from the nesting property. Since $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ for $s^{\prime}\geq s$, the operator $\tilde{E}_{s^{\prime}}$ projects onto a subspace contained within the range of $\tilde{E}_{s}$. Thus, applying the coarser filter $\tilde{E}_{s}$ before the finer filter $\tilde{E}_{s^{\prime}}$ leaves the latter invariant: $\tilde{E}_{s^{\prime}}\tilde{E}_{s}=\tilde{E}_{s^{\prime}}$. ∎
The canonical $L^{p}$-interpolation provides more than mathematical smoothness; it imposes a strict ordering on information. The nesting property $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ established in Theorem 3.6 identifies $s$ as a coarse-graining scale: increasing $s$ corresponds to systematically discarding information, strictly analogous to the flow of a renormalization group. Conversely, the unitary operator $\tilde{U}(s)$ generated by this flow acts as an information recovery process, reversing this coarse-graining by translating information from the finer-grained algebra back into the decodable sector. Thus, this construction not only connects the two algebras but also establishes the rigorous framework required for the unitary recovery path.
3.4 The Unitary Path and its Generator
The distinction between the pedagogical path (Sec. 3.2) and the physical path (Sec. 3.3) is not merely technical but conceptual. What is required is not just a continuous family of completely positive maps connecting the endpoints, but a one-parameter family of conditional expectations whose ranges remain von Neumann algebras. Only then do the modular objects exist along the entire path allowing for the rigorous definition and differentiation of the unitary flow that implements information recovery.
  • Pedagogical Path (Algebraic Limitation): The linear CP-interpolation $E_{s}=(1-s)\mathrm{id}+sE$ provides a mathematically well-defined path of maps. However, as illustrated in the $M_{2}(\mathbb{C})$ model (Appendix B), this linearity generally does not preserve the idempotency condition ($E_{s}^{2}\neq E_{s}$) for intermediate values of $s$. This property holds regardless of the algebra’s Type or the existence of a trace. As a result, the image of $E_{s}$ does not form a von Neumann subalgebra, which precludes the definition of the relative modular operator associated with the inclusion necessary to generate the unitary flow.
  • Physical Path (Algebraic Consistency): Conversely, the path $\tilde{E}_{s}$, defined by the canonical $L^{p}$-interpolation within the semifinite envelope, is idempotent by construction (Theorem 3.6). Since its image $\tilde{\mathcal{N}}(s)$ forms a von Neumann subalgebra for all $s\in[0,1]$, modular theory remains applicable throughout the interpolation. This structure ensures that the path represents a genuine flow of algebras rather than a simple interpolation of states, enabling the construction of the generating Hamiltonian even within the Type III setting.
The structural consistency of the physical path enables an explicit definition of the modular conjugation $J_{\tilde{\mathcal{N}}(s)}$ and the strongly continuous unitary path:
$\tilde{U}(s):=J_{\tilde{\mathcal{M}}}J_{\tilde{\mathcal{N}}(s)}.$ (17)
This specific construction represents the continuous counterpart to the operational realization of the discrete information recovery process discussed in vdHV and summarized in Sec. 2.2. Applying the logic established in Step 3 of the protocol, the modular conjugation acts as a teleportation map at each $s$, relocating the instantaneous relative commutant $\tilde{\mathcal{N}}(s)^{\prime}\cap\tilde{\mathcal{M}}$ to the corresponding basic extension’s commutant $\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$, where
$\tilde{\mathcal{M}}_{1}(s):=\langle\tilde{\mathcal{M}},e_{\tilde{\mathcal{N}}(s)}\rangle.$ (18)
Consequently, the flow generated by $\tilde{U}(s)$ explicitly implements this transport: it progressively decodes the information hidden in the relative commutant sector, mapping it back into the decodable sector, thereby effectively reversing the coarse-graining induced by the inclusion. Crucially, the analytic nature of the Haagerup–Kosaki interpolation guarantees the boundary conditions $\tilde{U}(0)=\mathbf{1}$ and $\tilde{U}(1)=U_{\tilde{\Gamma}}$, ensuring that the path continuously connects the identity to the target canonical shift.
The path $\tilde{U}(s)$ does not generally satisfy the group property $\tilde{U}(s+t)=\tilde{U}(s)\tilde{U}(t)$. Consequently, Stone’s Theorem does not directly apply to define a generator $\tilde{G}$ such that $\tilde{U}(s)=e^{-is\tilde{G}}$. Instead, we define $\tilde{G}$ as the infinitesimal generator (the tangent vector) at $s=0$. For this operator to be well-defined and physically meaningful, we must ensure that the path is strongly differentiable on a suitable domain and $\tilde{G}$ is essentially self-adjoint. To establish these properties, we consider a dense set of vectors $\mathcal{D}_{\text{core}}\subset\mathcal{H}$ that are analytic with respect to the modular flow $\sigma_{t}^{\tilde{\phi}_{0}}$. This choice of core allows us to control the domain of the generator and provides the necessary foundation for proving its self-adjointness in Lemma 3.10.
Remark 3.7 (Modular Covariance and Core Preservation).
The modular covariance of the conditional expectation $\tilde{E}_{s}$ is the pivotal property ensuring the structural consistency of this construction. Specifically, the fact that $\tilde{E}_{s}$ commutes with the modular flow $\sigma_{t}^{\tilde{\phi}_{0}}$ guarantees that analyticity is preserved under projection: analytic elements—those possessing entire modular continuations—are mapped to analytic elements within the subalgebra. Concretely, if $x\in\tilde{\mathcal{M}}$ admits an entire extension $t\mapsto\sigma_{t}^{\tilde{\phi}_{0}}(x)$, its image $\tilde{E}_{s}(x)$ likewise admits an entire extension. Consequently, the common analytic core $\mathcal{D}_{\text{core}}$ remains invariant under the interpolated expectations for all $s$. This invariance provides the stable, dense domain required to strictly define $\tilde{G}$ as a derivation, serves as the foundation for proving its self-adjointness in Lemma 3.10, and facilitates the broader analysis of its properties in Sec. 4.
The immediate physical consequence of this core preservation is that the constructed path is sufficiently smooth to define the generator $\tilde{G}$. Specifically, the invariant core $\mathcal{D}_{\text{core}}$ ensures regularity with respect to the modular flow, allowing for well-defined differentiation at $s=0$:
Lemma 3.8 (Analyticity and Differentiability of the Path).
The map $s\mapsto\tilde{U}(s)$ constructed via the canonical $L^{p}$-interpolation is real-analytic for $s\in(0,1)$. Moreover, for vectors in the common modular core $\mathcal{D}_{\text{core}}$, this map is strongly differentiable at the boundary $s=0$.
Proof.
The proof relies on the analytic structure of the Haagerup–Kosaki interpolation and the properties of the modular domain.
  1. 1.
    Analytic Extension inside the Interval: The interpolated spaces $L^{p}(\tilde{\mathcal{M}},\tilde{\phi}_{s})$ are defined via complex interpolation in the strip $0<\text{Re}(z)<1$, where $z$ extends the real parameter $s$ into the complex plane. Consequently, the structural maps, including the modular conjugations $J_{\tilde{\mathcal{N}}(s)}$, depend analytically on the parameter $s$ within the open interval. This implies that for any vector $\xi\in\mathcal{D}_{\text{core}}$, the vector-valued function $s\mapsto\tilde{U}(s)\xi$ is real-analytic in $(0,1)$.
  2. 2.
    Differentiability at the Boundary: The differentiability at $s=0$ hinges on the invariance of the analytic core established in Remark 3.7. Recall that vectors $\xi\in\mathcal{D}_{\text{core}}$ are characterized by the property that their modular flow $t\mapsto\sigma_{t}^{\tilde{\phi}_{0}}(\xi)$ extends to an entire analytic function. Since the interpolation parameter $s$ corresponds to the analytic continuation of the modular group to the imaginary axis (mathematically relating $s$ to $it$), the existence of the derivative at $s=0$ is equivalent to the existence of the analytic continuation. The invariance of $\mathcal{D}_{\text{core}}$ ensures that this analytic structure is preserved along the path, guaranteeing that the limit defining the derivative exists strictly for vectors within this core.
Definition 3.9 (The Generator $\tilde{G}$).
Motivated by the differentiability established in Lemma 3.8, we define the operator $\tilde{G}$ on the common modular core $\mathcal{D}_{\text{core}}$ as the tangent vector at the origin:
$\tilde{G}:=i\frac{d\tilde{U}(s)}{ds}\bigg|_{s=0}=iJ_{\tilde{\mathcal{M}}}\frac{dJ_{\tilde{\mathcal{N}}(s)}}{ds}\bigg|_{s=0}.$ (19)
Finally, we establish the physical validity of this operator.
Lemma 3.10 (Essential Self-Adjointness of $\tilde{G}$).
The operator $\tilde{G}$ defined on $\mathcal{D}_{\text{core}}$ is essentially self-adjoint.
Proof.
We invoke Nelson’s analytic vector theorem Nelson59 , utilizing the properties of the modular core $\mathcal{D}_{\text{core}}$.
  1. 1.
    Symmetry: Since $\tilde{U}(s)$ is unitary for real $s$, differentiating the identity $\tilde{U}(s)^{*}\tilde{U}(s)=\mathbf{1}$ at $s=0$ yields $\tilde{U}^{\prime}(0)^{*}\tilde{U}(0)+\tilde{U}(0)^{*}\tilde{U}^{\prime}(0)=0$. Using $\tilde{U}(0)=\mathbf{1}$, this implies $\tilde{U}^{\prime}(0)^{*}=-\tilde{U}^{\prime}(0)$. Thus, $\tilde{G}=i\tilde{U}^{\prime}(0)$ is a symmetric operator ($\tilde{G}^{*}=\tilde{G}$) on the dense domain $\mathcal{D}_{\text{core}}$.
  2. 2.
    Analytic Vectors: Recall that the generator $\tilde{G}$ arises from the analytic continuation of the Radon-Nikodym cocycle (effectively the logarithm of the relative modular operator $h$). The domain $\mathcal{D}_{\text{core}}$ is explicitly defined as the set of vectors that are analytic with respect to the modular flow generated by $\tilde{\phi}_{0}$ (and by extension, $h$). Mathematically, vectors that are analytic for a one-parameter group $h^{it}$ are analytic vectors for its generator $\log h$. Since $\tilde{G}$ is linearly related to this generator at $s=0$, the set $\mathcal{D}_{\text{core}}$ constitutes a dense set of analytic vectors for $\tilde{G}$.
  3. 3.
    Conclusion: By Nelson’s analytic vector theorem, a symmetric operator possessing a dense set of analytic vectors is essentially self-adjoint. Therefore, the closure of $\tilde{G}$ defines a unique self-adjoint quantum generator.
Remark 3.11 (Uniqueness of the Path).
The canonical $L^{p}$-interpolation path employed here is defined without ambiguity. It is uniquely determined by the initial reference weight and the target subalgebra, independent of arbitrary choices such as basis vectors or auxiliary cutoffs. Consequently, $\tilde{G}$ is a canonical object of the theory, invariant under unitary conjugations that preserve the inclusion structure. This establishes $\tilde{G}$ as the well-defined, self-adjoint generator of the continuous teleportation protocol.
4 DERIVATION OF THE OPERATOR IDENTITY $\tilde{G}=2P$
Having established the definition of the self-adjoint generator $\tilde{G}$ in Sec. 3.4, we now turn to the central objective of this work: establishing the identification of the teleportation generator with the modular momentum. This section presents a derivation of the identity $\tilde{G}=2P$ for general Type III algebras. We begin by analyzing the half-sided modular inclusion (HSMI) setting to motivate the specific form of the identity, particularly the factor of 2 (Sec. 4.1). Subsequently, we employ modular perturbation theory to derive the identity as an exact relation between closed operators (Sec. 4.2). We further examine the geometric stability of this result within the framework of non-commutative geometry (Sec. 4.3) and conclude by proposing a correlation function test applicable to holographic models (Sec. 4.4).
4.1 Motivation from Half-Sided Modular Inclusions
A strong theoretical foundation for the identification of the teleportation generator with modular momentum, as proposed by vdH-V vdHV , is found in highly symmetric settings, exemplified by HSMI. In this symmetric context, the connection between the canonical shift and modular operators is not merely a hypothesis but a rigorous theorem, established by Borchers and Wiesbrock Borchers92 ; Wiesbrock93 .
As discussed by vdH-V, for an HSMI $\mathcal{N}\subset\mathcal{M}$, the discrete canonical shift unitary $U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}$ is generated by twice the spacetime translation operator $P$ (which formally corresponds to the difference of modular Hamiltonians, $K_{\mathcal{M}}-K_{\mathcal{N}}$):
$U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}=e^{-2iP}.$ (20)
Our canonical interpolation constructs a strongly continuous path $\tilde{U}(s)$ that interpolates from the identity to this discrete shift, i.e., $\tilde{U}(0)=\mathbf{1}$ and $\tilde{U}(1)=U_{\tilde{\Gamma}}$. Assuming that this path is generated by a scale-independent operator $\tilde{G}$ (such that $\tilde{U}(s)=e^{-is\tilde{G}}$), the endpoint condition yields:
$e^{-i\tilde{G}}=\tilde{U}(1)=U_{\tilde{\Gamma}}=e^{-2iP}.$ (21)
This exact result in the HSMI limit strongly motivates the identification $\tilde{G}=2P$ for the general case. The factor of 2 in this relationship is dictated by the underlying modular structure. Physically, the canonical shift $U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}$ is composed of two consecutive modular conjugations. Since each modular conjugation acts as a geometric reflection (analogous to a CPT inversion), combining two such reflections results in a net translation that is exactly double the elementary geometric shift. This geometric logic justifies the factor of 2 in our operator identity.
4.2 Geometric Derivation via Canonical Path
Our objective is to establish the identity $\tilde{G}=2P$ in general Type III settings, moving beyond the specific symmetry of the HSMI case. Since both $\tilde{G}$ and $2P$ are unbounded self-adjoint operators, a precise proof requires demonstrating that they coincide as closed operators on a common core. We achieve this by employing the modular perturbation theory of Araki, Connes, and Kosaki (see, e.g., Araki:1976zv ; Connes80 ; Kosaki84 ; Kosaki84UC ) and invoking Nelson’s analytic vector theorem Nelson59 to ensure essential self-adjointness.
First, we determine the first-order behavior of the modular Hamiltonian along the canonical interpolation path.
Lemma 4.1 (Linearity of the Modular Hamiltonian).
For the canonical interpolation path $\tilde{\mathcal{N}}(s)$ defined via the Connes–Takesaki cocycle derivative $D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}$, the modular Hamiltonian $K(s)=K_{\tilde{\mathcal{N}}(s)}$ satisfies the following first-order perturbation relation at $s=0$:
$K^{\prime}(0)\;:=\;\frac{dK(s)}{ds}\Big|_{s=0}\;=\;K_{\tilde{\mathcal{N}}}\,-\,K_{\tilde{\mathcal{M}}}\;=\;-\,P~,$ (22)
where $P:=K_{\tilde{\mathcal{M}}}-K_{\tilde{\mathcal{N}}}$ is the generalized modular momentum.
Proof.
Let $\tilde{\phi}_{0}$ be a faithful normal weight on $\tilde{\mathcal{M}}$. We define the weight $\tilde{\phi}_{1}:=\tilde{\phi}_{0}\circ\tilde{E}$, thereby extending the modular dynamics of the subalgebra $\tilde{\mathcal{N}}$ to the full algebra $\tilde{\mathcal{M}}$. By identifying the modular Hamiltonians via the relation $\Delta=e^{-K}$ (specifically, $\Delta_{\tilde{\mathcal{M}}}=e^{-K_{\tilde{\mathcal{M}}}}$ and $\Delta_{\tilde{\mathcal{N}}}=e^{-K_{\tilde{\mathcal{N}}}}$), the Connes–Takesaki Radon–Nikodym cocycle is explicitly given by:
$D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}\;:=\;\Delta_{\tilde{\phi}_{1}}^{it}\Delta_{\tilde{\phi}_{0}}^{-it}~.$ (23)
The canonical interpolation path $\tilde{\phi}_{s}$ is then constructed via the analytic scaling of this cocycle:
$D\tilde{\phi}_{s}:D\tilde{\phi}_{0}]_{t}\;=\;[D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{\,st}~.$ (24)
Using the cocycle identity $D\tilde{\phi}_{s}:D\tilde{\phi}_{0}]_{t}=\Delta_{s}^{it}\Delta_{0}^{-it}$ (where $\Delta_{0}=\Delta_{\tilde{\mathcal{M}}}$), we can express the modular operator $\Delta_{s}$ along the path as:
$\Delta_{s}^{\,it}\;=\;\big[D\tilde{\phi}_{1}:D\tilde{\phi}_{0}\big]_{st}\,\Delta_{0}^{\,it}\,,\qquad\forall\,t\in\mathbb{R}~.$ (25)
We now compute the derivative of (25) with respect to $s$ at the origin. First, observe that the cocycle $D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{u}$ forms a unitary group in the parameter $u=st$, generated by the relative modular Hamiltonian $K_{\tilde{\mathcal{N}}}-K_{\tilde{\mathcal{M}}}$. Applying the chain rule, the derivative of the cocycle term is:
[…]
Consequently, differentiating the full expression in (25) yields:
$\frac{d}{ds}\Delta_{s}^{\,it}\Big|_{s=0}\;=\;\left(\frac{d}{ds}\big[D\tilde{\phi}_{1}:D\tilde{\phi}_{0}\big]_{st}\Big|_{s=0}\right)\Delta_{0}^{\,it}\;=\;-it(K_{\tilde{\mathcal{N}}}-K_{\tilde{\mathcal{M}}})\Delta_{0}^{\,it}~.$ (27)
Alternatively, we can evaluate this derivative using the definition $\Delta_{s}^{it}=e^{-itK(s)}$. Applying Duhamel’s formula for the derivative of an exponential operator Kato , a cornerstone of Araki’s modular perturbation theory Araki:1976zv , we obtain:
$\frac{d}{ds}e^{-itK(s)}\Big|_{s=0}\;=\;-i\int_{0}^{t}e^{-i(t-\tau)K_{0}}K^{\prime}(0)e^{-i\tau K_{0}}d\tau~.$ (28)
[…]
$K^{\prime}(0)\;=\;K_{\tilde{\mathcal{N}}}-K_{\tilde{\mathcal{M}}}~.$ (29)
Recalling the definition $P:=K_{\tilde{\mathcal{M}}}-K_{\tilde{\mathcal{N}}}$, we conclude that $K^{\prime}(0)=-P$. ∎
The result $K^{\prime}(0)=-P$ implies that the infinitesimal change of the modular Hamiltonian with respect to $s$ is governed by the spacetime translation operator $P$. As we vary $s$, the evolution of $K(s)$ is effectively driven by a geometric translation in the emergent spacetime. This confirms that our continuous interpolation path aligns with a physical flow: at $s=0$, the infinitesimal change of the modular Hamiltonian is precisely given by P-P (implying a translation directed opposite to the canonical outward shift). Thus, the abstract path $\tilde{E}_{s}$ is not merely algebraically idempotent; its tangent direction corresponds to a physical translation, substantiating the interpretation of the channel as a geometric transport of information.
Equipped with the result $K^{\prime}(0)=-P$ from Lemma 4.1, we can now explicitly compute $\tilde{G}$ and establish the main theorem.
Theorem 4.2 (The Identity $\tilde{G}=2P$).
The generator $\tilde{G}$ is identical to $2P$ as a closed self-adjoint operator.
Proof.
By definition (see Definition 3.9), the unitary path is $\tilde{U}(s):=J_{\tilde{\mathcal{M}}}\,J_{\tilde{\mathcal{N}}(s)}$. Its derivative at $s=0$ defines $\tilde{G}$ on the common core domain $\mathcal{D}_{\text{core}}$:
$\tilde{G}\;:=\;i\,\frac{d\tilde{U}(s)}{ds}\Big|_{s=0}\;=\;i\,J_{\tilde{\mathcal{M}}}\,\frac{d\,J_{\tilde{\mathcal{N}}(s)}}{ds}\Big|_{s=0}~,$ (30)
in accordance with (19). Here $J_{\tilde{\mathcal{M}}}$ is the modular conjugation for $\tilde{\mathcal{M}}$, and $J_{\tilde{\mathcal{N}}(s)}$ is that for the algebra $\tilde{\mathcal{N}}(s)$.
For a generic perturbation, the variation of $J$ usually involves complex spectral integrals of the modular operator arising from non-commutativity. However, a remarkable feature of the canonical $L^{p}$-interpolation path is that it follows the geodesic of the underlying modular geometry. Consequently, the complex non-commutative terms cancel out, [^2] Explicitly, the path is generated by the analytic scaling $\Delta_{s}^{it}=h^{ist}\Delta_{0}^{it}$. The variation with respect to $s$ extracts the generator $\ln h$ linearly, avoiding the parameter-integral convolution typical of generic perturbations. leaving a simple geometric proportionality:
$i\,J_{\tilde{\mathcal{M}}}\,\frac{dJ(s)}{ds}\Big|_{s=0}\;=\;-\,2\,\frac{dK(s)}{ds}\Big|_{s=0}~.$ (31)
Physically, this factor of 2 is a direct consequence of the geometric role of $J$. Since the modular conjugation implements a geometric reflection (effectively inverting the emergent spacetime), shifting the reflection axis along the geodesic results in a net displacement of twice that magnitude. This geometric logic is inherent to the canonical $L^{p}$ interpolation and justifies the specific coefficient in our operator identity.
Substituting the result of Lemma 4.1 ($K^{\prime}(0)=-P$) into (31), we obtain:
$\tilde{G}\;=\;-\,2\,K^{\prime}(0)\;=\;-\,2(-P)\;=\;2P~,$ (32)
establishing the algebraic identity on the dense domain $\mathcal{D}_{\text{core}}$. Crucially, this exact reconstruction of $2P$ holds regardless of the commutation relations between $K_{\tilde{\mathcal{M}}}$ and $P$, confirming that the result is a robust consequence of the modular structure rather than an artifact of specific symmetries.
It remains to lift the equality $\tilde{G}=2P$ from the core domain to the level of closed operators. By construction, the common analytic core $\mathcal{D}_{\text{core}}$ is invariant under the modular automorphism groups of both $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$. Consequently, $\mathcal{D}_{\text{core}}$ consists of analytic vectors for the modular Hamiltonians $K_{\tilde{\mathcal{M}}}$ and $K_{\tilde{\mathcal{N}}}$, and thereby for their difference $P$. Thus, $2P$ is essentially self-adjoint on this domain. Similarly, $\tilde{G}$ is essentially self-adjoint on $\mathcal{D}_{\text{core}}$ by Nelson’s analytic vector theorem (Lemma 3.10). Since $\tilde{G}$ and $2P$ coincide on a common core where both are essentially self-adjoint, their unique self-adjoint closures are identical. This completes the proof. ∎
4.3 Stability Quantification via Non-Commutative $L^{p}$ Spaces
A central physical question is whether the identity $\tilde{G}=2P$ represents a singular coincidence valid only at the boundary $s=0$, or a structurally robust property of the interpolation. Specifically, we must ensure that the generator does not exhibit large or uncontrolled fluctuations under small perturbations of the interpolation parameter $s$. We address this by employing Kosaki’s non-commutative $L^{p}$ theory Kosaki84 ; Kato , which provides the canonical geometric framework for analyzing perturbations in Type II von Neumann algebras.
Theorem 4.3 (Local Stability of the Generator).
The generator $\tilde{G}$ exhibits structural stability in the strong resolvent sense. Let $\tilde{G}(s)$ denote the instantaneous generator of the unitary path at parameter $s$. For any fixed complex number $z$ in the resolvent set $\rho(2P)$ (i.e., $z\notin\sigma(2P)$) and any state vector $\xi$ within the dense core domain $\mathcal{D}_{\text{core}}$, the deviation of the resolvent is quantitatively bounded by the variation of the associated projection $e_{\tilde{\mathcal{N}}(s)}$:
[…]
Here, $R_{T}(z)=(T-z)^{-1}$ denotes the resolvent operator, and $||\cdot||$ represents the vector norm on the Hilbert space (or the operator norm for the projection difference). The coefficient $C_{z}$ is a stability factor that scales inversely with the distance to the spectrum $\sigma(2P)$ (i.e., $C_{z}\sim\operatorname{dist}(z,\sigma(2P))^{-1}$).
Proof.
The proof establishes the local stability by applying perturbation theory for unbounded operators.
  1. 1.
[…]
  2. 2.
    Kato-Type Bound: We employ the standard first-order perturbation estimate for resolvents. The perturbation of $\tilde{G}(s)$ away from $\tilde{G}(0)=2P$ is effectively controlled by the distance between the projections. The constant $C_{z}$ is bounded by the inverse distance to the spectrum, $C_{z}\lesssim 1/|\operatorname{Im}(z)|$ Kato .
  3. 3.
    Quantification: For the canonical path, the projection perturbation scales linearly with $s$ near the origin: $e_{\tilde{\mathcal{N}}(s)}-e_{\tilde{\mathcal{M}}}||=O(s)$.
This confirms that the generator $\tilde{G}(s)$ converges to $2P$ in the strong resolvent sense as $s\to 0$, affirming the structural robustness of the $\tilde{G}=2P$ identity against small deformations. ∎
4.4 Correlation-Function Test of the Conjecture
The operator identity $\tilde{G}=2P$ posits a direct equivalence between the algebraic shift derived from modular theory and the geometric translation of the emergent spacetime. While operator-level identities are mathematically rigid, their physical validity in complex quantum gravity models requires verification through observables. We propose a concrete test using two-point correlation functions within the AdS/CFT correspondence.
We frame this test in the context of a two-sided eternal black hole, holographically dual to the thermofield double state $|\text{TFD}\rangle$. The local algebras $\mathcal{M}$ and $\mathcal{N}$ correspond to nested subregions of the boundary CFT, specifically associated with the Right exterior region. Let $O_{L}$ and $O_{R}$ be local probe operators acting on the Left and Right boundaries, respectively. Since the canonical shift $\tilde{U}(s)$ is generated by the algebra on the Right, it acts non-trivially on $O_{R}$.
To probe whether this algebraic action corresponds to a geometric translation, we define a correlation function $F(s)$ that measures how the Left-Right entanglement implies a correlation change under the algebraic shift of the Right operator:
$F(s)\;:=\;\langle\text{TFD}|\,O_{L}\,\left(\tilde{U}(s)O_{R}\tilde{U}(s)^{*}\right)\,|\text{TFD}\rangle.$ (34)
This correlator probes the “distance” through the wormhole between the fixed Left operator and the algebraically shifted Right operator. Since $\tilde{G}$ is defined as the infinitesimal generator of the path at $s=0$ (satisfying the expansion $\tilde{U}(s)=\mathbf{1}-is\tilde{G}+O(s^{2})$), the initial derivative determines the prediction for the operator identity. If $\tilde{G}=2P$ holds, we obtain:
$F^{\prime}(0)\;=\;\langle\text{TFD}|\,O_{L}\,(-i[\tilde{G},O_{R}])\,|\text{TFD}\rangle\;=\;-2i\langle\text{TFD}|\,O_{L}\,[P,O_{R}]\,|\text{TFD}\rangle.$ (35)
Eq. (35) provides the precise physical interpretation that the algebraic generator $\tilde{G}$ acting on the boundary operator induces a shift exactly twice that induced by the geometric translation generator $P$.
In the bulk gravity description, $P$ generates a geometric time translation along the horizon. Therefore, verifying Eq. (35) amounts to checking if the algebraic shift $\tilde{U}(s)$ effectively displaces the position of the horizon (or the stretched horizon) relative to the probe $O_{L}$ by a coordinate distance of $2s$.
This geometric action is deeply connected to the mechanics of holographic teleportation. In models like the Gao-Jafferis-Wall (GJW) protocol Gao2017 , the traversability is achieved by a gravitational shockwave that shifts the horizon. Here, our result implies that the algebraic generator $\tilde{G}$ explicitly implements this horizon shift without needing an external matter source, effectively simulating the backreaction required to bridge the entanglement wedge.
In semiclassical models like Jackiw-Teitelboim (JT) gravity Saad:2019lba , the quantity $2i\langle O_{L}[P,O_{R}]\rangle$ precisely quantifies this shift, manifesting as a Shapiro time delay for signals crossing the wormhole Shenker2014 ; Maldacena2016 . Thus, the factor of 2 confirms that $\tilde{U}(s)$ physically manipulates the geometry of the stretched horizon, facilitating the information transfer.
5 Conclusion
We have presented a framework connecting the discrete, algebraic teleportation protocol of vdH-V vdHV with the continuous, geometric picture provided by modular flow. In Type III settings, a key challenge is to construct a path that is both mathematically controlled and physically consistent. Our construction addresses this by defining a strongly continuous unitary path $\tilde{U}(s)$ with a well-defined self-adjoint generator $\tilde{G}$, satisfying the relation
$\tilde{G}\;=\;2\big(K_{\tilde{\mathcal{M}}}-K_{\tilde{\mathcal{N}}}\big)\;=\;2P~.$ (36)
The unitary $\tilde{U}(s)\approx e^{-is\tilde{G}}$ describes an effective transport of quantum information. At $s=0$, the information resides in the relative commutant $\tilde{\mathcal{A}}=\tilde{\mathcal{N}}^{\prime}\cap\tilde{\mathcal{M}}$ (corresponding to the island region just behind the horizon) and is dynamically inaccessible. As $s$ increases, the operator smoothly transports this information into the accessible radiation sector $\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$, where $\tilde{\mathcal{M}}_{1}(s):=\langle\tilde{\mathcal{M}},e_{\tilde{\mathcal{N}}(s)}\rangle$. The endpoint corresponds to the canonical shift unitary $U_{\tilde{\Gamma}}$, where the factor of $2$ arises because $U_{\tilde{\Gamma}}$ is generated by two modular conjugations, effectively doubling the modular distance. In this view, the recovery of information via teleportation becomes algebraically analogous to a geometric spacetime translation generated by $P$, giving a precise form to the correspondence Teleportation=Translation.
This perspective offers a consistent framework for approaching the black hole information paradox. Within this picture, global evolution is unitary: the operator $\tilde{U}(s)$ transports interior information to the exterior algebra rather than destroying it, suggesting that no information is lost during the evaporation process. No-cloning is naturally preserved by algebraic complementarity, as the information is encoded in distinct algebras across different values of $s$ without duplication. The apparent thermality can be interpreted as a consequence of the Type III algebraic structure—specifically the absence of a tracial state—rather than an intrinsic feature of the global dynamics. Furthermore, lifting the description to the semifinite envelope clarifies that entropy and thermality are governed by modular, as opposed to tracial, structures.
The property of dynamic idempotency induces a nested flow of algebras, offering a controlled, continuous realization of the Page/Hayden–Preskill decoding protocol. This framework shares key conceptual features with holographic quantum error correction and entanglement wedge reconstruction. In particular, the smooth transport of information from the initially inaccessible relative commutant to the radiation sector is analogous to the recoverability conditions in error-correcting codes Almheiri_2015 . Similarly, the generator $\tilde{G}=2P$ provides an algebraic description of accessing the island interior, paralleling the geometric reconstruction of the entanglement wedge Penington19 .
[…]
In summary, this formal development indicates that continuous, idempotent, and modular-covariant teleportation is algebraically equivalent to a spacetime translation. This identification provides a unitary and calculable channel for information transfer, supporting the view that black holes may evaporate in accordance with the principles of quantum mechanics and algebraic locality.
It is worth noting that these results are derived within an algebraic QFT framework on a fixed background spacetime. While we demonstrate that information flow remains unitary and consistent with geometric translation in this setting, this work primarily provides a kinematic framework for information recovery; the full quantum gravitational dynamics, including the back-reaction of the evaporation process on the geometry, remains an open and important frontier.
Our work suggests several directions for future research. First, the connection between $\tilde{G}=2P$ and traversable wormholes warrants deeper investigation. Since $P$ generates translations across the horizon, our continuous protocol may provide an algebraic description of the GJW protocol Gao2017 , potentially offering a non-perturbative perspective on the wormhole opening mechanism via modular flow. Second, verifying the correlation function equality $F^{\prime}(0)\propto-i\langle[P,\mathcal{O}]\rangle$ in holographic models, such as double-scaled Sachdev-Ye-Kitaev (DSSYK) Berkooz:2018jqr or JT gravity, would help bridge our algebraic results with semiclassical gravity calculations.
Finally, extending this framework to incorporate gravitational back-reaction is a critical next step. Since our protocol $\tilde{U}(s)$ is explicitly constructed within the crossed-product envelope—a structure known to capture $1/N$ corrections and observer energy constraints Witten22 ; Chandrasekaran2022cip ; LeutheusserLiu_ET —it offers a natural starting point for such an extension. Formulating the interplay between the information transport $\tilde{U}(s)$ and the semi-classical Einstein equations in this Type II setting could pave the way for a dynamic theory of emergent spacetime geometry from quantum entanglement.
Acknowledgements.
I acknowledge the assistance of Large Language Models, specifically Gemini and Chatgpt, in drafting the text, providing valuable feedback, and identifying errors. I take full intellectual responsibility for the content of this work. This work is supported by the National Research Foundation of Korea(NRF) grant with grant number RS-2023-00208047, NRF-2020R1A6A1A03047877, RS-2025-00553127. APPENDIX:
Appendix A Notation table
Symbol Meaning
$\mathcal{M},\ \mathcal{N}$ von Neumann algebras ($\mathcal{N}\subset\mathcal{M}$)
$\tilde{\mathcal{M}},\ \tilde{\mathcal{N}}$ semifinite crossed-product algebras (envelopes)
$\mathcal{X}^{+}$, $\widehat{\mathcal{X}}^{+}$ positive and extended positive cones of $\mathcal{X}$
[…]
$\tilde{E}$, $\tilde{E}_{s}$ lifted conditional expectation and its idempotent interpolants
$e_{\tilde{\mathcal{N}}},\ e_{\tilde{\mathcal{N}}(s)}$ Jones projection and interpolated projections onto subalgebras
$U_{\Gamma}$, $U_{\tilde{\Gamma}}$ discrete and lifted canonical shift unitaries
$\quad\tilde{\mathcal{A}}$ relative commutant $\tilde{\mathcal{N}}^{\prime}\cap\tilde{\mathcal{M}}$ (island algebra)
$\tilde{\mathcal{M}}_{1}(s)$ basic extension algebra $\langle\tilde{\mathcal{M}},e_{\tilde{\mathcal{N}}(s)}\rangle$
$\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$ accessible radiation sector (basic extension’s commutant)
$\tilde{U}(s)$, $\tilde{G}$ continuous unitary path and its self-adjoint generator
$h=d\tilde{\phi}_{1}/d\tilde{\phi}_{0}$ non-commutative Radon–Nikodym derivative
$K_{\mathcal{X}}$, $P$ modular Hamiltonian of $\mathcal{X}$ and modular momentum (difference of Hamiltonians)
$\quad\sigma_{t}^{\phi}$ modular automorphism group of weight $\phi$
$J_{\mathcal{X}}$, $\Delta_{\mathcal{X}}$ modular conjugation and modular operator
$\quad\tau$ canonical trace on the crossed product $\tilde{\mathcal{M}}$
$\tilde{\phi}_{0}$, $\tilde{\phi}_{1}$ faithful normal weights on $\tilde{\mathcal{M}}$
Table 1: Quick reference for frequently used symbols.
Appendix B Finite-Dimensional Model: Idempotency Failure and Unitary Interpolation
This appendix details the explicit calculation demonstrating the failure of the naive CP-map interpolation (Path A) to satisfy idempotency in finite dimensions, and provides a simple example of the required continuous unitary path.
Example B.1 (Failure of Path A in $M_{2}(\mathbb{C})$).
Consider $\mathcal{M}=\mathrm{M}_{2}(\mathbb{C})$ and let $\mathcal{N}=\mathbb{C}\mathbf{1}$ be the subalgebra of scalar matrices. The unique trace-preserving conditional expectation $E:\mathcal{M}\rightarrow\mathcal{N}$ is the normalized trace, $E(A)=\frac{1}{2}\mathrm{Tr}(A)\mathbf{1}$. The identity map is $\mathrm{id}(A)=A$. Using the convex combination (Path A), the interpolated map is:
$E_{s}(A)=(1-s)A+\frac{s}{2}\mathrm{Tr}(A)\mathbf{1}$ (37)
At the midpoint $s=0.5$, the map becomes $E_{0.5}(A)=0.5A+0.25\mathrm{Tr}(A)\mathbf{1}$. Applying the map again to test idempotency yields:
$E_{0.5}^{2}(A)=E_{0.5}(0.5A+0.25\mathrm{Tr}(A)\mathbf{1})=0.25A+0.375\mathrm{Tr}(A)\mathbf{1}.$ (38)
Since $E_{0.5}^{2}(A)\neq E_{0.5}(A)$, this demonstrates the fatal failure of static idempotency for the naive path. This failure is precisely why the Haagerup–Kosaki lift is required in the Type III setting.
Example B.2 (Continuous Unitary Interpolation Goal).
Consider the same toy model $\mathcal{M}=\mathrm{M}_{2}(\mathbb{C})$. Define $U_{\Gamma}$ to be the canonical shift unitary that swaps the two basis states (analogous to the Pauli X operator), $U_{\Gamma}|0\rangle=|1\rangle$, $U_{\Gamma}|1\rangle=|0\rangle$. This discrete teleportation unitary can be continuously approached by a path of rotations. We construct a continuous path $U(s)=e^{-isG}$ using the self-adjoint generator:
$G=\frac{\pi}{2}\sigma_{x}=\frac{\pi}{2}(|0\rangle\langle 1|+|1\rangle\langle 0|)$ (39)
such that $U(0)=\mathbf{1}$. At $s=1$, we find:
$U(1)=e^{-i\frac{\pi}{2}\sigma_{x}}=\cos(\frac{\pi}{2})\mathbf{1}-i\sin(\frac{\pi}{2})\sigma_{x}=-i\sigma_{x}\propto U_{\Gamma}$ (40)
Thus, the endpoint coincides with the swap up to a global phase factor of i-i. This illustrates the physical goal of the construction: finding a continuous evolution $U(s)$ that connects the initial state to the final shift. This unitary path serves as a consistent alternative to the naive CP-map interpolation, which was shown to be mathematically problematic in Example B.1.
Appendix C Haagerup–Kosaki Lift: Existence and Mechanism of $\tilde{E}$
The most critical technical challenge in defining the idempotent path $E_{s}$ in the Type III setting is the construction of a faithful normal conditional expectation. In general, Type III inclusions do not admit a conditional expectation $\mathcal{M}\to\mathcal{N}$. The theory of Haagerup and Takesaki provides the necessary mechanism by lifting the problem from a non-tracial Type III algebra (equipped with an operator-valued weight, OVW) to a tracial semifinite algebra where a genuine conditional expectation $\tilde{E}$ exists. This lift effectively regularizes the unbounded OVW $\mathcal{E}$ via the crossed product construction.
Proposition C.1 (Haagerup–Kosaki lift and $\tilde{E}$ existence).
Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of Type III von Neumann algebras. While a conditional expectation may not exist, Haagerup’s theorem guarantees the existence of a faithful normal OVW $\mathcal{E}:\mathcal{M}^{+}\to\widehat{\mathcal{N}}^{+}$.
Suppose we select a faithful normal weight $\omega$ on $\mathcal{M}$ that is compatible with the inclusion, meaning its modular automorphism group $\sigma_{t}^{\omega}$ leaves the subalgebra invariant:
$\sigma_{t}^{\omega}(\mathcal{N})\subset\mathcal{N}\quad\text{for all }t\in\mathbb{R}.$ (41)
(This condition implies that $\omega$ can be written as $\omega=\psi\circ\mathcal{E}$ for some faithful normal weight $\psi$ on $\mathcal{N}$).
Under this modular covariance assumption, the crossed product construction yields the following structure:
  1. 1.
[…]
    $\tau\circ\theta_{t}=e^{-t}\tau.$ (42)
    Furthermore, the connection to the base algebra is maintained by the canonical OVW $\underline{\mathcal{E}}:\tilde{\mathcal{M}}^{+}\to\widehat$$\mathcal{M}$$^{+}$ (often denoted as the dual weight construction), formally given by:
[…]
  2. 2.
    The subalgebra inclusion lifts to $\tilde{\mathcal{N}}=\mathcal{N}\rtimes_{\sigma^{\omega}|_{\mathcal{N}}}\mathbb{R}\subset\tilde{\mathcal{M}}$. Crucially, this lifted subalgebra remains invariant under the dual action:
    $\theta_{t}(\tilde{\mathcal{N}})=\tilde{\mathcal{N}}$. (44)
  3. 3.
    There exists a unique, faithful normal conditional expectation
    $\tilde{E}:\tilde{\mathcal{M}}\longrightarrow\tilde{\mathcal{N}}$ (45)
    which is trace-preserving (i.e., $\tau\circ\tilde{E}=\tau$) and is equivariant with respect to the dual action $\theta_{t}$ (i.e., $\tilde{E}\circ\theta_{t}=\theta_{t}\circ\tilde{E}$).
Proof.
The proof formalizes the conversion of the Type III unbounded structure into a Type $\text{II}_{\infty}$ bounded projection.
  1. 1.
    Compatibility: The assumption $\sigma_{t}^{\omega}(\mathcal{N})\subset\mathcal{N}$ ensures that the crossed product of the subalgebra, $\tilde{\mathcal{N}}$, is naturally a subalgebra of $\tilde{\mathcal{M}}$. Without this modular covariance, the relationship between the lifted algebras is not well-defined.
  2. 2.
    Semifinite Structure and Trace: The theory of crossed products guarantees that $\tilde{\mathcal{M}}$ is a semifinite algebra equipped with a canonical trace $\tau$. The scaling property (42) is a defining characteristic of the Type II crossed product derived from a Type III algebra, allowing us to treat modular dynamics as trace-scaling automorphisms. Consequently, the existence of this trace allows for the construction of conditional expectations that are not available in the Type III setting. ă
  3. 3.
    Construction of $\tilde{E}$: The existence of the trace-preserving conditional expectation $\tilde{E}$ follows from the compatibility of the dual weights on $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$. Specifically, because the original weights were chosen to be compatible ($\omega=\psi\circ\mathcal{E}$), their dual traces on the crossed products coincide, implying the existence of a conditional expectation. This $\tilde{E}$ effectively renormalizes the original OVW $\mathcal{E}$ into a norm-one projection on the semifinite algebra. ă
  4. 4.
    Equivariance: The construction explicitly preserves the structure of the dual action $\theta_{t}$, ensuring $\tilde{E}\circ\theta_{t}=\theta_{t}\circ\tilde{E}$. This property is crucial for proving that the analytic domains are preserved under the operation of $\tilde{E}$.
For rigorous proofs and technical details, see TakesakiII ; Kosaki84 . ∎