### v4

# Teleportation=Translation: Continuous recovery of black hole information

Abstract
The Teleportation=Translation conjecture posits that the recovery of information from a black hole is dual to a geometric translation in the emergent spacetime. In this paper, we establish this equivalence for general local quantum field theories by constructing a continuous unitary interpolation that bridges discrete algebraic teleportation protocols and continuous modular flow. We resolve the failure of dynamic idempotency, fundamentally inherent in Type III von Neumann algebras, by employing the Haagerup-Kosaki crossed-product construction. This lift to the semifinite Type II envelope yields a canonical, dynamically consistent path. Crucially, we prove that its unique infinitesimal generator $\tilde{G}$ is exactly twice the geometric modular momentum ($\tilde{G}=2P$). We establish this identity as a closed operator equivalence using Nelson’s analytic vector theorem and quantify its structural robustness via non-commutative $L^{p}$ theory. Ultimately, our results demonstrate that unitary information recovery fundamentally manifests as a continuous geometric translation. This provides a rigorous operator-algebraic mechanism for resolving the black hole information paradox, offering a kinematic framework naturally extendable to include gravitational back-reaction.
1 Introduction
The black hole information paradox exposes a fundamental tension between the apparent thermality of Hawking radiation, derived from semiclassical gravity, and the unitary evolution required by quantum theory Hawking1976 . Although Hawking’s original calculation suggested that a semiclassical black hole evaporates via thermal radiation in a manner violating unitarity, recent developments indicate that information recovery is possible even within the semiclassical framework, provided the problem is analyzed with refined conceptual and technical tools Almheiri2020 ; Raju2022 ; Akil:2025bhip .
In particular, the AdS/CFT correspondence has provided a semiclassical framework utilizing replica wormholes and quantum extremal surfaces to derive the Page curve Almheiri2019 ; Almheiri2020Page ; Penington2022 . While these results strongly support the preservation of unitarity, they raise critical questions regarding the interpretation of the gravitational path integral as an ensemble average SaadShenkerStanford and its compatibility with spacetime factorization HarlowJafferis2020 ; Marolf2020 . To address these foundational issues, quantum information and operator-algebraic approaches have been increasingly employed to clarify the underlying assumptions behind these entropy computations Raju2022 ; Almheiri_2015 ; Penington19 ; Kibe2022 .
However, calculating the entropy curve does not, by itself, explain the dynamical mechanism of information retrieval. To address the retrieval process, Hayden and Preskill reformulated the problem as a quantum decoding task, demonstrating that a black hole acting as a fast scrambler allows for rapid information retrieval, provided the observer has access to the early radiation HaydenPreskill . This recovery process is operationally analogous to quantum teleportation. Building on this operational viewpoint, van den Heijden and Verlinde (vdH-V) reformulated the problem entirely within the language of operator algebras vdHV , with mathematical foundations for such protocols further developed in Conlon2022 . In their framework, passing to the continuum limit transforms the canonical shift into a continuous teleportation operator. Furthermore, they demonstrated that in the special symmetric case of a half-sided modular inclusion (HSMI), this naturally leads to the identification Teleportation = Translation.
Nonetheless, while the continuous limit naturally emerges under the special symmetric conditions of a HSMI, extending this protocol to the continuum limit for general local quantum field theories faces fundamental obstructions, as acknowledged in vdHV . Local operator algebras in quantum field theory are Type III von Neumann factors, distinguished by the absence of a tracial state Connes80 ; Witten . This algebraic feature reflects the physical reality that vacuum entanglement is effectively unbounded. Consequently, the maximally entangled resources prerequisite for standard teleportation are mathematically ill-defined in this general setting.
In this work, we generalize this paradigm to arbitrary local quantum field theories. We construct a mathematically rigorous and physically consistent smooth unitary interpolation that connects the discrete canonical shift with continuous spacetime evolution. A central challenge is ensuring that this interpolation path is canonical rather than arbitrary. To resolve this in the general Type III context, we employ the Haagerup–Kosaki framework Connes80 ; Kosaki84 and lift the inclusion to a Type II crossed product via Takesaki duality Connes80 ; TakesakiII . Crucially, the conditional expectation in this setting is uniquely determined by the invariance of the reference weights defining the $L^{p}$-interpolation Kosaki84 ; Takesaki72 . This rigidity guarantees that our unitary path is intrinsic to the modular structure of the system. Furthermore, by exploiting the analytic properties of modular flow, we ensure the path is sufficiently smooth to rigorously define the infinitesimal generator $\widetilde{G}$ at the interpolation boundary $s=0$ as the unique self-adjoint operator driving this flow.
Building on this construction, we derive our main result: the generator of the teleportation flow bears a direct physical identification with the modular momentum. Specifically, we establish an exact operator identity asserting that the infinitesimal generator equals twice the difference of the modular Hamiltonians, $\widetilde{G}=2P$. The factor of 2 is dictated by Borchers’ commutation relations Borchers92 , ensuring consistency with the scaling of modular flows. Our proof establishes this equality on a dense core of analytic modular vectors and extends it to the closed operators via Nelson’s analytic vector theorem Nelson59 . Moreover, we quantify the stability of this identification under small deformations using non-commutative $L^{p}$-theory Kosaki84 ; Kato .
On a physical level, this result directly addresses the ultimate motivation of this work: the black hole information paradox. It demonstrates that infalling information is not lost, but rather unitarily recovered through a continuous path. From an algebraic viewpoint, this recovery process constitutes quantum teleportation, which our exact operator identity proves is dynamically equivalent to a geometric translation in spacetime. By rigorously establishing this Teleportation = Translation identity for general Type III inclusions—without relying on the symmetries of a HSMI—our work confirms that information recovery fundamentally manifests as a continuous geometric translation. Finally, to facilitate verification in complex physical models, we propose a straightforward test based on two-point correlation functions.
The remainder of this paper is organized as follows. In Sec. 2, we summarize the necessary elements of Tomita–Takesaki theory and introduce the Jones projection together with the canonical shift. In Sec. 3, we construct the continuous unitary interpolation by lifting the Type III inclusion to the crossed product via the Haagerup–Kosaki framework. Sec. 4 is devoted to the proof of our main result, identifying the generator of this flow with the modular momentum ($\widetilde{G}=2P$). We conclude with a discussion of the results in Sec. 5.
Technical details are deferred to the appendices. Appendix A collects the notation used throughout the paper. Appendix B illustrates the failure of naive interpolation in finite dimensions and provides a concrete continuous unitary path. Finally, Appendix C details the Haagerup-Takesaki construction of faithful conditional expectations in Type III settings via operator-valued weights.
2 The Discrete Protocol: Canonical Shift
2.1 Preliminaries: Tomita–Takesaki Theory
We work on a fixed Gelfand-Naimark-Segal (GNS) Hilbert space $\mathcal{H}$ TakesakiI built from a faithful normal state $\omega$ (with cyclic and separating vector $|\Omega\rangle$) for the von Neumann algebra $\mathcal{X}$. The core objects of Tomita–Takesaki theory TakesakiII are:
  • The Tomita operator $S_{\mathcal{X}}$, defined on the dense domain $\mathcal{X}|\Omega\rangle$ by $S_{\mathcal{X}}x|\Omega\rangle=x^{*}|\Omega\rangle$ for $x\in\mathcal{X}$.
  • Its polar decomposition $S_{\mathcal{X}}=J_{\mathcal{X}}\Delta_{\mathcal{X}}^{1/2}$, where $J_{\mathcal{X}}$ is the anti-unitary modular conjugation and $\Delta_{\mathcal{X}}$ is the positive self-adjoint modular operator.
  • The modular Hamiltonian $K_{\mathcal{X}}$, defined by $\Delta_{\mathcal{X}}=e^{-K_{\mathcal{X}}}$.
  • The modular automorphism group $\sigma_{t}^{\mathcal{X}}(a)=\Delta_{\mathcal{X}}^{it}a\Delta_{\mathcal{X}}^{-it}$.
These objects satisfy $J_{\mathcal{X}}\mathcal{X}J_{\mathcal{X}}=\mathcal{X}^{\prime}$, where $\mathcal{X}^{\prime}$ is the commutant of $\mathcal{X}$.
2.2 The Canonical Shift Protocol
The discrete vdH-V protocol is built from three components vdHV .
Step 1: Conditional Expectation. Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of von Neumann algebras and $E:\mathcal{M}\to\mathcal{N}$ be a normal, faithful conditional expectation that is $\omega$-preserving: $\omega(E(a))=\omega(a)$ for all $a\in\mathcal{M}$. This map effectively discards information in $\mathcal{M}$ that is not in $\mathcal{N}$.
Step 2: Jones Projection. In the GNS representation where $|a\rangle:=a|\Omega\rangle$, $E$ is implemented by the Jones projection $e_{\mathcal{N}}$, an operator $e_{\mathcal{N}}\in\mathcal{B}(\mathcal{H})$ satisfying
$e_{\mathcal{N}}|a\rangle=|E(a)\rangle.$ (1)
$E$ being a $\omega$-preserving conditional expectation ensures that $e_{\mathcal{N}}$ is a well-defined self-adjoint projection ($e_{\mathcal{N}}^{2}=e_{\mathcal{N}}$, $e_{\mathcal{N}}^{\dagger}=e_{\mathcal{N}}$) and satisfies $e_{\mathcal{N}}ae_{\mathcal{N}}=E(a)e_{\mathcal{N}}$ for all $a\in\mathcal{M}$.
Step 3: Canonical Shift. The protocol is enacted by the unitary operator
$U_{\Gamma}:=J_{\mathcal{M}}J_{\mathcal{N}},$ (2)
which induces the automorphism (the shift) $\Gamma(a)=U_{\Gamma}aU_{\Gamma}^{\dagger}$. This map $\Gamma$ acts as a teleportation map, sending the relative commutant $\mathcal{N}^{\prime}\cap\mathcal{M}$ (information in $\mathcal{M}$ hidden from $\mathcal{N}$) to the relative commutant $\mathcal{M}^{\prime}\cap\mathcal{M}_{1}$ (information in the basic extension $\mathcal{M}_{1}:=\langle\mathcal{M},e_{\mathcal{N}}\rangle$ hidden from $\mathcal{M}$):
$\Gamma(\mathcal{N}^{\prime}\cap\mathcal{M})=\mathcal{M}^{\prime}\cap\mathcal{M}_{1}.$ (3)
Information is not lost, but relocated from one relative commutant to another. This suggests that the algebraic shift operator $U_{\Gamma}$ acts as a geometric translation that maps the interior sector $\mathcal{N}^{\prime}\cap\mathcal{M}$ to the radiation sector $\mathcal{M}^{\prime}\cap\mathcal{M}_{1}$. This identification between algebraic relocation and spacetime translation will be established precisely in the continuum limit in the subsequent sections.
3 From a Discrete Protocol to a Continuous Path
Our core strategy is to construct a continuous teleportation flow. To this end, we first discuss the physical requirement of dynamic idempotency in Sec. 3.1. In Sec. 3.2, we present a pedagogical interpolation (Path A), which, despite satisfying several key physical properties, ultimately fails to maintain idempotency. To resolve this, we lift the problem to the semifinite crossed-product (Haagerup–Kosaki envelope) and construct a canonical physical interpolating path via non-commutative $L^{p}$-space interpolation theory (Path B), which rigorously satisfies the idempotency requirement (Sec. 3.3). Finally, by employing Tomita–Takesaki machinery in this semifinite setting, we build a unitary path $\tilde{U}(s)$ and analytically establish its generator $\tilde{G}$ as the tangent vector at $s=0$, proving its self-adjointness through the theory of analytic vectors (Sec. 3.4).
3.1 The Physical Requirement: Dynamical Idempotency
To define a continuous unitary path $U(s)=J_{\mathcal{M}}J_{\mathcal{N}(s)}$ for $s\in[0,1]$, we must first construct a continuous family of von Neumann subalgebras $\mathcal{N}(s)$ interpolating between the full algebra $\mathcal{N}(0)=\mathcal{M}$ and the target subalgebra $\mathcal{N}(1)=\mathcal{N}$. This path of algebras is operationally characterized by a corresponding family of conditional expectations $E_{s}:\mathcal{M}\to\mathcal{N}(s)$, where the connection between the map $E_{s}$ and the subalgebra $\mathcal{N}(s)$ is fundamentally established by the property of idempotency.
However, for the parameter $s$ to be physically interpreted as a coarse-graining scale or an information resolution flow, a condition stronger than simple static idempotency ($E_{s}^{2}=E_{s}$) is required. A consistent physical flow must satisfy dynamic idempotency (or the semigroup property): $E_{s^{\prime}}\circ E_{s}=E_{s^{\prime}}$ for $s^{\prime}\geq s$. This condition ensures that the sequence of coarse-graining operations is consistent—strictly analogous to the renormalization group flow—and implies a strict nesting of information resources. Mathematically, this prevents information leakage and guarantees that the image of the map at every step retains the structure of a von Neumann algebra. Consequently, it ensures the monotonicity of the information recovery flow from the black hole interior to the exterior.
Definition 3.1.
Dynamic idempotency for a family of maps $\{E_{s}\}_{s\in[0,1]}$ is defined by the following conditions:
$E_{s}^{2}=E_{s}\qquad\text{(static idempotency)},$
$E_{s^{\prime}}\circ E_{s}=E_{s^{\prime}}\quad\text{for }s^{\prime}\geq s\qquad\text{(semigroup property)}.$ (4)
By Tomiyama’s theorem Tomiyama57 , any normal, unital, completely positive (CP) map satisfying static idempotency $E^{2}_{s}=E_{s}$ (along with state invariance) constitutes a faithful normal conditional expectation onto the von Neumann subalgebra $\mathcal{N}(s):=E_{s}(\mathcal{M})$. Our objective is thus reduced to constructing a continuous path of such idempotent CP maps that further satisfies the dynamic consistency requirement.
3.2 Path A: Pedagogical Interpolation
The most intuitive approach to connecting the identity map ($\mathrm{id}$) to the target conditional expectation ($E$) is to construct a linear interpolation. The primary constraint for any physical process is that every intermediate map must constitute a valid quantum channel—that is, a unital CP map.
The most direct construction satisfying this requirement is the convex combination of the endpoint maps:
$E_{s}=(1-s)\mathrm{id}+sE,\quad s\in[0,1].$ (5)
Since both $\mathrm{id}$ and $E$ are unital CP maps, their convex combination $E_{s}$ inherently preserves these properties. Thus, $E_{s}$ remains a well-defined quantum channel for all $s$, even in the Type III setting.
  • Finite Dimensions: In matrix algebras, this construction is equivalent to linearly interpolating the corresponding Choi matrices Choi75 . Since the Choi matrices of the endpoint maps $\mathrm{id}$ and $E$ are positive semidefinite, their convex combination preserves this property, ensuring complete positivity along the entire path.
  • Infinite Dimensions: For general von Neumann algebras (including Type III factors), Stinespring’s dilation theorem Stinespring55 guarantees that every intermediate map $E_{s}$ admits a physical realization. Specifically, there exists a dilation Hilbert space $\mathcal{K}_{s}$, a representation $\pi_{s}:\mathcal{M}\to\mathcal{B}(\mathcal{K}_{s})$, and an isometry $V_{s}:\mathcal{H}\to\mathcal{K}_{s}$ such that the map acts as a compression:
    $E_{s}(x)=V_{s}^{\dagger}\pi_{s}(x)V_{s},\quad\forall x\in\mathcal{M}.$
These constructions provide a continuous path of maps $E_{s}$ that satisfy the fundamental mathematical properties of is mathematically well-definedness, boundedness, and self-adjointness. We first establish these properties for the map $E_{s}$ itself, and subsequently for the corresponding operator $e_{\mathcal{N}(s)}$ on the GNS Hilbert space.
Lemma 3.2.
(Properties of the Interpolating Map $E_{s}$) For all $s\in[0,1]$, the map $E_{s}$ defined above is:
  1. 1.
    $\omega$-preserving: $\omega(E_{s}(a))=\omega(a)$ for all $a\in\mathcal{M}$.
  2. 2.
    Self-adjoint with respect to the GNS inner product associated with $\omega$: $\omega(y^{*}E_{s}(x))=\omega((E_{s}(y))^{*}x)$ for all $x,y\in\mathcal{M}$.
Proof.
(1) State-preserving: By assumption, $\omega\circ E=\omega$. Since $E_{s}$ is a linear combination,
$\omega(E_{s}(a))=(1-s)\omega(a)+s\,\omega(E(a))=\omega(a).$ (6)
(2) GNS Self-adjointness: The condition for self-adjointness with respect to the GNS inner product holds for both $\mathrm{id}$ and the $\omega$-preserving conditional expectation $E$. By linearity,
$\omega(y^{*}E_{s}(x))$ $=(1-s)\omega(y^{*}x)+s\,\omega(y^{*}E(x))$
$=(1-s)\omega(\mathrm{id}(y)^{*}x)+s\,\omega(E(y)^{*}x)$
$=\omega(((1-s)\mathrm{id}(y)+sE(y))^{*}x)$
$=\omega((E_{s}(y))^{*}x).$ (7)
Using this lemma, we can rigorously define the interpolated operator on the Hilbert space.
Theorem 3.3.
(Existence and Properties of $e_{\mathcal{N}}(s)$) For each $s\in[0,1]$, the operator $e_{\mathcal{N}}(s)$ defined on the dense domain $D=\{|a\rangle\mid a\in\mathcal{M}\}\subset\mathcal{H}$ by
$e_{\mathcal{N}(s)}|a\rangle:=|E_{s}(a)\rangle$ (8)
extends uniquely to a bounded, self-adjoint linear operator on $\mathcal{H}$ with norm $\|e_{\mathcal{N}}(s)\|\leq 1$. Furthermore, it satisfies the boundary conditions $e_{\mathcal{N}(0)}=\mathbf{1}$ and $e_{\mathcal{N}(1)}=e_{\mathcal{N}}$.
Proof.
(1) Boundedness: Using the Kadison-Schwarz inequality for the unital CP map $E_{s}$, we have $E_{s}(a)^{*}E_{s}(a)\leq E_{s}(a^{*}a)$. Evaluating in the state $\omega$ (which is preserved by $E_{s}$):
$\||E_{s}(a)\rangle\|^{2}=\omega(E_{s}(a)^{*}E_{s}(a))\leq\omega(E_{s}(a^{*}a))=\omega(a^{*}a)=\||a\rangle\|^{2}.$ (9)
This implies well-definedness (respecting GNS equivalence classes) and boundedness ($\|e_{\mathcal{N}(s)}\|\leq 1$), allowing a unique extension to $\mathcal{H}$.
(2) Self-adjointness: The sesquilinear form is symmetric due to the self-adjointness of $E_{s}$ (Lemma 3.2):
$\langle|a\rangle|e_{\mathcal{N}(s)}|b\rangle\rangle=\omega(a^{*}E_{s}(b))=\omega(E_{s}(a)^{*}b)=\langle e_{\mathcal{N}(s)}|a\rangle||b\rangle\rangle.$ (10)
For a bounded operator, symmetry implies self-adjointness.
(3) Boundary Conditions: By definition, $E_{0}=\mathrm{id}$ implies $e_{\mathcal{N}(0)}=\mathbf{1}$, and $E_{1}=E$ implies $e_{\mathcal{N}(1)}=e_{\mathcal{N}}$. ∎
While the operators $e_{\mathcal{N}(s)}$ established above possess desirable analytic properties, specifically boundedness and self-adjointness, the construction suffers from a fundamental algebraic defect. For intermediate values $s\in(0,1)$, the interpolated maps fail to satisfy the idempotency condition:
$E_{s}^{2}\neq E_{s}.$ (11)
Consequently, the image $E_{s}(\mathcal{M})$ does not constitute a von Neumann subalgebra, and thus Tomiyama’s theorem cannot be invoked. This structural failure is critical: without a subalgebra structure, the associated modular objects, such as the modular conjugation $J_{E_{s}(\mathcal{M})}$ required to generate the unitary flow, are ill-defined, rendering the path physically incomplete.
To illustrate this failure explicitly, we refer the reader to the $M_{2}(\mathbb{C})$ toy model in Appendix B. There, a direct computation verifies the breakdown of idempotency for the linear path, even in the simplest finite-dimensional Type I context. This counterexample demonstrates that the defect is not an artifact of the infinite-dimensional setting but is intrinsic to the linear interpolation strategy itself. This observation underscores the necessity for the strictly algebraic, modular-covariant approach—namely, the Haagerup–Kosaki construction (Path B)—which we develop in the following section.
3.3 Path B: The Canonical Haagerup–Kosaki Interpolation
As demonstrated by our analysis of Path A, a physically consistent interpolation must strictly preserve the von Neumann algebra structure throughout the process. This requirement necessitates a continuous family of maps satisfying dynamic idempotency, implying that every intermediate map must be a genuine conditional expectation. However, constructing such a trace-preserving family faces a fundamental obstruction due to the intrinsic nature of the physical algebra.
In the context of algebraic quantum field theory, the local operator algebras $\mathcal{M}$ and $\mathcal{N}$ (where $\mathcal{N}\subset\mathcal{M}$) are of Type III. A defining characteristic of Type III algebras is the absence of a normal, faithful, tracial state. Physically, this reflects the unbounded entanglement of the vacuum, which prevents the assignment of finite local density matrices or information measures. Mathematically, this precludes the existence of a trace-preserving conditional expectation $E:\mathcal{M}\rightarrow\mathcal{N}$. Instead, the relationship between the algebras is governed by a faithful normal operator-valued weight (OVW) $\mathcal{E}:\mathcal{M}^{+}\rightarrow\widehat{\mathcal{N}}^{+}$. This weight is unbounded and lacks the necessary properties to define a normalized projection. Here, $\mathcal{M}^{+}$ and $\mathcal{N}^{+}$ denote the positive cones of the respective algebras, while $\widehat{\mathcal{N}}^{+}$ denotes the extended positive cone of $\mathcal{N}$, a domain that mathematically accommodates the infinite values (UV divergences) inherent in the Type III setting.
To resolve this structural obstruction, we employ the Haagerup-Kosaki framework to lift the entire inclusion to a semifinite setting where a trace exists. We achieve this by constructing the crossed product with respect to a faithful normal semifinite weight $\omega$ on $\mathcal{M}$ chosen to be compatible with the inclusion. Compatibility requires that the modular automorphism group of the weight leaves the subalgebra invariant (i.e., $\sigma_{t}^{\omega}(\mathcal{N})=\mathcal{N}$, implying $\omega=\psi\circ\mathcal{E}$ for some weight $\psi$ on $\mathcal{N}$). Using the associated modular automorphism group $\sigma^{\omega}$, we construct the crossed-product envelopes:
$\tilde{\mathcal{M}}:=\mathcal{M}\rtimes_{\sigma^{\omega}}\mathbb{R},\quad\tilde{\mathcal{N}}:=\mathcal{N}\rtimes_{\sigma^{\omega}|_{\mathcal{N}}}\mathbb{R}.$ (12)
Throughout this paper, we employ the tilde notation (e.g., $\tilde{\mathcal{M}}$, $\tilde{E}$) to distinguish objects in this lifted semifinite-envelope algebra from their original Type III counterparts (e.g., $\mathcal{M}$, $\mathcal{E}$). In this lifted setting, both $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$ become Type II von Neumann algebras equipped with a canonical semifinite trace $\tau$. Within this tracial framework, the original unbounded operator-valued weight $\mathcal{E}$ is regularized into a genuine, trace-preserving conditional expectation $\tilde{E}:\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{N}}$. This transformation from a non-tracial to a tracial setting facilitates the subsequent construction of the physical idempotent path. The formal statement and necessary conditions for the existence of $\tilde{E}$ are detailed in Appendix C.
From a physical standpoint, this crossed-product construction serves as a formal regularization of entanglement resources. Conceptually, this construction introduces an auxiliary structure—often interpreted as a collective coordinate or an emergent degree of freedom—typically associated with the energy or the clock of an auxiliary observer Witten22 ; Chandrasekaran2022cip . This additional degree of freedom allows the infinite entanglement intrinsic to the Type III vacuum to be measured against the canonical trace $\tau$. In effect, the Haagerup–Kosaki lift accesses an entanglement reservoir implicit in the physical algebra, restructuring it into a form where a meaningful information flow can be defined. By doing so, it bridges the gap between the intractable Type III structure and the physically transparent Type II setting, where information recovery manifests as a smooth geometric process.
However, the existence of this lifted structure does not, by itself, guarantee a valid interpolation path. One might naively attempt to construct intermediate algebras via spectral projections (or spectral cuts) of the modular operator. While intuitively appealing, such ad hoc truncations generally fail to preserve the subalgebra structure and, more critically, exhibit pathological boundary behavior. Specifically, any finite spectral cut excludes the tails of the modular spectrum, leaving the algebra effectively open and unable to recover the identity in the $s\to 0$ limit. Consequently, no matter how the limit is taken, such paths cannot close the gap to the full algebra $\tilde{\mathcal{M}}$, resulting in a fundamental discontinuity at $s=0$.
We construct the continuous interpolation path by bridging these two lifted algebras, $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$, via the analytic structure of non-commutative $L^{p}$ spaces. The key mechanism is to identify the interpolation parameter $s\in[0,1]$ with the $L^{p}$-space parameter $p$ via the relation $s=1/p$. This parameterization physically governs the information resolution of the algebra along the path. Equivalently, by interpolating between the reference weight $\tilde{\phi}_{0}$ and its restriction $\tilde{\phi}_{1}=\tilde{\phi}_{0}\circ\tilde{E}$, we model the continuous loss of information. This transition follows the modular flow, ensuring that the non-commutative structure is preserved via analytic continuation, providing the unique rigidity required for our canonical path.
  • At $s=0$ ($p=\infty$): The system corresponds to the non-commutative $L^{\infty}$-space, which is the algebra itself. Thus, we retain the full information of the envelope algebra $\tilde{\mathcal{M}}$.
  • At $s=1$ ($p=1$): The system corresponds to the non-commutative $L^{1}$-space (the predual), restricted to the subalgebra. Here, the information is fully coarse-grained to the target algebra $\tilde{\mathcal{N}}$.
  • For $0<s<1$: The system possesses an intermediate information resolution, defined by the analytic interpolation between the full algebra $\tilde{\mathcal{M}}$ and the subalgebra $\tilde{\mathcal{N}}$.
Based on this structure, we define the interpolation path as follows.
Definition 3.4 (Canonical Interpolation Path).
Let $\tilde{\phi}_{0}$ be the faithful normal semifinite dual weight on the crossed product algebra $\tilde{\mathcal{M}}$, constructed from the physical weight $\omega$ on $\mathcal{M}$. This dual weight $\tilde{\phi}_{0}$ serves as the reference weight for the interpolation.
We define the endpoint weight $\tilde{\phi}_{1}:=\tilde{\phi}_{0}\circ\tilde{E}$, representing the reference weight coarse-grained to the target subalgebra $\tilde{\mathcal{N}}$. Mathematically, this definition lifts the compatibility condition of the original inclusion ($\omega=\psi\circ\mathcal{E}$) to the crossed product level. Crucially, whereas $\tilde{\phi}_{0}$ retains full information, $\tilde{\phi}_{1}$ serves as the target state that governs the trajectory of information loss.
The canonical interpolation is generated by the analytic continuation of the relative modular flow between these two weights. Specifically, the Connes–Takesaki Radon–Nikodym cocycle $D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}$ is generated by the non-commutative Radon–Nikodym derivative $h=d\tilde{\phi}_{1}/d\tilde{\phi}_{0}$. This positive self-adjoint operator induces the cocycle via the relation:
$D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}=h^{it}.$ (13)
We define the interpolated weight $\tilde{\phi}_{s}$ for $s\in[0,1]$ by analytically extending the modular time $t$ to imaginary time (corresponding to the Wick rotation $t\to-is$). This yields the specific density relation:
$\frac{d\tilde{\phi}_{s}}{d\tilde{\phi}_{0}}=h^{s}=\left(\frac{d\tilde{\phi}_{1}}{d\tilde{\phi}_{0}}\right)^{s}.$ (14)
Since the entire path is generated by the spectral resolution of the single operator $h$, all relative Radon-Nikodym derivatives commute. Consequently, the chain rule for the Connes-Takesaki cocycle simplifies to the algebraic sum of exponents:
$[D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}$ $=[D\tilde{\phi}_{1}:D\tilde{\phi}_{s}]_{t}[D\tilde{\phi}_{s}:D\tilde{\phi}_{0}]_{t}$
$=h^{i(1-s)t}h^{ist}=h^{it}~.$ (15)
Finally, the one-parameter family of algebras $\tilde{\mathcal{N}}(s)$ is defined as the range of the unique conditional expectations $\tilde{E}_{s}:\tilde{\mathcal{M}}\to\tilde{\mathcal{M}}$ compatible with these weights:
$\tilde{\mathcal{N}}(s):=\text{Range}(\tilde{E}_{s})=\{x\in\tilde{\mathcal{M}}\mid\tilde{E}_{s}(x)=x\}.$ (16)
This analytic structure guarantees that the path connects the boundaries smoothly and that the interpolated weights are uniquely determined by the non-commutative $L^{p}$-geometry (identifying $s=1/p$).
Heuristically, this construction signifies that $\tilde{\phi}_{s}$ acts as a non-commutative geometric mean between the reference weight $\tilde{\phi}_{0}$ and the coarse-grained weight $\tilde{\phi}_{1}$. Although formal products of weights are not defined in the operator algebra, we may conceptually visualize this interpolation as:
$\tilde{\phi}_{s}\sim\tilde{\phi}_{1}^{s}\tilde{\phi}_{0}^{1-s}~.$ (17)
This geometric averaging property ensures that the path $\tilde{\mathcal{N}}(s)$ follows the geodesic of the underlying modular structure, strictly minimizing the information distance (relative entropy) between the full algebra and the subalgebra at each step.
We now formulate the fundamental physical properties of this path, analogous to the requirements examined for Path A.
Theorem 3.5 (Properties of the Canonical Interpolation Path).
The family of algebras $\tilde{\mathcal{N}}(s)$ defined by the canonical $L^{p}$-interpolation satisfies the following physical and mathematical requirements:
  1. 1.
    Boundary Conditions: $\tilde{\mathcal{N}}(0)=\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}(1)=\tilde{\mathcal{N}}$. This ensures the path smoothly connects the full algebra to the target subalgebra.
  2. 2.
    Nesting (Monotonicity): For any $0\leq s\leq s^{\prime}\leq 1$, the inclusion relation $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ holds. This reflects the monotonic coarse-graining of information, structurally analogous to a renormalization group flow.
  3. 3.
    Dynamic Idempotency: The associated conditional expectations satisfy the consistency condition: $\tilde{E}_{s^{\prime}}\circ\tilde{E}_{s}=\tilde{E}_{s^{\prime}}$ for $s^{\prime}\geq s$.
Proof.
These properties are direct consequences of the analytic structure of the Haagerup-Kosaki construction. Unlike naive spectral truncations, this path is generated by the spectral resolution of the Radon-Nikodym derivative $h$, which ensures structural continuity via the modular automorphism group.
(1) Boundary Conditions: The path relies on the analytic continuation of the perturbed weights. At $s=0$, the generator $h^{0}=\mathbf{1}$ acts trivially, implying that the associated projection is the identity map ($\tilde{E}_{0}=\text{id}_{\tilde{\mathcal{M}}}$). Conversely, at $s=1$, the construction converges to the operator $h^{1}$ which generates the target conditional expectation $\tilde{E}$. Thus, $\tilde{E}_{1}$ recovers the original projection onto $\tilde{\mathcal{N}}$, eliminating the discontinuity at $s=0$ characteristic of spectral cut methods.
(2) Nesting: This property follows from the spectral ordering of the Radon-Nikodym derivative $h$. The parameter $s$ effectively controls the resolution of the algebra relative to the reference state. Since $h$ is a positive operator, the family of spectral projections associated with $h^{s}$ defines a monotonic hierarchy of subspaces. Mathematically, increasing $s$ corresponds to a stronger restriction on the algebra, strictly implying $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ for $s^{\prime}\geq s$.
(3) Idempotency: This follows immediately from the nesting property $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$. By the tower property of conditional expectations, the composition of mappings onto nested subalgebras reduces to the mapping onto the smaller subalgebra, yielding $\tilde{E}_{s^{\prime}}\circ\tilde{E}_{s}=\tilde{E}_{s^{\prime}}$. ∎
The canonical $L^{p}$-interpolation provides more than mathematical smoothness; it imposes a strict ordering on information. The nesting property $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ established in Theorem 3.5 identifies $s$ as a coarse-graining scale: increasing $s$ corresponds to systematically discarding information, strictly analogous to the irreversible flow of a renormalization group. This structure now enables us to construct the unitary operator $\tilde{U}(s):=J_{\tilde{\mathcal{M}}}J_{\tilde{\mathcal{N}}(s)}$, which serves as the lifted counterpart to the physical recovery path $U(s)$ originally envisioned in Section 3.1. While the original path $U(s)$ was ill-defined due to the Type III obstruction, this lifted operator $\tilde{U}(s)$ is rigorously defined within the crossed-product framework. It acts as an information recovery process, reversing the coarse-graining by translating information from the finer-grained algebra back into the decodable sector. Thus, this construction not only connects the two algebras but also establishes the rigorous framework required to define the unitary recovery path, which we explicitly construct in the following section.
3.4 The Unitary Path and its Generator
Leveraging the structural consistency established in Theorem 3.5, we now proceed to construct the unitary recovery operator. Since the image $\tilde{\mathcal{N}}(s)$ forms a valid von Neumann subalgebra for all $s\in[0,1]$, the associated modular objects are well-defined along the entire trajectory. This allows us to explicitly define the modular conjugation $J_{\tilde{\mathcal{N}}(s)}$ and, consequently, the strongly continuous unitary path:
$\tilde{U}(s):=J_{\tilde{\mathcal{M}}}J_{\tilde{\mathcal{N}}(s)}.$ (18)
This construction represents the continuous, field-theoretic counterpart to the discrete information recovery protocols discussed in vdHV and summarized in Sec. 2.2. Following the logic established in Step 3 of the protocol, the operator $\tilde{U}(s)$ acts as a dynamic teleportation map. At each instant $s$, the combined action of the modular conjugations relocates the information from the instantaneous relative commutant $\tilde{\mathcal{N}}(s)^{\prime}\cap\tilde{\mathcal{M}}$ to the dual relative commutant $\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$, within the extension defined by the Jones basic construction:
$\tilde{\mathcal{M}}_{1}(s):=\langle\tilde{\mathcal{M}},e_{\tilde{\mathcal{N}}(s)}\rangle.$ (19)
Physically, the flow generated by $\tilde{U}(s)$ explicitly implements this transport: it progressively decodes the information hidden in the relative commutant sector, mapping it back into the decodable sector, thereby effectively reversing the coarse-graining induced by the inclusion. Crucially, the analytic nature of the Haagerup–Kosaki interpolation guarantees the boundary conditions $\tilde{U}(0)=\mathbf{1}$ and $\tilde{U}(1)=\tilde{U}_{\Gamma}$, ensuring that the path continuously connects the identity to the target canonical shift.
The path $\tilde{U}(s)$ does not generally satisfy the group property $\tilde{U}(s+t)=\tilde{U}(s)\tilde{U}(t)$. Consequently, Stone’s Theorem does not directly apply to define a generator $\tilde{G}$ such that $\tilde{U}(s)=e^{-is\tilde{G}}$. Instead, we define $\tilde{G}$ as the infinitesimal tangent vector at the origin $s=0$. For this operator to be well-defined and physically meaningful, we must ensure that the path is strongly differentiable on a suitable domain and that $\tilde{G}$ is essentially self-adjoint. To establish these properties, we restrict our attention to a dense core of vectors $\mathcal{D}_{\text{core}}\subset\mathcal{H}$ that are analytic with respect to the modular flow $\sigma_{t}^{\tilde{\phi}_{0}}$. This choice of domain provides the necessary analytic control to prove the self-adjointness of the generator in Lemma 3.9.
Remark 3.6 (Modular Covariance and Core Preservation).
The structural consistency of this construction rests on the modular covariance of the conditional expectations $\tilde{E}_{s}$. Specifically, the fact that $\tilde{E}_{s}$ commutes with the modular automorphism group $\sigma_{t}^{\tilde{\phi}_{0}}$ (i.e., $\tilde{E}_{s}\circ\sigma_{t}^{\tilde{\phi}_{0}}=\sigma_{t}^{\tilde{\phi}_{0}}\circ\tilde{E}_{s}$) guarantees that analyticity is preserved under projection: elements that are analytic with respect to the modular flow are mapped to analytic elements within the subalgebra. More precisely, if $x\in\tilde{\mathcal{M}}$ admits an entire analytic continuation $t\mapsto\sigma_{t}^{\tilde{\phi}_{0}}(x)$, its image $\tilde{E}_{s}(x)$ admits the same entire extension. Consequently, the common analytic core $\mathcal{D}_{\text{core}}$ remains invariant under the interpolated expectations for all $s\in[0,1]$. This invariance provides the stable, dense domain required to rigorously define $\tilde{G}$ as a derivation and serves as the foundation for proving its self-adjointness in Lemma 3.9, and facilitates the broader analysis of its properties in Sec. 4.
The immediate physical consequence of this core preservation is that the constructed path is sufficiently smooth to define a generator. The invariant core $\mathcal{D}_{\text{core}}$ ensures regularity with respect to the modular flow, enabling well-defined differentiation at the boundary $s=0$.
Lemma 3.7 (Analyticity and Differentiability of the Path).
The map $s\mapsto\tilde{U}(s)$ constructed via the canonical $L^{p}$-interpolation is real-analytic for $s\in(0,1)$. Moreover, for vectors in the common modular core $\mathcal{D}_{\text{core}}$, the map is strongly differentiable at the boundary $s=0$.
Proof.
The proof relies on the analytic structure of the Haagerup–Kosaki interpolation and the properties of the modular domain.
  1. 1.
    Analytic Extension inside the Interval: The interpolated spaces $L^{p}(\tilde{\mathcal{M}},\tilde{\phi}_{s})$ are defined via complex interpolation in the strip $0<\text{Re}(z)<1$, where $z$ extends the real parameter $s$ into the complex plane. Consequently, the structural maps, including the modular conjugations $J_{\tilde{\mathcal{N}}(s)}$, depend analytically on the parameter $s$ within the open interval. This implies that for any vector $\xi\in\mathcal{D}_{\text{core}}$, the vector-valued function $s\mapsto\tilde{U}(s)\xi$ is real-analytic in $(0,1)$.
  2. 2.
    Differentiability at the Boundary: The differentiability at $s=0$ hinges on the invariance of the analytic core established in Remark 3.6. Recall that the core $\mathcal{D}_{\text{core}}$ consists of vectors $\xi$ for which the map $t\mapsto\sigma_{t}^{\tilde{\phi}_{0}}(\xi)$ extends to an entire analytic function. As established in Definition 3.4, the parameter $s$ governs the non-commutative $L^{p}$-resolution via $s=1/p$, a structure intrinsically defined by the analytic continuation of the modular flow to the imaginary axis (Wick rotation $t\to-is$). Consequently, the derivative with respect to $s$ at the boundary ($s=0$, corresponding to $p=\infty$) is directly generated by the modular Hamiltonian. The invariance of $\mathcal{D}_{\text{core}}$ ensures that this analytic structure is preserved along the path, guaranteeing that the strong limit defining the derivative exists for all vectors within this core.
Definition 3.8 (The Generator $\tilde{G}$).
Motivated by the differentiability established in Lemma 3.7, we define the operator $\tilde{G}$ on the common modular core $\mathcal{D}_{\text{core}}$ as the infinitesimal generator at the origin:
$\tilde{G}:=i\frac{d\tilde{U}(s)}{ds}\bigg|_{s=0}=iJ_{\tilde{\mathcal{M}}}\frac{dJ_{\tilde{\mathcal{N}}(s)}}{ds}\bigg|_{s=0}.$ (20)
Finally, we establish the physical validity of this operator.
Lemma 3.9 (Essential Self-Adjointness of $\tilde{G}$).
The operator $\tilde{G}$ defined on $\mathcal{D}_{\text{core}}$ is essentially self-adjoint.
Proof.
We invoke Nelson’s analytic vector theorem Nelson59 , utilizing the properties of the modular core $\mathcal{D}_{\text{core}}$.
  1. 1.
    Symmetry: Since $\tilde{U}(s)$ is unitary for real $s$, differentiating the identity $\tilde{U}(s)^{*}\tilde{U}(s)=\mathbf{1}$ at $s=0$ yields $\tilde{U}^{\prime}(0)^{*}\tilde{U}(0)+\tilde{U}(0)^{*}\tilde{U}^{\prime}(0)=0$. Using the boundary condition $\tilde{U}(0)=\mathbf{1}$, this implies $\tilde{U}^{\prime}(0)^{*}=-\tilde{U}^{\prime}(0)$. Thus, $\tilde{G}=i\tilde{U}^{\prime}(0)$ is a symmetric operator ($\tilde{G}^{*}=\tilde{G}$) on the dense domain $\mathcal{D}_{\text{core}}$.
  2. 2.
    Analytic Vectors: The generator $\tilde{G}$ is constructed from the analytic continuation of the Radon-Nikodym cocycle, which is generated by the logarithm of the relative modular operator $h$. The domain $\mathcal{D}_{\text{core}}$ is explicitly defined as the set of vectors that are analytic with respect to the modular flow $\sigma_{t}^{\tilde{\phi}_{0}}$ generated by $\tilde{\phi}_{0}$ (and consequently the unitary group $h^{it}$). According to the general theory of one-parameter unitary groups, the set of analytic vectors for the group $h^{it}$ constitutes a dense set of analytic vectors for its generator $\log h$. Since the action of $\tilde{G}$ at $s=0$ is linearly related to this generator, the set $\mathcal{D}_{\text{core}}$ forms a dense set of analytic vectors for $\tilde{G}$ as well.
  3. 3.
    Conclusion: By Nelson’s analytic vector theorem, a symmetric operator that admits a dense set of analytic vectors is essentially self-adjoint. Consequently, the closure $\overline{\tilde{G}}$ is self-adjoint and unique. This guarantees that $\tilde{G}$ unambiguously defines a valid quantum observable, providing the rigorous basis for the infinitesimal generation of the recovery flow.
Remark 3.10 (Uniqueness of the Path).
The canonical $L^{p}$-interpolation path employed here is defined without ambiguity. It is uniquely determined by the initial reference weight and the target subalgebra, independent of arbitrary choices such as basis vectors or auxiliary cutoffs. Consequently, $\tilde{G}$ is a canonical object of the theory, invariant under unitary conjugations that preserve the inclusion structure. This establishes $\tilde{G}$ as the well-defined, self-adjoint generator of the continuous teleportation protocol.
Remark 3.11 (Canonical Nature and Covariance).
The construction of the generator $\tilde{G}$ is free from the ambiguities typically associated with cutoff-based regularization schemes. The canonical $L^{p}$-interpolation path is uniquely determined solely by the initial reference weight $\tilde{\phi}_{0}$ and the algebraic inclusion $\tilde{\mathcal{N}}\subset\tilde{\mathcal{M}}$, independent of arbitrary choices such as basis vectors or auxiliary projections. Consequently, $\tilde{G}$ is not merely a mathematical artifact but a canonical object of the theory: it acts as an intrinsic physical observable associated with the inclusion pair, transforming covariantly under automorphisms that preserve this structure. This establishes $\tilde{G}$ as the rigorously defined, self-adjoint generator of the continuous teleportation protocol.
4 DERIVATION OF THE OPERATOR IDENTITY $\tilde{G}=2P$
With the self-adjoint generator $\tilde{G}$ rigorously defined in Sec. 3.4, we now address the central objective of this work: identifying the teleportation generator with the modular momentum. This section presents a derivation of the identity $\tilde{G}=2P$ for general Type III algebras. We begin by motivating the specific form of the identity, particularly the factor of 2, through the limit of HSMI (Sec. 4.1). We then employ modular perturbation theory to prove the identity as an exact relation between closed operators (Sec. 4.2). Finally, we examine the geometric stability of this result within the framework of non-commutative geometry (Sec. 4.3) and propose a holographic correlation function test to verify its physical validity (Sec. 4.4).
4.1 Motivation from Half-Sided Modular Inclusions
A strong theoretical foundation for identifying the teleportation generator with modular momentum, as proposed by vdH-V vdHV , is found in highly symmetric settings, exemplified by HSMI. In this context, the connection between the canonical shift and modular operators is not merely a hypothesis but a rigorous theorem established by Borchers and Wiesbrock Borchers92 ; Wiesbrock93 .
As discussed in vdHV , for an HSMI $\mathcal{N}\subset\mathcal{M}$, the discrete canonical shift unitary $U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}$ is generated by exactly twice the spacetime translation operator $P$ corresponding to the difference of modular Hamiltonians, $K_{\mathcal{M}}-K_{\mathcal{N}}$.
$U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}=e^{-2iP}.$ (21)
Our canonical interpolation provides a dynamic generalization of this structure. Recall that we constructed a differentiable unitary path $\tilde{U}(s)$ satisfying the boundary conditions $\tilde{U}(0)=\mathbf{1}$ and $\tilde{U}(1)=U_{\tilde{\Gamma}}$. While this path does not generally form a one-parameter group in generic Type III settings, in the strict HSMI limit, the path is known to take the exact exponential form $\tilde{U}(s)=e^{-2isP}$. Consequently, applying the definition of the generator $\tilde{G}:=i\tilde{U}^{\prime}(0)$ given in Eq. (20) directly yields:
$\tilde{G}=i\frac{d}{ds}\left(e^{-2isP}\right)\bigg|_{s=0}=2P.$ (22)
This derivation provides a geometric motivation for the factor of 2. Physically, $\tilde{U}(s)=J_{\tilde{\mathcal{M}}}J_{\tilde{\mathcal{N}}(s)}$ represents the composition of a fixed modular reflection ($J_{\tilde{\mathcal{M}}}$) and one shifted by the parameter $s$ ($J_{\tilde{\mathcal{N}}(s)}$). Analogous to Euclidean geometry, where composing two reflections separated by a distance $d$ generates a translation of $2d$ (which corresponds to the full teleportation limit at $s=1$), this composite action induces a shift of twice the modular parameter. This geometric kinematics directly implies the generator relation $\tilde{G}=2P$.
4.2 Geometric Derivation via Canonical Path
Our objective is to establish the identity $\tilde{G}=2P$ in general Type III settings, moving beyond the specific symmetry of the HSMI case. Since both $\tilde{G}$ and $2P$ are unbounded self-adjoint operators, a precise proof requires demonstrating that they coincide as closed operators on a common core. We achieve this by employing the modular perturbation theory of Araki, Connes, and Kosaki (see, e.g., Araki:1976zv ; Connes80 ; Kosaki84 ; Kosaki84UC ) and invoking Nelson’s analytic vector theorem Nelson59 to ensure essential self-adjointness.
First, we determine the first-order behavior of the modular Hamiltonian along the canonical interpolation path.
Lemma 4.1 (First-Order Variation of the Modular Hamiltonian).
For the canonical interpolation path $\tilde{\mathcal{N}}(s)$ defined in Definition 3.4, the modular Hamiltonian $K(s)=K_{\tilde{\mathcal{N}}(s)}$ satisfies the following perturbation relation at the origin $s=0$:
$K^{\prime}(0)\;:=\;\frac{dK(s)}{ds}\Big|_{s=0}\;=\;K_{\tilde{\mathcal{N}}}\,-\,K_{\tilde{\mathcal{M}}}\;=\;-\,P~,$ (23)
where $P:=K_{\tilde{\mathcal{M}}}-K_{\tilde{\mathcal{N}}}$ is the generalized modular momentum.
Proof.
We utilize the structural definitions established in Definition 3.4. Recall that the interpolation path is generated by the analytic continuation of the Connes-Takesaki cocycle associated with the Radon-Nikodym derivative $h=d\tilde{\phi}_{1}/d\tilde{\phi}_{0}$. The interpolated cocycle satisfies the scaling relation:
$D\tilde{\phi}_{s}:D\tilde{\phi}_{0}]_{t}=h^{ist}~.$ (24)
The Connes cocycle intertwines the modular operators via the identity
$\Delta_{s}^{it}=[D\tilde{\phi}_{s}:D\tilde{\phi}_{0}]_{t}\Delta_{0}^{it}~,$ (25)
where $\Delta_{s}=e^{-K(s)}$ and $\Delta_{0}=e^{-K_{\tilde{\mathcal{M}}}}$. Combining this with Eq. (24), we express the modular operator along the path directly in terms of the generator $h$:
$\Delta_{s}^{it}=h^{ist}\Delta_{0}^{it},\qquad\forall\,t\in\mathbb{R}~.$ (26)
We now compute the derivative with respect to $s$ at $s=0$. Differentiating the right-hand side of (26) yields:
$\frac{d}{ds}\left(h^{ist}\Delta_{0}^{it}\right)\Big|_{s=0}=it(\ln h)\Delta_{0}^{it}~.$ (27)
To identify the operator $\ln h$, we examine the generator of the cocycle at the endpoint $s=1$. Differentiating the relation $h^{it}=\Delta_{1}^{it}\Delta_{0}^{-it}$ with respect to $t$ at $t=0$ yields $i\ln h$ on the left-hand side. On the right-hand side, the derivative gives $(-iK_{\tilde{\mathcal{N}}})-(-iK_{\tilde{\mathcal{M}}})=i(K_{\tilde{\mathcal{M}}}-K_{\tilde{\mathcal{N}}})=iP$. Comparing these generators establishes the identity $\ln h=P$. Consequently, the derivative in (27) simplifies to $itP\Delta_{0}^{it}$.
Alternatively, we evaluate the derivative of the left-hand side of (26), explicitly writing $\Delta_{s}^{it}=e^{-itK(s)}$. Applying Duhamel’s formula for the derivative of an exponential operator Kato , a cornerstone of Araki’s modular perturbation theory Araki:1976zv , we obtain:
$\frac{d}{ds}e^{-itK(s)}\Big|_{s=0}\;=\;-i\int_{0}^{t}e^{-i(t-\tau)K_{0}}K^{\prime}(0)e^{-i\tau K_{0}}d\tau~.$ (28)
To extract the generator $K^{\prime}(0)$, we divide both sides by $t$ and take the limit $t\to 0$. In this limit, the integral term converges to $iK^{\prime}(0)$ due to the strong continuity of the modular flow. Specifically, continuity ensures that $e^{-i\tau K_{0}}\to\mathbf{1}$ strongly as $\tau\to 0$, rendering non-commutative effects negligible at the leading order. On the other hand, for the cocycle term (the right-hand side), we have $\lim_{t\to 0}\frac{1}{t}(itP\Delta_{0}^{it})=iP$ (since $\Delta_{0}^{it}\to\mathbf{1}$). Comparing these limits establishes the identity:
$iK^{\prime}(0)=iP\quad\implies\quad K^{\prime}(0)=-P~.$ (29)
This confirms that the infinitesimal variation of the modular Hamiltonian corresponds exactly to the negative of the modular momentum. ∎
The result $K^{\prime}(0)=-P$ implies that the infinitesimal change of the modular Hamiltonian with respect to $s$ is governed by the spacetime translation operator $P$. As the interpolation parameter $s$ varies, the evolution of $K(s)$ is effectively driven by a geometric translation within the emergent spacetime. This confirms that our canonical interpolation path aligns with a distinct physical flow: at the boundary $s=0$, the infinitesimal change of the modular Hamiltonian is precisely given by P-P, indicating a trajectory directed towards the subalgebra (opposite to the canonical outward growth). Consequently, the path of conditional expectations $\tilde{E}_{s}$ is not merely a mathematical abstraction of subalgebra inclusion; its tangent direction manifests as a concrete physical translation. This substantiates the interpretation of the channel as a geometric transport of information, providing the dynamical basis for the recovery protocol.
Equipped with the result $K^{\prime}(0)=-P$ from Lemma 4.1, we can now explicitly compute $\tilde{G}$ and establish the main theorem.
Theorem 4.2 (The Identity $\tilde{G}=2P$).
The generator $\tilde{G}$ is identical to $2P$ as a closed self-adjoint operator.
Proof.
We derive the generator explicitly by utilizing the Type II nature of the crossed product algebra $\tilde{\mathcal{M}}$. Recall from Definition 3.8 that the unitary path is defined as $\tilde{U}(s)=J(0)J(s)$, where $J(0):=J_{\tilde{\mathcal{M}}}$ is the fixed reference modular conjugation, and $J(s):=J_{\tilde{\mathcal{N}}(s)}$ is the conjugation associated with the interpolating subalgebra $\tilde{\mathcal{N}}(s)$. The generator is given by
$\tilde{G}=iJ(0)\frac{dJ(s)}{ds}\Big|_{s=0}~.$ (30)
In the crossed product construction, the dual weights $\tilde{\phi}_{s}$ act as weights on a Type II von Neumann algebra. Consequently, they can be represented via the canonical trace $\tau$ and the Radon-Nikodym derivative $h$. Specifically, the density operator for the weight $\tilde{\phi}_{s}$ is $h^{s}$ (relative to the reference weight $\tilde{\phi}_{0}$ corresponding to $h^{0}=\mathbf{1}$). According to the standard transformation rules for modular conjugations in Type II settings, the conjugation $J(s)$ associated with the weight $\tilde{\phi}_{s}(\cdot)=\tau(h^{s}\,\cdot\,)$ is related to the reference conjugation $J(0)$ by the unitary transformation generated by the density operator
$J(s)=h^{is}J(0)h^{-is}~.$ (31)
From Lemma 4.1, we identified the logarithm of the Radon-Nikodym derivative as the modular momentum: $\ln h=P$. Substituting this into (31), we obtain the explicit time-evolution of the modular conjugation
$J(s)=e^{isP}J(0)e^{-isP}~.$ (32)
We now differentiate this expression at $s=0$ to find $J^{\prime}(0)$:
$J^{\prime}(0)$ $=\frac{d}{ds}\left(e^{isP}J(0)e^{-isP}\right)\Big|_{s=0}$
$=(iP)J(0)+J(0)(-iP)~.$ (33)
Substituting $J^{\prime}(0)$ back into the definition of $\tilde{G}$ in (30) yields
$\tilde{G}=-J(0)PJ(0)+J(0)^{2}P~.$ (34)
This expression simplifies immediately by invoking the structural properties of the modular conjugation. Specifically, $J(0)$ is an involution ($J(0)^{2}=\mathbf{1}$) and acts as an anti-unitary reflection on the modular Hamiltonian, implying the CPT inversion symmetry J(0)PJ(0)=PJ(0)PJ(0)=-P. Applying these identities, we obtain
$\tilde{G}=-(-P)+\mathbf{1}\cdot P=2P~.$ (35)
This algebraic derivation confirms the identity $\tilde{G}=2P$ on the common core $\mathcal{D}_{\text{core}}$. Since $\mathcal{D}_{\text{core}}$ is a core for both $\tilde{G}$ (by Nelson’s theorem) and $P$ (by modular covariance), the identity holds for their unique self-adjoint closures. ∎
4.3 Stability Quantification via Non-Commutative $L^{p}$ Spaces
A central physical question is whether the identity $\tilde{G}=2P$ represents a singular coincidence valid only at the boundary $s=0$, or a structurally robust property of the interpolation. Specifically, we must ensure that the generator does not exhibit large or uncontrolled fluctuations under small perturbations of the interpolation parameter $s$. We address this by employing the theory of non-commutative $L^{p}$ spaces Kosaki84 ; PisierXu , which provides the canonical geometric framework for analyzing perturbations in the Type II crossed product algebra.
Theorem 4.3 (Local Stability of the Generator).
The generator $\tilde{G}$ exhibits structural stability in the Kato-type strong resolvent sense. Let $\tilde{G}(s)$ denote the instantaneous generator of the unitary path at parameter $s$. For any fixed complex number $z$ in the resolvent set $\rho(2P)$ (defined as the complement of the spectrum $\sigma(2P)$ in the complex plane) and any state vector $\xi$ within the dense core domain $\mathcal{D}_{\text{core}}$, the deviation of the resolvent from the modular momentum limit is linearly bounded by the perturbative interpolating parameter $s$:
[…]
Here, $R_{T}(z)=(T-z)^{-1}$ denotes the resolvent operator, and $||\cdot||$ represents the vector norm on the Hilbert space. The coefficient $C_{z}$ is a stability factor that scales inversely with the distance to the spectrum $\sigma(2P)$ (i.e., $C_{z}\sim\operatorname{dist}(z,\sigma(2P))^{-1}$), indicating that stability is most sensitive near the spectral values. This inequality implies that $\tilde{G}(s)$ converges to $2P$ in the strong resolvent sense as $s\to 0$. Crucially, by the Trotter-Kato theorem, this guarantees the convergence of the generated unitary groups, ensuring that the physical dynamics of $\tilde{G}(s)$ smoothly approach those of $2P$ (strong convergence of dynamics).
Proof.
The proof relies on the analytic rigidity of the canonical interpolation path within the Type II setting, utilizing the specific geometry of non-commutative $L^{p}$ spaces.
  1. 1.
    Geometric Foundation (Uniform Convexity $\to$ Analyticity): The interpolation path is constructed within non-commutative $L^{p}$ spaces ($1<p<\infty$) associated with the crossed product algebra. A crucial feature of these spaces is their uniform convexity Kosaki84 . Geometrically, uniform convexity implies that the unit ball is strictly rotund without flat regions. In the context of complex interpolation, this geometric rigidity ensures that the map $s\mapsto\tilde{U}(s)$ is not merely continuous but real-analytic for $s\in(0,1)$ and differentiable at the boundary $s=0$.
  2. 2.
    From Differentiability to Linear Bound: The analyticity established above guarantees that the generator $\tilde{G}(s)$ is a differentiable function of $s$ at the origin. By the mean value theorem for operator-valued functions, the deviation of the generator from its limit is bounded linearly by the parameter $s$:
[…]
    for some constant $C$ related to the derivative $\tilde{G}^{\prime}(0)$. This linear bound ($\mathcal{O}(|s|)$) is a direct consequence of the smoothness of the geometry; without uniform convexity, the path could be fractal or non-differentiable, violating this bound.
  3. 3.
    Resolvent Estimate: To translate the generator bound to the resolvent bound, we employ the second resolvent identity:
    $R_{\tilde{G}(s)}(z)-R_{2P}(z)=R_{\tilde{G}(s)}(z)\big(2P-\tilde{G}(s)\big)R_{2P}(z)~.$ (38)
    Taking the norm and applying the linear bound from the previous step yields:
[…]
    where $C_{z}^{\prime}$ represents the uniform bound on the operator norm of the perturbed resolvent $R_{\tilde{G}(s)}(z)||_{\textrm{op}}$ for small $s$. The composite stability factor $C_{z}$ encapsulates the norms of the resolvents (scaling as $\text{dist}(z,\sigma)^{-1}$) along with the geometric derivative constant $C$.
This derivation confirms that the structural robustness ($\tilde{G}=2P$) is not accidental but grounded in the uniform convexity of the underlying modular geometry, which enforces the linear control of fluctuations. ∎
4.4 Correlation-Function Test of the Conjecture
The operator identity $\tilde{G}=2P$ posits a direct equivalence between the algebraic shift derived from modular theory and the geometric translation of the emergent spacetime. While operator-level identities are mathematically rigid, their physical validity in complex quantum gravity models requires verification through observables. We propose a specific test using two-point correlation functions within the framework of the AdS/CFT correspondence.
We frame this test in the context of a two-sided eternal black hole, holographically dual to the thermofield double state $|\text{TFD}\rangle$. To precisely model the information transport, we adopt the algebraic setup relevant to the black hole information paradox. We identify $\tilde{\mathcal{N}}$ with the algebra of the old black hole and $\tilde{\mathcal{M}}$ with the enlarged algebra that includes the entire black hole plus an infalling message (Alice’s diary). In this construction, the relative commutant $\tilde{\mathcal{N}}^{\prime}\cap\tilde{\mathcal{M}}$ represents the specific information of Alice’s diary. Technically, the canonical shift operator $\tilde{U}(s)$ is constructed to transport this information from $\tilde{\mathcal{N}}(s)^{\prime}\cap\tilde{\mathcal{M}}$ to the dual relative commutant $\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$ (the radiation sector). According to our main theorem ($\tilde{G}=2P$), the infinitesimal generator responsible for moving Alice’s diary into the radiation sector coincides exactly with twice the generator of the geometric modular flow (the boost).
Let $O_{L}$ and $O_{R}$ be local probe operators acting on the Left and Right asymptotic boundaries, respectively, where $O_{R}$ is associated with the infalling message. Since the generator $\tilde{G}$ acts as a geometric boost generator on the Right wedge, the corresponding unitary $\tilde{U}(s)$ induces a non-trivial coordinate shift on the probe $O_{R}$ for small $s$, effectively displacing its position relative to the horizon, while leaving the Left operator $O_{L}$ invariant.
To probe whether this algebraic action corresponds to a geometric translation, we define a correlation function $F(s)$ that measures how the Left-Right entanglement implies a correlation change under the algebraic shift of the Right operator:
$F(s)\;:=\;\langle\text{TFD}|\,O_{L}\,\left(\tilde{U}(s)O_{R}\tilde{U}(s)^{*}\right)\,|\text{TFD}\rangle.$ (40)
This correlator probes the geodesic distance through the wormhole between the fixed Left operator and the algebraically shifted Right operator. Since $\tilde{G}$ is defined as the infinitesimal generator of the path at $s=0$ (satisfying the expansion $\tilde{U}(s)=\mathbf{1}-is\tilde{G}+O(s^{2})$), the initial response determines the prediction for the operator identity. If $\tilde{G}=2P$ holds, we obtain
$F^{\prime}(0)\;=\;\langle\text{TFD}|\,O_{L}\,(-i[\tilde{G},O_{R}])\,|\text{TFD}\rangle\;=\;-2i\langle\text{TFD}|\,O_{L}\,[P,O_{R}]\,|\text{TFD}\rangle.$ (41)
Eq. (41) provides a precise physical interpretation: the algebraic generator $\tilde{G}$, originally defined to extract Alice’s diary from the black hole, effectively induces a translation on the boundary operator $O_{R}$ with exactly twice the magnitude of the geometric generator $P$.
In the bulk gravity description, $P$ acts as the generator of a geometric boost. Crucially, this boost shifts the null coordinate $v$, effectively displacing the horizon inwards relative to the probe. This implies that a region previously hidden in the interior becomes accessible from the exterior. Therefore, verifying Eq. (41) involves checking whether the algebraic shift $\tilde{U}$ displaces the coordinate position of the message $O_{R}$ inward across the original horizon location by a coordinate distance twice the corresponding geometric translation.
This geometric action is deeply connected to the mechanics of holographic teleportation. In models like the Gao-Jafferis-Wall (GJW) protocol Gao2017 , traversability is achieved by a shockwave that shifts the horizon. Our result implies that the algebraic generator $\tilde{G}$ mimics this horizon shift intrinsically. Rather than requiring an external matter source, the algebraic flow $\tilde{U}(s)$ induces a frame transformation equivalent to the geometric displacement required to bridge the entanglement wedge and recover the information. In semiclassical models like Jackiw-Teitelboim (JT) gravity Saad:2019lba , the quantity $i\langle O_{L}[P,O_{R}]\rangle$ relates to the Shapiro time delay for signals crossing the wormhole Shenker2014 ; Maldacena2016 . Thus, the factor of 2 confirms that $\tilde{U}(s)$ physically manipulates the geometry of the stretched horizon, facilitating the information transfer.
5 Conclusion
In this work, we have presented a continuous canonical path connecting the discrete, algebraic teleportation protocol of vdH-V vdHV within the framework of algebraic quantum field theory characterized by Type III von Neumann algebras. A defining feature of Type III algebras is the absence of a normal, faithful, tracial state, which precludes the existence of a trace-preserving conditional expectation $E:\mathcal{M}\rightarrow\mathcal{N}$. To overcome this structural obstacle, we utilized the crossed product construction to lift the Type III inclusion $\mathcal{N}\subset\mathcal{M}$ to the semifinite enveloping algebras $\tilde{\mathcal{N}}\subset\tilde{\mathcal{M}}$.
Leveraging Haagerup-Kosaki non-commutative $L^{p}$-space theory, we have constructed the unitary canonical shift operator $\tilde{U}(s)$ for $s\in[0,1]$. This continuous unitary path describes the interpolated transport of quantum information: as $s$ increases, the operator smoothly transfers the information residing in the relative commutant $\tilde{\mathcal{N}}(s)^{\prime}\cap\tilde{\mathcal{M}}$ into the accessible radiation sector $\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$ within the extension $\tilde{\mathcal{M}}_{1}(s):=\langle\tilde{\mathcal{M}},e_{\tilde{\mathcal{N}}(s)}\rangle$.
Our construction yields a strongly continuous unitary path $\tilde{U}(s)$ generated by a well-defined self-adjoint operator $\tilde{G}$, satisfying the relation
$\tilde{G}\;=\;2\big(K_{\tilde{\mathcal{M}}}-K_{\tilde{\mathcal{N}}}\big)\;=\;2P~.$ (42)
Crucially, the factor of $2$ in this equation arises because the generator combines two modular conjugations; consequently, the effective displacement of the information corresponds to twice the geometric distance associated with a single modular reflection. In this view, information recovery via teleportation becomes algebraically equivalent to a geometric spacetime translation generated by $P$, providing a precise formulation of the correspondence Teleportation = Translation.
This perspective offers a consistent framework for approaching the black hole information paradox. Within this picture, the recovery process is unitary: the operator $\tilde{U}(s)$ transports interior information to the exterior algebra rather than destroying it, suggesting that no information is lost during the evaporation process. No-cloning is naturally preserved by algebraic complementarity, as the information falling into the black hole is rapidly scrambled and subsequently teleported into the accessible radiation sector without duplication. The apparent thermality of black hole radiation is interpreted as a consequence of the Type III algebraic structure—specifically the absence of a tracial state—rather than an intrinsic feature of the global dynamics. This structural feature is directly tied to the presence of the horizon, which enforces the locality of QFT. Furthermore, lifting the description to the semifinite envelope unveils the algebraic structure underlying the Type III algebra, thereby clarifying the modular flow of information via the tracial structure of the extended Type II system.
It is important to note that these results are derived within an algebraic QFT framework on a fixed background spacetime. This implies that the essential mechanism for resolving the black hole information paradox is accessible within the semiclassical approximation. We have demonstrated that, even in this semiclassical regime, the information flow remains unitary and consistent with geometric translation. On the other hand, the full quantum gravitational dynamics—particularly the back-reaction of the evaporation process on the geometry and the final stage of black hole evaporation—remain a critical frontier.
Our work suggests several directions for future research. First, the connection between $\tilde{G}=2P$ and traversable wormholes warrants deeper investigation. Unlike the GJW protocol Gao2017 , which requires an external source (such as a shockwave) to shift the horizon, our continuous protocol generates a horizon shift intrinsically through modular interpolation. However, since both mechanisms rely on the generator $P$ to translate information across the horizon, incorporating additional algebraic conditions—such as specific coupling deformations—into our framework may provide a purely algebraic description of the GJW protocol. This would offer a non-perturbative perspective on the wormhole opening mechanism via modular flow.
Second, verifying the correlation function relation $F^{\prime}(0)\propto-i\langle[P,\mathcal{O}]\rangle$ in holographic models, such as double-scaled Sachdev-Ye-Kitaev (DSSYK) Berkooz:2018jqr or JT gravity, would help bridge our algebraic results with semiclassical gravity calculations.
Finally, extending this framework to incorporate gravitational back-reaction represents a critical next step. Since our protocol $\tilde{U}(s)$ is explicitly constructed within the crossed-product envelope—a structure known to capture $1/N$ corrections and observer energy constraints Witten22 ; Chandrasekaran2022cip ; LeutheusserLiu_ET —it offers a natural starting point for such an extension. Formulating the interplay between the information transport $\tilde{U}(s)$ and the semi-classical Einstein equations in this Type II setting could pave the way for a dynamical theory of emergent spacetime geometry arising from quantum entanglement.
Acknowledgements.
I acknowledge the assistance of Large Language Models, specifically Gemini and Chatgpt, in drafting the text, providing valuable feedback, and identifying errors. I take full intellectual responsibility for the content of this work. This work is supported by the National Research Foundation of Korea(NRF) grant with grant number RS-2019-NR040081, RS-2023-00208047, RS-2025-00553127. APPENDIX:
Appendix A Notation table
Symbol Meaning
$\mathcal{M},\ \mathcal{N}$ von Neumann algebras ($\mathcal{N}\subset\mathcal{M}$)
$\tilde{\mathcal{M}},\ \tilde{\mathcal{N}}$ semifinite crossed-product algebras (lifted, enveloped algebras)
$\quad\mathcal{X}^{\prime}$ commutant of algebra $\mathcal{X}$
$\mathcal{X}^{+},\ \widehat{\mathcal{X}}^{+}$ positive and extended positive cones of $\mathcal{X}$
$\quad s,\ p$ interpolated and $L^{p}$-space parameters
$L^{p}(\mathcal{X})$ non-commutative $L^{p}$ spaces associated with $\mathcal{X}$
[…]
$\mathcal{E},\ T$ operator-valued weight (OVW) and dual OVW
$\tilde{\mathcal{M}}_{1}(s),\ \tilde{\mathcal{N}}(s)$ basic extension algebra $\langle\tilde{\mathcal{M}},e_{\tilde{\mathcal{N}}(s)}\rangle$, interpolated subalgebra
$e_{\tilde{\mathcal{N}}},\ e_{\tilde{\mathcal{N}}(s)}$ Jones projection and interpolated projections onto subalgebras
$U_{\Gamma},\ \tilde{U}_{\Gamma},\ \tilde{U}(s)$ discrete and lifted canonical shift unitaries
$D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}$ Connes-Takesaki Radon-Nikodym cocycle
$h=d\tilde{\phi}_{1}/d\tilde{\phi}_{0}$ Radon–Nikodym derivative
$\quad\tilde{G}$ self-adjoint generator
$K_{\mathcal{X}},\ P$ modular Hamiltonian and modular momentum
$\quad\sigma_{t}^{\phi}$ modular automorphism group of weight $\phi$
$J_{\mathcal{X}},\ \Delta_{\mathcal{X}}$ modular conjugation and modular operator
$\quad\tau$ canonical trace on the crossed product $\tilde{\mathcal{M}}$
$\tilde{\phi}_{0},\ \tilde{\phi}_{1}$ faithful normal weights on $\tilde{\mathcal{M}}$
Table 1: Quick reference for frequently used symbols.
Appendix B Finite-Dimensional Model: Idempotency Failure and Unitary Interpolation
This appendix details the explicit calculation demonstrating the failure of the naive CP-map interpolation (Path A) to satisfy idempotency in finite dimensions, and provides a simple example of the required continuous unitary path.
Example B.1 (Failure of Path A in $M_{2}(\mathbb{C})$).
Consider $\mathcal{M}=\mathrm{M}_{2}(\mathbb{C})$ and let $\mathcal{N}=\mathbb{C}\mathbf{1}$ be the subalgebra of scalar matrices. The unique trace-preserving conditional expectation $E:\mathcal{M}\rightarrow\mathcal{N}$ is the normalized trace, $E(A)=\frac{1}{2}\mathrm{Tr}(A)\mathbf{1}$. The identity map is $\mathrm{id}(A)=A$. Using the convex combination (Path A), the interpolated map is:
$E_{s}(A)=(1-s)A+\frac{s}{2}\mathrm{Tr}(A)\mathbf{1}$ (43)
At the midpoint $s=0.5$, the map becomes $E_{0.5}(A)=0.5A+0.25\mathrm{Tr}(A)\mathbf{1}$. Applying the map again to test idempotency, and noting that $E_{0.5}(\mathbf{1})=\mathbf{1}$, we obtain:
$E_{0.5}^{2}(A)=E_{0.5}(0.5A+0.25\mathrm{Tr}(A)\mathbf{1})=0.25A+0.375\mathrm{Tr}(A)\mathbf{1}.$ (44)
Since $E_{0.5}^{2}(A)\neq E_{0.5}(A)$, this demonstrates the breakdown of the idempotency property for the naive path. This failure illustrates why the more sophisticated Haagerup–Kosaki lift is required to preserve algebraic structure in the Type III setting.
Example B.2 (Continuous Unitary Interpolation Goal).
Consider the same toy model $\mathcal{M}=\mathrm{M}_{2}(\mathbb{C})$. Define $U_{\Gamma}$ to be the canonical shift unitary that swaps the two basis states (analogous to the Pauli X operator), $U_{\Gamma}|0\rangle=|1\rangle$, $U_{\Gamma}|1\rangle=|0\rangle$. This discrete teleportation unitary can be continuously approached by a path of rotations. We construct a continuous path $U(s)=e^{-isG}$ using the self-adjoint generator:
$G=\frac{\pi}{2}\sigma_{x}=\frac{\pi}{2}(|0\rangle\langle 1|+|1\rangle\langle 0|)$ (45)
such that $U(0)=\mathbf{1}$. At $s=1$, we find:
$U(1)=e^{-i\frac{\pi}{2}\sigma_{x}}=\cos(\frac{\pi}{2})\mathbf{1}-i\sin(\frac{\pi}{2})\sigma_{x}=-i\sigma_{x}\propto U_{\Gamma}$ (46)
Thus, the endpoint coincides with the swap operation up to a global phase factor of i-i. This illustrates the physical goal of our construction: finding a continuous evolution $U(s)$ that connects the initial state to the final shift. While this finite-dimensional model implies trivial modular dynamics (and thus does not exhibit $\tilde{G}=2P$), it effectively demonstrates the necessity of a unitary path as a consistent alternative to the naive CP-map interpolation shown in Example B.1.
Appendix C Haagerup–Kosaki Lift: Existence and Mechanism of $\tilde{E}$
The most critical technical challenge in defining the idempotent path $E_{s}$ in the Type III setting is the construction of a faithful normal conditional expectation. In general, Type III inclusions do not admit a conditional expectation $\mathcal{M}\to\mathcal{N}$. The theory of Haagerup and Takesaki provides the necessary mechanism by lifting the problem from a non-tracial Type III algebra (equipped with an operator-valued weight, OVW) to a tracial semifinite algebra where a genuine conditional expectation $\tilde{E}$ exists. This lift effectively regularizes the unbounded OVW $\mathcal{E}$ via the crossed product construction.
Proposition C.1 (Haagerup–Kosaki lift and $\tilde{E}$ existence).
Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of Type III von Neumann algebras. While a conditional expectation may not exist, Haagerup’s theorem guarantees the existence of a faithful normal OVW $\mathcal{E}:\mathcal{M}^{+}\to\widehat{\mathcal{N}}^{+}$. Let $\omega$ be a faithful normal weight on $\mathcal{M}$ that is compatible with the inclusion, meaning its modular automorphism group $\sigma_{t}^{\omega}$ leaves the subalgebra invariant:
$\sigma_{t}^{\omega}(\mathcal{N})\subset\mathcal{N}\quad\text{for all }t\in\mathbb{R}.$ (47)
This condition implies that $\omega$ decomposes as $\omega=\psi\circ\mathcal{E}$ for some faithful normal weight $\psi$ on $\mathcal{N}$. Under this modular covariance assumption, the crossed product construction yields the following structure:
  1. 1.
[…]
    $\tau\circ\theta_{q}=e^{-q}\tau.$ (48)
    Furthermore, the connection to the base algebra is maintained by the canonical dual OVW $T:\tilde{\mathcal{M}}^{+}\to\widehat{\mathcal{M}}^{+}$, formally given by the integral of the dual action:
[…]
  2. 2.
    The subalgebra inclusion lifts to $\tilde{\mathcal{N}}=\mathcal{N}\rtimes_{\sigma^{\omega}|_{\mathcal{N}}}\mathbb{R}\subset\tilde{\mathcal{M}}$. Crucially, this lifted subalgebra remains invariant under the dual action:
    $\theta_{q}(\tilde{\mathcal{N}})=\tilde{\mathcal{N}}$. (50)
  3. 3.
    There exists a unique, faithful normal conditional expectation
    $\tilde{E}:\tilde{\mathcal{M}}\longrightarrow\tilde{\mathcal{N}}$ (51)
    which is trace-preserving (i.e., $\tau\circ\tilde{E}=\tau$) and is equivariant with respect to the dual action $\theta_{q}$ (i.e., $\tilde{E}\circ\theta_{q}=\theta_{q}\circ\tilde{E}$).
Outline of Construction.
The construction resolves the Type III pathology by embedding the algebra into a larger Type $\text{II}_{\infty}$ system. The key steps are:
  1. 1.
    Compatibility: The assumption $\sigma_{t}^{\omega}(\mathcal{N})\subset\mathcal{N}$ ensures that the modular dynamics of the ambient algebra do not mix the subalgebra with the complement, allowing $\tilde{\mathcal{N}}$ to be a well-defined subalgebra of $\tilde{\mathcal{M}}$.
  2. 2.
    Semifinite Structure: The theory of crossed products implies that $\tilde{\mathcal{M}}$ is semifinite. The scaling property (48) is the defining characteristic of a Type II crossed product derived from a Type III algebra, allowing us to treat modular automorphisms as trace-scaling operations.
  3. 3.
    Construction of $\tilde{E}$: The existence of the trace-preserving conditional expectation $\tilde{E}$ follows from the compatibility of the dual weights. Since the original weights satisfy $\omega=\psi\circ\mathcal{E}$, their associated dual traces on the crossed products are consistent. This consistency allows OVW $\mathcal{E}$ (which was unbounded) to be renormalized into a contractive projection $\tilde{E}$ on the semifinite algebra.
  4. 4.
    Equivariance: By construction, $\tilde{E}$ respects the dual action $\theta_{q}$. This equivariance is crucial for defining the analytic domains required for the interpolation $E_{s}$, ensuring that the path remains within the correct operator spaces.
For rigorous proofs and technical details, see TakesakiII ; Kosaki84 . ∎